### v2

# A $2$-distance $(2\Delta+7)$-coloring of planar graphs

Abstract.
A vertex coloring of a graph $G$ is called a $2$-distance coloring if any two vertices at a distance at most $2$ from each other receive different colors. Recently, Bousquet et al. (Discrete Mathematics, 346(4), 113288, 2023) proved that $72\Delta+$ 7 colors are sufficient for the $2$-distance coloring of planar graphs with maximum degree $\Delta\geq 9$. In this paper, we strengthen their result by removing the maximum degree constraint and show that all planar graphs admit a 2-distance $(2\Delta+7)$-coloring. This particularly improves the result of Van den Heuvel and McGuinness (Journal of Graph Theory, 42(2), 110-124, 2003).
Key words and phrases:
Coloring, 2-distance coloring, girth, planar graph.
1. Introduction
Among the numerous problems and concepts associated with graphs, coloring stands out as a fundamental and extensively studied topic. A vertex coloring of a graph involves assigning colors to its vertices to ensure that adjacent vertices have distinct colors. One particularly interesting variant of vertex coloring is 2-distance coloring, where any two vertices at a distance of two have distinct colors. Its motivation arises from the necessity in some real-life problems to assign different colors not only to adjacent vertices but also to those in close proximity [11, 14]. The concept of 2-distance coloring was first introduced in [9, 10], and it has received considerable attention due to the fact that certain problems, such as the Total Coloring Conjecture (see [4] for details), can be formulated as a 2-distance coloring of specific graphs. A comprehensive survey on 2-distance coloring and its related coloring concepts has been presented by Cranston [4].
The smallest number of colors for which graph $G$ admits a 2-distance coloring is known as the 2-distance chromatic number $\chi_{2}(G)$ of $G$. In 1977, Wegner [17] proposed the following conjecture.
Conjecture 1.1.
For every planar graph $G$ with maximum degree $\Delta$, $\chi_{2}(G)\leq 7$ if $\Delta=3$, $\chi_{2}(G)\leq\Delta+5$ if $4\leq\Delta\leq 7$, and $\chi_{2}(G)\leq\big{\lfloor}\frac{3\Delta}{2}\big{\rfloor}+1$ if $\Delta\geq 8$.
Wegner’s Conjecture is one of the most popular problems in graph coloring, and it has remained open for many years, except for the case when $\Delta=3$, which was solved by Thomassen [16] (independently by Hartke et al. [7]). For general planar graphs, Van den Heuvel and McGuinness [8] showed that $\chi_{2}(G)\leq 2\Delta+25$, while the bound $\chi_{2}(G)\leq\big{\lceil}\frac{5\Delta}{3}\big{\rceil}+78$ was proved by Molloy and Salavatipour [15]. On the other hand, some improved results are presented in [2, 3, 5, 6, 13] with a certain degree or girth restrictions. In particular, Bousquet et al. [1] recently showed that $\chi_{2}(G)\leq 2\Delta+7$ if $\Delta\geq 9$, which significantly improves upon the work in [12].
In this paper, we strengthen the result of Bousquet et al. [1] by removing the maximum degree constraint and show that all planar graphs admit a 2-distance $(2\Delta+7)$-coloring, which particularly improves the best-known bound of $252\Delta+$ 25 provided by Van den Heuvel and McGuinness [8].
Theorem 1.2.
For every planar graph $G$, we have $\chi_{2}(G)\leq 2\Delta+7$.
All graphs in this paper are assumed to be simple, and we refer to [18] for terminology and notation not defined here. When $G$ is a graph, we use $V(G),E(G),F(G),$ and $\Delta(G)$ to denote the vertex, edge and face set, and the maximum degree of $G$, respectively. If there is no confusion in the context, we abbreviate $\Delta(G)$ to $\Delta$. For a positive integer $t$, we denote by $[t]$ the set of integers between $1$ and $t$. Given a planar graph $G$, we denote by $\ell(f)$ the length of a face $f$ and by $d(v)$ the degree of a vertex $v$. A $k$-vertex is a vertex of degree $k$. A $k^{-}$-vertex is a vertex of degree at most $k$ while a $k^{+}$-vertex is a vertex of degree at least $k$. A $k$ ($k^{-}$ or $k^{+}$)-face is defined similarly. A vertex $u\in N(v)$ is called $k$-neighbour (resp. $k^{-}$-neighbour, $k^{+}$-neighbour) of $v$ if $d(u)=k$ (resp. $d(u)\leq k$, $d(u)\geq k$).
[…]
2. The Proof of Theorem 1.2
Let $G$ be a minimal counterexample to Theorem 1.2 such that $|V(G)|+|E(G)|$ is minimum. Thus, $G$ does not admit any $2$-distance $(2\Delta+7)$-coloring, but any planar graph $G^{\prime}$ obtained from $G$ with smaller $|V(G^{\prime})|+|E(G^{\prime})|$ admits a 2-distance $(2\Delta+7)$-coloring.
Clearly, $G$ is a connected graph. Due to the results of Bousquet et al. [1] and Zou et al. [19] (also [5]), we only need to consider the case $6\leq\Delta\leq 8$.
We call a graph $H$ proper with respect to $G$ if $H$ is obtained from $G$ by deleting some edges or vertices and adding some edges, ensuring that for every pair of vertices $x_{1}$ and $x_{2}$ in $V(G)\cap V(H)$ having distance at most 2 in $G$, they also have distance at most $2$ in $H$. If $f$ is a 2-distance coloring of such a graph $H$, then $f$ can be extended to the whole graph $G$, provided that each of the remaining uncolored vertices has an available color.
For a given vertex $v$ with $d(v)=k$, we assume that $v_{1},v_{2},\ldots,v_{k}$ are the neighbours of $v$ in clockwise order throughout this section.
2.1. The case $\Delta=6$
Recall that $G$ does not admit a $2$-distance $19$-coloring, but any planar graph $G^{\prime}$ obtained from $G$ with a smaller $|V(G^{\prime})|+|E(G^{\prime})|$ admits a 2-distance $19$-coloring.
Lemma 2.1.
We have $\delta(G)\geq 3$.
Proof.
If $v$ is a vertex of degree at most $2$, then $d_{2}(v)\leq 12$. Let $G^{\prime}$ be the graph obtained from $G-v$ by adding an edge between the vertices in $N(v)$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring $f$ with $19$ colors. Notice that $G^{\prime}$ is proper with respect to $G$. Since $d_{2}(v)\leq 12$, we can color $v$ with an available color, a contradiction. ∎
Lemma 2.2.
Let $v$ be a $3$-vertex. Then $m_{3}(v)=0$, $m_{4}(v)\leq 1$, and each neighbour of $v$ is a $6(4^{-})$-vertex.
Proof.
Let $v$ be a $3$-vertex, and denote by $v_{1},v_{2},v_{3}$ the neighbours of $v$ in clockwise order. Assume for a contradiction that $v$ is incident to a $3$-face, say $f_{1}=v_{1}vv_{2}$. It is clear that $d_{2}(v)\leq 16$. If we set $G^{\prime}=G-v+\{v_{1}v_{3}\}$, then $G^{\prime}$ would be proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring with $19$ colors. Since $d_{2}(v)\leq 16$, we can color $v$ with an available color, a contradiction.
Suppose now that $m_{4}(v)\geq 2$, i.e., $v$ is incident to two $4$-faces, say $f_{1}=v_{1}vv_{2}x$ and $f_{2}=v_{2}vv_{3}y$. Clearly, we have $d_{2}(v)\leq 16$. If we set $G^{\prime}=G-v+\{v_{1}v_{3}\}$, then $G^{\prime}$ would be proper with respect to $G$. Similarly as above, we get a contradiction.
Next we claim that each neighbour of $v$ is a $6(4^{-})$-vertex. Note that $v$ cannot have any $6(5^{+})$-neighbour since $m_{3}(v)=0$. Moreover, $v$ cannot have any $5^{-}$-neighbour. Indeed, if $v_{1}$ is such a vertex, then we set $G^{\prime}=G-v+\{v_{1}v_{2},v_{1}v_{3}\}$. In such a case, $d_{2}(v)\leq 17$, and $G^{\prime}$ is proper with respect to $G$. Similarly as above, we get a contradiction. Thus, each neighbour of $v$ is a $6(4^{-})$-vertex. ∎
We now address some vertices having some special forms: If $v$ is a $4(1)$- or a $4(2)$-vertex, then we call $v$ as a bad $4$-vertex. Similarly, if $v$ is a $5(4)$- or a 5(5)555(5)5 ( 5 )-vertex, then we call $v$ as a bad $5$-vertex.
Lemma 2.3.
If $v$ is a $4$-vertex, then $m_{3}(v)\leq 2$, and $v$ has no 6(6)666(6)6 ( 6 )-neighbour.
Proof.
Let $v$ be a $4$-vertex, and denote by $f_{1},f_{2},\ldots,f_{4}$ the faces incident to $v$. Assume for a contradiction that $m_{3}(v)\geq 3$. It follows that we have $f_{i}=v_{i}vv_{i+1}$ for $i\in[3]$. Clearly, $d_{2}(v)\leq 18$. If we set $G^{\prime}=G-v+\{v_{1}v_{4}\}$ (assuming $v_{1}v_{4}\notin E(G)$), then, by minimality, the graph $G^{\prime}$ has a 2-distance coloring with $19$ colors. Observe that every vertex in $V(G)$, except $v$, is colored with the same color that appears on the same vertex in $G^{\prime}$. Namely, $G^{\prime}$ is proper with respect to $G$. Since $d_{2}(v)\leq 18$, we assign $v$ with an available color, a contradiction.
Suppose now that $v$ has a 6(6)666(6)6 ( 6 )-neighbour. Without loss of generality, assume that $v_{2}$ is a 6(6)666(6)6 ( 6 )-vertex. In such a case, $v$ must be incident to two adjacent $3$-faces $f_{1},f_{2}$, say $f_{1}=v_{1}vv_{2}$ and $f_{2}=v_{2}vv_{3}$. Moreover, both $v_{1}v_{2}$ and $v_{2}v_{3}$ must be contained in two $3$-faces. This implies that $d_{2}(v)\leq 18$. If we set $G^{\prime}=G-v+\{v_{2}v_{4}\}$, then $G^{\prime}$ would be proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring with $19$ colors. Since $d_{2}(v)\leq 18$, we assign $v$ with an available color, a contradiction. ∎
Lemma 2.4.
Let $v$ be a $4(1)$-vertex. If $v$ has a $4$-neighbour, then the other neighbours of $v$ are different from $4$- and $5(4)$-vertices.
Proof.
Suppose that $v$ has a $4$-neighbour, say $v_{1}$, and assume by contradiction that $v$ has another neighbour consisting of $4$- or $5(4)$-vertices. In this case, we have $d_{2}(v)\leq 18$. If we set $G^{\prime}=G-v+\{v_{1}v_{2},v_{1}v_{3},v_{1}v_{4}\}$, then $G^{\prime}$ would be proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring with $19$ colors. Since $d_{2}(v)\leq 18$, we can color $v$ with an available color, a contradiction. ∎
Lemma 2.5.
Let $v$ be a $4(1)$-vertex. If $m_{4}(v)=r$ for $0\leq r\leq 3$, then $n_{6}(v)\geq r+1$.
Proof.
Let $v$ be a $4(1)$-vertex, and denote by $f_{1},f_{2},\ldots,f_{4}$ the faces incident to $v$. Suppose that $f_{1}$ is a $3$-face with $f_{1}=v_{1}vv_{2}$. Let $m_{4}(v)=r$ for $0\leq r\leq 3$. Assume for a contradiction that $n_{6}(v)\leq r$. We then deduce that $d_{2}(v)\leq 18$, since $v$ has $4-r$ many $5^{-}$-neighbours, and $v$ is incident to $r$ many $4$-faces. This particularly implies that either $v$ has two $5^{-}$-neighbours or $v$ is incident to three $4$-faces.
If $v$ has two $5^{-}$-neighbours $v_{i},v_{j}$ such that $\{i,j\}\cap\{1,2\}\neq\emptyset$, then we set $G^{\prime}=G-v+\{v_{k}v_{3},v_{k}v_{4}\}$ for $k\in\{i,j\}\cap\{1,2\}$. If $v$ has two $5^{-}$-neighbours $v_{i},v_{j}$ with $i<j$ such that $\{i,j\}\cap\{1,2\}=\emptyset$, then we set $G^{\prime}=G-v+\{v_{i}v_{j},v_{i}v_{2},v_{j}v_{1}\}$. If $v$ is incident to exactly three $4$-faces, then we set $G^{\prime}=G-v+\{v_{2}v_{3},v_{1}v_{4}\}$. In each case, we observe that $G^{\prime}$ is proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring with $19$ colors. Since $d_{2}(v)\leq 18$, we can color $v$ with an available color, a contradiction. ∎
Lemma 2.6.
Let $v$ be a $4(2)$-vertex. Then $m_{4}(v)\leq 1$. In particular, if $m_{4}(v)=r$ for $0\leq r\leq 1$, then $n_{6}(v)\geq r+3$.
Proof.
Let $v$ be a $4(2)$-vertex, and denote by $f_{1},f_{2},\ldots,f_{4}$ the faces incident to $v$. Suppose that $f_{1}$ and $f_{2}$ are $3$-faces such that $f_{1}=v_{i}vv_{i+1}$ and $f_{2}=v_{j}vv_{j+1}$ with $i<j$ and $i,j\in[4]$ in a cyclic fashion. Assume for a contradiction that $v$ is incident to two $4$-faces. In this case, we have $d_{2}(v)\leq 18$. If $f_{1},f_{2}$ are adjacent, then we set $G^{\prime}=G-v+\{v_{j}v_{k}\}$ for $k\in N(v)\setminus\{v_{i},v_{j},v_{j+1}\}$. Else, we set $G^{\prime}=G-v+\{v_{i}v_{j+1},v_{i+1}v_{j}\}$. In both cases, $G^{\prime}$ is proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring with $19$ colors. Since $d_{2}(v)\leq 18$, we can color $v$ with an available color, a contradiction.
On the other hand, if $m_{4}(v)=0$ and $n_{6}(v)\leq 2$, then $v$ has two $5^{-}$-neighbours, and so $d_{2}(v)\leq 18$. Similarly, if $m_{4}(v)=1$ and $n_{6}(v)\leq 3$, then $v$ has a $5^{-}$-neighbour, and so $d_{2}(v)\leq 18$, as $m_{4}(v)=1$. By applying the same argument as above, we get a contradiction. ∎
Lemma 2.7.
Let $v$ be a $4(2)$-vertex. Then, $v$ has neither a $4$-neighbour nor a bad $5$-neighbour.
Proof.
Let $v$ be a $4$-vertex with $m_{3}(v)=2$. Suppose first that one of $v_{i}$’s is a $4$-vertex, say $v_{1}$. Obviously, we have $d_{2}(v)\leq 18$, and let $G^{\prime}=G-v+\{v_{1}v_{2},v_{1}v_{3},v_{1}v_{4}\}$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring $f$ with $19$ colors. Observe that $G^{\prime}$ is proper with respect to $G$. Since $d_{2}(v)\leq 18$, we can color $v$ with an available color, a contradiction.
Suppose now that $v$ has a bad $5$-neighbour, say $v_{i}$. Then, there exists an edge $v_{i}v_{j}$ for $v_{j}\in N(v)$ such that $v_{i}v_{j}$ is contained in two $3$-faces. It follows that we have $d_{2}(v)\leq 18$. If we set $G^{\prime}=G-v+\{v_{i}v_{p},v_{i}v_{t}\}$ for $v_{p},v_{t}\in N(v)\setminus\{v_{i},v_{j}\}$, then $G^{\prime}$ would be proper with respect to $G$. Similarly as above, we get a contradiction. ∎
Lemma 2.8.
Let $v$ be a $5(4)$-vertex. If $v$ has a $4$-neighbour, then the other neighbours of $v$ are different from $4$- and $5(4)$-vertices.
Proof.
Let $v$ be a $5(4)$-vertex and denote by $f_{1},f_{2},\ldots,f_{5}$ the faces incident to $v$ such that $f_{i}=v_{i}vv_{i+1}$ for $i\in[4]$. Suppose that $v$ has a $4$-neighbour, and assume by contradiction that $v$ has another neighbour consisting of $4$- or $5(4)$-vertices. We then deduce that $d_{2}(v)\leq 18$. If we set $G^{\prime}=G-v+\{v_{1}v_{5}\}$, then $G^{\prime}$ would be proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring with $19$ colors. Since $d_{2}(v)\leq 18$, we can color $v$ with an available color, a contradiction. ∎
The following is a direct consequence of Lemma 2.8, since a $5(4)$-vertex cannot have both a $4$-neighbour and a $5(4)$-neighbour simultaneously.
Corollary 2.9.
If $v$ is a $4(1)$-vertex, then $v$ has no two $5(4)$-neighbours.
Lemma 2.10.
Let $v$ be a $5(4)$-vertex.
  • $(a)$
    If $m_{4}(v)=0$, then $v$ has two $6(5^{-})$-neighbours.
  • $(b)$
    If $m_{4}(v)=1$, then $v$ has three $6(5^{-})$-neighbours.
Proof.
Let $v$ be a $5(4)$-vertex and denote by $f_{1},f_{2},\ldots,f_{5}$ the faces incident to $v$ such that $f_{i}=v_{i}vv_{i+1}$ for $i\in[4]$.
$(a)$ Let $m_{4}(v)=0$. Assume for a contradiction that $v$ has at most one $6(5^{-})$-neighbour. In this case, $v$ has four neighbours consisting of 6(6)666(6)6 ( 6 )- or $5^{-}$-vertices, and so we have $d_{2}(v)\leq 18$. If we set $G^{\prime}=G-v+\{v_{1}v_{5}\}$, then $G^{\prime}$ is proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring with $19$ colors. Since $d_{2}(v)\leq 18$, we can color $v$ with an available color, a contradiction.
$(b)$ Let $m_{4}(v)=1$. By contradiction, assume that $v$ has at most two $6(5^{-})$-neighbours. In this case, $v$ has three neighbours consisting of 6(6)666(6)6 ( 6 )- or $5^{-}$-vertices, and so $d_{2}(v)\leq 18$. Similarly as above, we get a contradiction. ∎
Proposition 2.11.
A $5(4)$-vertex cannot have two non-adjacent $5(4)$-neighbours.
Proof.
Let $v$ be a $5(4)$-vertex and denote by $f_{1},f_{2},\ldots,f_{5}$ the faces incident to $v$ such that $f_{i}=v_{i}vv_{i+1}$ for $i\in[4]$.
Assume that $v$ has two $5(4)$-neighbours $v_{i},v_{j}$ such that $|i-j|\geq 2$. Then $v_{i}v_{i-1}$ or $v_{i}v_{i+1}$ (in a cyclic ordering) is contained in two $3$-faces. Moreover, $v_{j}v_{j-1}$ or $v_{j}v_{j+1}$ (in a cyclic ordering) is contained in two $3$-faces. We then deduce that $d_{2}(v)\leq 18$. If we set $G^{\prime}=G-v+\{v_{1}v_{5}\}$, then $G^{\prime}$ would be proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring with $19$ colors. Since $d_{2}(v)\leq 18$, we can color $v$ with an available color, a contradiction. ∎
Lemma 2.12.
Let $v$ be a 5(5)555(5)5 ( 5 )-vertex. Then $v$ has neither $4$-neighbour nor bad $5$-neighbour nor 6(6)666(6)6 ( 6 )-neighbour. In particular, $n_{5}(v)\leq 1$.
Proof.
Let $v$ be a 5(5)555(5)5 ( 5 )-vertex and denote by $f_{1},f_{2},\ldots,f_{5}$ the $3$-faces incident to $v$ such that $f_{i}=v_{i}vv_{i+1}$ for $i\in[5]$ in a cyclic fashion.
Recall that $G$ has no $4(3^{+})$-vertex by Lemma 2.3, also $v$ cannot have any $4(1^{-})$-neighbour as $m_{3}(v)=5$. Moreover, $v$ cannot have any $4(2)$-neighbour by Lemma 2.7. Consequently, we deduce that $v$ has no $4$-neighbour. On the other hand, if $v$ has a bad $5$-neighbour or a 6(6)666(6)6 ( 6 )-neighbour or two $5$-neighbours, then we have $d_{2}(v)\leq 18$. When we set $G^{\prime}=G-v$, the graph $G^{\prime}$ remains proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring with $19$ colors. Since $d_{2}(v)\leq 18$, we can color $v$ with an available color, a contradiction. ∎
The following is an easy consequence of Lemma 2.12.
Corollary 2.13.
If $v$ is a 5(5)555(5)5 ( 5 )-vertex, then $v$ has four $6(5^{-})$-neighbours.
An edge $uv$ is said to be special if $v$ is a 5(5)555(5)5 ( 5 )-vertex, and $uv$ is contained in two $3$-faces $f_{1},f_{2}$ such that each of those faces is adjacent to a $4^{+}$-face (see Figure 1). Recall that $u$ is a $4^{+}$-vertex by Lemma 2.2.
[…]
Figure 1. A special edge $uv$ with a 5(5)555(5)5 ( 5 )-vertex $v$ and a $4^{+}$-vertex $u$.
Proposition 2.14.
Let $v$ be a 5(5)555(5)5 ( 5 )-vertex, and let $uw$ be an edge for $u,w\in N(v)$.
  • $(a)$
    If $uw$ is contained in two $3$-faces, then all neighbours of $v$ are $6(5^{-})$-vertices. In particular, for every vertex $z\in N(v)\setminus\{u,w\}$, the edge $vz$ is a special edge.
  • $(b)$
    If no edge in $G[N(v)]$ is contained in two $3$-faces, then $v$ has four $6(4^{-})$-neighbours, and each $vv_{i}$ is a special edge for $v_{i}\in N(v)$.
Proof.
Let $v$ be a 5(5)555(5)5 ( 5 )-vertex, and denote by $f_{1},f_{2},\ldots,f_{5}$ the $3$-faces incident to $v$ such that $f_{i}=v_{i}vv_{i+1}$ for $i\in[5]$ in a cyclic fashion.
$(a)$ Suppose first that there exists $v_{i}\in N(v)$ such that $v_{i}v_{i+1}$ is contained in two $3$-faces. If $v$ has a $5^{-}$- or 6(6)666(6)6 ( 6 )-neighbours, then $d_{2}(v)\leq 18$. When we set $G^{\prime}=G-v$, the graph $G^{\prime}$ remains proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring with $19$ colors. Since $d_{2}(v)\leq 18$, we can color $v$ with an available color, a contradiction. Thus, all neighbours of $v$ are $6(5^{-})$-vertices. In particular, if there also exists $v_{j}\in N(v)\setminus\{v_{i},v_{i+1}\}$ such that $v_{j}v_{j-1}$ or $v_{j}v_{j+1}$ is contained in two $3$-faces, then we have again $d_{2}(v)\leq 18$. Similarly as above, we get a contradiction. Therefore, for every vertex $v_{j}\in N(v)\setminus\{v_{i},v_{i+1}\}$, the edge $vv_{j}$ is a special edge.
$(b)$ Suppose that no edge in $G[N(v)]$ is contained in two $3$-faces. This implies that each $vv_{i}$ is a special edge for $v_{i}\in N(v)$. So, the claim follows from Corollary 2.13. ∎
Lemma 2.15.
Let $v$ be a $6(5)$-vertex. Then $n_{4}(v)\leq 3$. In particular,
  • $(a)$
    if $n_{4}(v)=3$, then $v$ has no bad $5$-neighbour,
  • $(b)$
    if $n_{4}(v)=2$, then $v$ has at most two bad $5$-neighbours,
  • $(c)$
    if $n_{4}(v)=2$ and $v$ has two bad $5$-neighbours, then $v$ has two $6(4^{-})$-neighbours,
  • $(d)$
    if $n_{4}(v)=1$, then $v$ has at most three bad $5$-neighbours,
  • $(e)$
    if $n_{4}(v)=1$ and $v$ has three bad $5$-neighbours, then $v$ has a $6(4^{-})$-neighbour,
  • $(f)$
    if $n_{4}(v)=0$, then $v$ has at most four bad $5$-neighbours,
  • $(g)$
    if $n_{4}(v)=0$, $m_{4}(v)=1$ and $v$ has four bad $5$-neighbours, then $v$ has two $6(4^{-})$-neighbours.
Proof.
Let $v$ be a $6(5)$-vertex, and denote by $f_{1},f_{2},\ldots,f_{6}$ the faces incident to $v$ such that $f_{i}=v_{i}vv_{i+1}$ for $i\in[5]$. Recall first that $v$ has no $3$-neighbour by Lemma 2.2.
We now claim that $n_{4}(v)\leq 3$. Otherwise, $v$ has four $4$-neighbours, and so $d_{2}(v)\leq 18$. By assumption, one of $v_{2},v_{3},v_{4}$ must be $4$-vertex, say $v_{2}$. If we set $G^{\prime}=G-v+\{v_{2}v_{4},v_{2}v_{5},v_{2}v_{6}\}$, then $G^{\prime}$ would be proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring $f$ with $19$ colors. Since $d_{2}(v)\leq 18$, we can color $v$ with an available color, a contradiction. Hence, $n_{4}(v)\leq 3$.
$(a)$ Let $n_{4}(v)=3$. By contradiction, assume that $v$ has a bad $5$-neighbour, say $v_{i}$. This forces that either $v_{i}v_{i-1}$ or $v_{i}v_{i+1}$ is contained in two $3$-faces. In each case, we have $d_{2}(v)\leq 18$. By assumption, one of $v_{2},v_{3},v_{4},v_{5}$ is a $4$-vertex, say $v_{2}$. Let $G^{\prime}=G-v+\{v_{2}v_{4},v_{2}v_{5},v_{2}v_{6}\}$. Clearly, $G^{\prime}$ is proper with respect to $G$. Similarly as above, we get a contradiction.
$(b)$ Let $n_{4}(v)=2$. Assume to the contrary that $v$ has three bad $5$-neighbours. This forces that there exist two distinct vertices $v_{i},v_{j}\in N(v)$ such that $v_{i}v_{i+1}$ and $v_{j}v_{j+1}$ are contained in two $3$-faces. It is clear that $d_{2}(v)\leq 17$. By assumption, $v$ has two $4$-neighbours. If one of $v_{2},v_{3},v_{4},v_{5}$ is a $4$-vertex, say $v_{2}$, then we set $G^{\prime}=G-v+\{v_{2}v_{4},v_{2}v_{5},v_{2}v_{6}\}$, and so $G^{\prime}$ would be proper with respect to $G$. Otherwise, suppose that both $v_{1}$ and $v_{6}$ are $4$-vertices. Consider a $5$-neighbour of $v$, say $v_{2}$, if we set $G^{\prime}=G-v+\{v_{1}v_{6},v_{2}v_{4},v_{2}v_{6}\}$, then $G^{\prime}$ would be proper with respect to $G$. Similarly as above, we get a contradiction.
$(c)$ Let $n_{4}(v)=2$, and suppose that $v$ has two bad $5$-neighbours, say $v_{i},v_{j}$. This forces that either $v_{i}v_{i-1}$ or $v_{i}v_{i+1}$ is contained in two $3$-faces. Moreover, if $v_{i}$ and $v_{j}$ are non-adjacent, then there exists another edge forming $v_{j}v_{j-1}$ or $v_{j}v_{j+1}$ such that it is contained in two $3$-faces. We remark that if $v$ has at most one $6(4^{-})$-neighbour, then the last neighbour of $v$ other than $v_{i},v_{j}$, $4$-vertex and $6(4^{-})$-vertex would be either $5$-vertex or a $6(5)$-vertex or a 6(6)666(6)6 ( 6 )-vertex. In each case, we have $d_{2}(v)\leq 18$. By applying a similar process to $(b)$, we get a contradiction.
$(d)$ Let $n_{4}(v)=1$. Assume that $v$ has four bad $5$-neighbours. This forces that there exist two distinct vertices $v_{i},v_{j}\in N(v)$ with $|i-j|>1$ such that $v_{i}v_{i+1}$ and $v_{j}v_{j+1}$ are contained in two $3$-faces. Clearly we have $d_{2}(v)\leq 18$. By assumption, $v$ has a $4$-vertex. If one of $v_{2},v_{3},v_{4},v_{5}$ is a $4$-vertex, say $v_{2}$, then we set $G^{\prime}=G-v+\{v_{2}v_{4},v_{2}v_{5},v_{2}v_{6}\}$, and so $G^{\prime}$ would be proper with respect to $G$. Otherwise, suppose that $v_{1}$ (or $v_{6}$) is a $4$-vertex. Since $v$ has four $5$-neighbours, one of $v_{3},v_{5}$ must be $5$-vertex, say $v_{3}$. Let $G^{\prime}=G-v+\{v_{1}v_{6},v_{3}v_{1},v_{3}v_{5}\}$. Clearly, $G^{\prime}$ is proper with respect to $G$. Similarly as above, we get a contradiction.
$(e)$ Let $n_{4}(v)=1$, and suppose that $v$ has three bad $5$-neighbours, say $v_{i},v_{j},v_{k}$. This forces that there exist two distinct vertices $v_{p},v_{q}\in N(v)$ such that $v_{p}v_{p+1}$ and $v_{q}v_{q+1}$ are contained in two $3$-faces. If $v$ has no $6(4^{-})$-neighbour, then the last two neighbours of $v$ other than $v_{i},v_{j},v_{k}$ and $4$-vertex consist of $5$-, $6(5)$- or 6(6)666(6)6 ( 6 )-vertices. In each case, we have $d_{2}(v)\leq 18$. Similar to $(d)$, we get a contradiction.
$(f)$ Let $n_{4}(v)=0$. Assume to the contrary that $v$ has five bad $5$-neighbours. This forces that there exist three distinct vertices $v_{i},v_{j},v_{t}\in N(v)$ such that each of $v_{i}v_{i+1}$, $v_{j}v_{j+1}$, $v_{t}v_{t+1}$ is contained in two $3$-faces. Clearly, we have $d_{2}(v)\leq 18$. Since $v$ has five bad $5$-neighbours, one of $v_{1},v_{6}$ is a $5$-vertex, say $v_{1}$. Moreover, one of $v_{3},v_{5}$ is a $5$-vertex, say $v_{3}$. Let $G^{\prime}=G-v+\{v_{1}v_{6},v_{3}v_{1},v_{3}v_{5}\}$. Clearly, $G^{\prime}$ is proper with respect to $G$. Similarly as above, we get a contradiction.
$(g)$ Suppose that $n_{4}(v)=0$, $m_{4}(v)=1$, and $v$ has four bad $5$-neighbours. Since $v$ has four bad $5$-neighbours, there exist two distinct vertices $v_{i},v_{j}\in N(v)$ with $|i-j|>1$ such that $v_{i}v_{i+1}$ and $v_{j}v_{j+1}$ are contained in two $3$-faces. By contradiction, assume that $v$ has at most one $6(4^{-})$-neighbour. Then the sixth neighbour of $v$ other than bad $5$-vertex and $6(4^{-})$-vertex is either a $5(t)$-vertex with $t\leq 3$ or a $6(5^{+})$-vertex. Clearly, we have $d_{2}(v)\leq 18$. Moreover, this particularly implies that either $v$ has totally five $5$-neighbours or there exist three distinct vertices $v_{i},v_{j},v_{t}\in N(v)$ such that each of $v_{i}v_{i+1}$, $v_{j}v_{j+1}$, $v_{t}v_{t+1}$ is contained in two $3$-faces. If $v$ has five $5$-neighbours, then one of $v_{1},v_{6}$ must be a $5$-vertex, say $v_{1}$. Also, one of $v_{3},v_{5}$ must be a $5$-vertex, say $v_{3}$. Let $G^{\prime}=G-v+\{v_{1}v_{6},v_{3}v_{1},v_{3}v_{5}\}$. Clearly, $G^{\prime}$ is proper with respect to $G$. Similarly as above, we get a contradiction.
We now suppose that $v$ has exactly four $5$-neighbours. If $v_{1}$ (resp. $v_{6}$) is a $5$-vertex, and one of $v_{3},v_{5}$ (resp. $v_{2},v_{4}$) is a $5$-vertex, then we apply the same argument as above. Therefore, we may assume that $v_{1},v_{6}$ are $6$-vertices and $v_{2},v_{3},v_{4},v_{5}$ are bad $5$-vertices. In fact, $v_{2},v_{3},v_{4},v_{5}$ must be $5(4)$-vertices, since a 5(5)555(5)5 ( 5 )-vertex cannot have any bad $5$-neighbour by Lemma 2.12. However, this yields a contradiction by Proposition 2.11. ∎
We will now apply discharging to show that $G$ does not exist. First, we assign to each vertex $v$ a charge $\mu(v)=d(v)-4$ and to each face $f$ a charge $\mu(f)=\ell(f)-4$. By Euler’s formula, we have
$\sum_{v\in V(G)}\left(d(v)-4\right)+\sum_{f\in F(G)}(\ell(f)-4)=-8.$
We next present some rules and redistribute accordingly. Once the discharging finishes, we check the final charge $\mu^{*}(v)$ and $\mu^{*}(f)$. If $\mu^{*}(v)\geq 0$ and $\mu^{*}(f)\geq 0$, we get a contradiction that no such a counterexample can exist.
Discharging Rules
We apply the following discharging rules.
  • R1:
    Every $3$-face receives $\frac{1}{3}$ from each of its incident vertices.
  • R2:
    Let $f$ be a $5^{+}$-face. Then,
    • $(a)$
      $f$ gives $\frac{1}{3}$ to each of its incident $3$-vertices and bad $4$-vertices,
    • $(b)$
      $f$ gives $\frac{1}{5}$ to each of its incident $5(4)$-vertices,
    • $(c)$
      $f$ gives $\frac{1}{5}$ to each of its incident $6(5)$-vertex $u$ if $n_{4}(u)=0$.
  • R3:
    Every $6(k)$-vertex for $0\leq k\leq 3$ gives $\frac{1}{6}$ to each of its neighbours.
  • R4:
    Let $v$ be a $6(4)$-vertex. Then,
    • $(a)$
      $v$ gives $\frac{1}{12}$ to each of its $4$- and $6(5)$-neighbours,
    • $(b)$
      $v$ gives $\frac{1}{9}$ to each of its $3$- and $5(4)$-neighbours,
    • $(c)$
      $v$ gives $\frac{1}{6}$ to each of its 5(5)555(5)5 ( 5 )-neighbour $u$ if $uv$ is a special edge,
    • $(d)$
      $v$ gives $\frac{1}{9}$ to each of its 5(5)555(5)5 ( 5 )-neighbour $u$ if $uv$ is not a special edge.
  • R5:
    Let $v$ be a $6(5)$-vertex. Then,
    • $(a)$
      $v$ gives $\frac{1}{12}$ to each of its bad $4$-neighbours,
    • $(b)$
      $v$ gives $\frac{1}{9}$ to each of its bad $5$-neighbours.
Checking $\mu^{*}(v),\mu^{*}(f)\geq 0$ for $v\in V(G),f\in F(G)$.
First we show that $\mu^{*}(f)\geq 0$ for each $f\in F(G)$. Given a face $f\in F(G)$, if $f$ is a $3$-face, then it receives $\frac{1}{3}$ from each of incident vertices by R1, and so $\mu^{*}(f)=-1+3\times\frac{1}{3}=0$. If $f$ is a $4$-face, then $\mu(f)=\mu^{*}(f)=0$.
Let $f$ be a $5$-face. Suppose first that $f$ is not incident to any $3$-vertex. Note that if $f$ is not incident to any bad $4$-vertices, then $\mu^{*}(v)\geq 1-5\times\frac{1}{5}=0$ after $f$ sends at most $\frac{1}{5}$ to each of its incident vertices by R2(b),(c). Therefore we further suppose that $f$ is incident to at least one bad $4$-vertex, let $u$ be such a bad $4$-vertex incident to $f$. Recall that $u$ has no 6(6)666(6)6 ( 6 )-neighbour by Lemma 2.3. If $u$ is a $4(1)$-vertex, then it has at most one neighbour consisting of bad $4$-vertex or $5(4)$-vertex by Lemma 2.4 and Corollary 2.9. Besides, if $u$ is a $4(2)$-vertex, then the neighbours of $u$ are different from $4$- and $5(4)$-vertices by Lemma 2.7. Consequently, if $f$ is incident to three bad $4$-vertices, then the other vertices incident to $f$ are different from $4$-, $5(4)$- and 6(6)666(6)6 ( 6 )-vertices. Moreover, if $f$ is incident to exactly two bad $4$-vertices, then $f$ is incident to at most one vertex consisting of $5(4)$- or 6(6)666(6)6 ( 6 )-vertices. On the other hand, if $f$ is incident to exactly one bad $4$-vertex $u$, then $u$ has no 6(6)666(6)6 ( 6 )-neighbour, and has at most one $5(4)$-neighbour, so $f$ is incident to at most three vertices consisting of $5(4)$- or 6(6)666(6)6 ( 6 )-vertices. We then conclude that $\mu^{*}(f)\geq 0$, since the face $f$ sends $\frac{1}{3}$ to each of its incident bad $4$-vertices by R2(a), $\frac{1}{5}$ to each of its incident $5(4)$-vertices by R2(b) and $\frac{1}{5}$ to each of its incident 6(6)666(6)6 ( 6 )-vertices having no $4$-neighbour by R2(c).
Now we suppose that $f$ is incident to a $3$-vertex. By Lemma 2.2, $f$ can be incident to at most two $3$-vertices. If $f$ is incident to exactly one $3$-vertex, say $v_{1}$, then the neighbours of $v_{1}$ lying on $f$, denote by $v_{2},v_{5}$, are different from $5^{-}$- and $6(5^{+})$-vertex by Lemma 2.2. Thus, $\mu^{*}(v)\geq 1-3\times\frac{1}{3}=0$ after $f$ sends at most $\frac{1}{3}$ to each of its incident vertices other than $v_{2},v_{5}$ by R2. Next we assume that $f$ is incident to exactly two $3$-vertices, say $v_{1},v_{3}$, then all vertices incident to $f$ other than $v_{1},v_{3}$ are different from $5^{-}$- and $6(5^{+})$-vertex by Lemma 2.2. Thus, $\mu^{*}(v)\geq 1-2\times\frac{1}{3}>0$ after $f$ sends $\frac{1}{3}$ to each of $v_{1},v_{3}$ by R2(a).
On the other hand, if $f$ is a $6^{+}$-face, then we have $\mu^{*}(v)\geq\ell(f)-4-\ell(f)\times\frac{1}{3}\geq 0$ after $v$ sends at most $\frac{1}{3}$ to each of its incident vertices by R2.
We now pick a vertex $v\in V(G)$ with $d(v)=k$. By Lemma 2.1, we have $k\geq 3$.
(1). Let $k=3$. By Lemma 2.2, $m_{3}(v)=0$, $m_{4}(v)\leq 1$, and each neighbour of $v$ is a $6(4^{-})$-vertex. Note that $v$ is incident to at least two $5^{+}$-faces, and each of them gives $\frac{1}{3}$ to $v$ by R2(a). On the other hand, $v$ is incident to three $6(4^{-})$-vertices, and each of them gives at least $\frac{1}{9}$ to $v$ by R3 and R4(b). Thus, we have $\mu^{*}(v)\geq-1+2\times\frac{1}{3}+3\times\frac{1}{9}\geq 0$.
(2). Let $k=4$. By Lemma 2.3, $v$ is incident to at most two $3$-faces, in particular, $v$ has no 6(6)666(6)6 ( 6 )-neighbour. Clearly, we have $\mu^{*}(v)\geq 0$ when $v$ is not incident to any $3$-face, since $v$ does not give a charge to any vertex or face. We may therefore assume that $1\leq m_{3}(v)\leq 2$; namely, $v$ is a bad $4$-vertex.
First suppose that $m_{3}(v)=1$. Let $m_{4}(v)=r$ for $0\leq r\leq 3$. By Lemma 2.5, $n_{6}(v)\geq r+1$, i.e., $v$ has at least $r+1$ many $6(5^{-})$-neighbours, and each of them gives at least $\frac{1}{12}$ to $v$ by R3, R4(a) and R5(a). On the other hand, $v$ is incident to $3-r$ many $5^{+}$-faces, and each of them gives $\frac{1}{3}$ to $v$ by R2(a). Consequently, $v$ receives totally $(3-r)\times\frac{1}{3}$ from its incident $5^{+}$-faces, and totally $(r+1)\times\frac{1}{12}$ from its $6$-neighbours. Thus, $\mu^{*}(v)\geq(3-r)\times\frac{1}{3}+(r+1)\times\frac{1}{12}-\frac{1}{3}\geq 0$ after $v$ sends $\frac{1}{3}$ to its incident $3$-face by R1.
We now suppose that $m_{3}(v)=2$, and let $m_{4}(v)=r$. Then, by Lemma 2.6, we have $0\leq r\leq 1$ and $n_{6}(v)\geq r+3$. Since $v$ is not adjacent to any 6(6)666(6)6 ( 6 )-vertex, $v$ has at least $r+3$ many $6(5^{-})$-neighbours, and each of them gives at least $\frac{1}{12}$ to $v$ by R3, R4(a) and R5(a). On the other hand, $v$ is incident to $2-r$ many $5^{+}$-faces, and each of them gives $\frac{1}{3}$ to $v$ by R2(a). Consequently, $v$ receives totally $(2-r)\times\frac{1}{3}$ from its incident $5^{+}$-faces, and totally $(r+3)\times\frac{1}{12}$ from its $6$-neighbours. Thus, $\mu^{*}(v)\geq(2-r)\times\frac{1}{3}+(r+3)\times\frac{1}{12}-2\times\frac{1}{3}\geq 0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
(3). Let $k=5$. We first note that if $m_{3}(v)\leq 3$, then $\mu^{*}(v)\geq 1-3\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1. Therefore, we may assume that $m_{3}(v)\geq 4$; namely, $v$ is a bad $5$-vertex.
Let $m_{3}(v)=4$. If $m_{4}(v)=0$, then $v$ has at least two $6(5^{-})$-neighbours by Lemma 2.10(a), and $v$ receives $\frac{1}{9}$ from each of its $6(5^{-})$-neighbours by R3, R4(b) and R5(b). Also, $v$ receives $\frac{1}{5}$ from its incident $5^{+}$-face by R2(b). Thus, $\mu^{*}(v)\geq 1+2\times\frac{1}{9}+\frac{1}{5}-4\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1. Suppose now that $m_{4}(v)=1$. Then $v$ has at least three $6(5^{-})$-neighbours by Lemma 2.10(b). It follows from applying R3, R4(b) and R5(b) that $v$ receives at least $\frac{1}{9}$ from each of its $6(5^{-})$-neighbours. Thus, $\mu^{*}(v)\geq 1+3\times\frac{1}{9}-4\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
Let $m_{3}(v)=5$. Recall that $v$ has no 6(6)666(6)6 ( 6 )-neighbour by Lemma 2.12. Suppose first that there exists $uw\in G[N(v)]$ such that $uw$ is contained in two $3$-faces. It then follows from Proposition 2.14(a) that all neighbours of $v$ are $6(5^{-})$-vertices, and each $vv_{i}$ for $v_{i}\in N(v)\setminus\{u,w\}$ is a special edge. This particularly implies that each $v_{i}\in N(v)\setminus\{u,w\}$ is a $6(4^{-})$-vertex. Then, $v$ receives $\frac{1}{6}$ from each $v_{i}$ for $i\in[3]$ by R3 and R4(c), and $\frac{1}{9}$ from each of $u,w$ by R4(d) and R5(b). Thus, $v$ totally receives at least $\frac{2}{3}$ from its $6$-neighbours. Next we suppose that no edge $uw\in G[N(v)]$ is contained in two $3$-faces. By Proposition 2.14(b), there exist four $6(4^{-})$-vertices $v_{1},v_{2},v_{3},v_{4}\in N(v)$ such that each $vv_{i}$ is a special edge. Then, $v$ receives $\frac{1}{6}$ from each $v_{i}$ for $i\in[4]$ by R3 and R4(c). Clearly, $v$ receives totally at least $\frac{2}{3}$ from its $6$-neighbours. Hence, $\mu^{*}(v)\geq 1+\frac{2}{3}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
(4). Let $k=6$. Notice first that if $m_{3}(v)\leq 3$, then $\mu^{*}(v)\geq 2-3\times\frac{1}{3}-6\times\frac{1}{6}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and at most $\frac{1}{6}$ to each of its neighbours by R3. In addition, if $m_{3}(v)=6$, then $\mu^{*}(v)\geq 2-6\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1. Therefore we may assume that $4\leq m_{3}(v)\leq 5$.
(4.1). Let $m_{3}(v)=4$. Obviously, $v$ has one of the three configurations depicted in Figure 2. Notice that $v$ only gives charge to its incident $3$-faces, and $4^{-}$-, $5(4^{+})$-, $6(5)$-neighbours by R1, R4. We first note that if $v$ has no 5(5)555(5)5 ( 5 )-neighbour $v_{i}$ such that $vv_{i}$ is a special edge, then $v$ gives at most $\frac{1}{9}$ to each of its neighbours by R4(a),(b),(d), and so $\mu^{*}(v)\geq 2-6\times\frac{1}{9}-4\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1. We may further assume that $v$ has at least one 5(5)555(5)5 ( 5 )-neighbour $v_{i}$ such that $vv_{i}$ is a special edge. In fact, we will determine the final charge of $v$ based on the number of 5(5)555(5)5 ( 5 )-neighbors $v_{i}$ of $v$ for which each edge $vv_{i}$ is a special edge. It can be easily observed that $v$ cannot have three 5(5)555(5)5 ( 5 )-neighbours by Lemma 2.12.
$v$$v_{3}$$v_{4}$$v_{5}$$v_{6}$$v_{1}$$v_{2}$
$v$$v_{3}$$v_{4}$$v_{5}$$v_{6}$$v_{1}$$v_{2}$
$v$$v_{3}$$v_{4}$$v_{5}$$v_{6}$$v_{1}$$v_{2}$
Figure 2. Three configurations of $6(4)$-vertices.
Suppose first that $v$ has two 5(5)555(5)5 ( 5 )-neighbours $v_{i}$ and $v_{j}$ such that both $vv_{i}$ and $vv_{j}$ are special edges. By Lemma 2.12, $v_{i}$ and $v_{j}$ are non-adjacent, so the configuration in Figure 2(b) is not possible for $v$. Moreover, $v$ cannot have the form in Figure 2(c) as well, since $vv_{i}$ and $vv_{j}$ are special edges. Hence, we deduce that the neighbours of $v$ can only form as depicted in Figure 2(a), where $v_{i}=v_{2}$ and $v_{j}=v_{5}$. By Lemmas 2.2 and 2.12, $v_{2}$ and $v_{5}$ cannot have a $4^{-}$-neighbour, bad $5$-neighbour and 6(6)666(6)6 ( 6 )-neighbour, so the other neighbours of $v$ are $6(5^{-})$-vertices or $5(t)$-vertices for $0\leq t\leq 3$. Thus, $v$ gives $\frac{1}{6}$ to each of $v_{2},v_{5}$ by R4(c), and at most $\frac{1}{12}$ to each of the other neighbours of $v$ by R4(a). Consequently, $\mu^{*}(v)\geq 2-2\times\frac{1}{6}-4\times\frac{1}{12}-4\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
Now we suppose that $v$ has exactly one 5(5)555(5)5 ( 5 )-neighbour $v_{i}$ such that $vv_{i}$ is a special edge. Then, similarly as above, we deduce that the neighbours of $v$ can only form as depicted in Figure 2(a) where $v_{i}=v_{2}$ (or $v_{i}=v_{5}$). By Lemmas 2.2 and 2.12, $v_{2}$ cannot have a $4^{-}$-neighbour, bad $5$-neighbour and 6(6)666(6)6 ( 6 )-neighbour, so the neighbours of $v_{2}$ other than $v$ are $6(5^{-})$-vertices or $5(t)$-vertices for $0\leq t\leq 3$. Thus, $v$ gives $\frac{1}{6}$ to $v_{2}$, at most $\frac{1}{12}$ to each of $v_{1},v_{3}$ by R4(a), at most $\frac{1}{9}$ to $v_{5}$ by R4(a),(b),(d), and at most $\frac{1}{12}$ to each of $v_{4},v_{6}$ by R4(a),(b),(d). Hence, $\mu^{*}(v)\geq 2-\frac{1}{6}-2\times\frac{1}{12}-\frac{1}{9}-2\times\frac{1}{12}-4\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
(4.2). Let $m_{3}(v)=5$. By Lemma 2.15, we have $n_{4}(v)\leq 3$.
Let $n_{4}(v)=3$. Note that $v$ has no bad $5$-neighbours by Lemma 2.15(a). So, $v$ only gives charge to its bad $4$-neighbours and its incident $3$-faces. Thus, $\mu^{*}(v)\geq 2-5\times\frac{1}{3}-3\times\frac{1}{12}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and $\frac{1}{12}$ to each of its bad $4$-neighbours by R5(a).
Let $n_{4}(v)=2$. By Lemma 2.15(b), $v$ has at most two bad $5$-neighbours. If $v$ has such two neighbours, then the remaining neighbours of $v$ are $6(4^{-})$-vertices by Lemma 2.15(c), i.e., $v$ has two $6(4^{-})$-neighbours. In such a case, $v$ gives $\frac{1}{12}$ to each of its bad $4$-neighbours by R5(a), and $\frac{1}{9}$ to each of its bad $5$-neighbours by R5(b). On the other hand, $v$ receives at least $\frac{1}{12}$ from each of its $6(4^{-})$-neighbours by R3 and R4(a). Thus, $\mu^{*}(v)\geq 2-2\times\frac{1}{12}-2\times\frac{1}{9}+2\times\frac{1}{12}-5\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1. Suppose now that $v$ has at most one bad $5$-neighbour. Then, $v$ gives $\frac{1}{12}$ to each of its bad $4$-neighbours by R5(a), and $\frac{1}{9}$ to its bad $5$-neighbour R5(b) (if exists). Thus, $\mu^{*}(v)\geq 2-2\times\frac{1}{12}-\frac{1}{9}-5\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
Let $n_{4}(v)=1$. By Lemma 2.15(d), $v$ has at most three bad $5$-neighbours. If $v$ has such three neighbours, then $v$ would have a $6(4^{-})$-neighbour by Lemma 2.15(e), and so $v$ receives at least $\frac{1}{12}$ from its $6(4^{-})$-neighbour by R3 and R4(a). On the other hand, $v$ gives $\frac{1}{12}$ to its bad $4$-neighbour by R5(a), and $\frac{1}{9}$ to each of its bad $5$-neighbours by R5(b). Thus, $\mu^{*}(v)\geq 2+\frac{1}{12}-\frac{1}{12}-3\times\frac{1}{9}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1. Suppose now that $v$ has at most two bad $5$-vertices. Then, $\mu^{*}(v)\geq 2-\frac{1}{12}-2\times\frac{1}{9}-5\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{12}$ to its bad $4$-neighbour R5(a), $\frac{1}{9}$ to each of its bad $5$-neighbours R5(b), and $\frac{1}{3}$ to each of its incident $3$-faces by R1.
Let $n_{4}(v)=0$. Notice first that $v$ has at most four bad $5$-neighbours by Lemma 2.15(f). If $v$ has at most three bad $5$-neighbours, then $v$ gives $\frac{1}{9}$ to each of its bad $5$-neighbours by R5(b), and so $\mu^{*}(v)\geq 2-3\times\frac{1}{9}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1. Assume further that $v$ has exactly four bad $5$-neighbours. If $m_{4}(v)=0$, i.e, $v$ is incident to a $5^{+}$-face, then $v$ receives $\frac{1}{5}$ from its incident $5^{+}$-face by R2(c). Thus, $\mu^{*}(v)\geq 2+\frac{1}{5}-4\times\frac{1}{9}-5\times\frac{1}{3}>0$ after $v$ gives $\frac{1}{9}$ to each of its bad $5$-neighbours by R5(b), and $\frac{1}{3}$ to each of its incident $3$-faces by R1. If $m_{4}(v)=1$, then $v$ has two $6(4^{-})$-neighbours by Lemma 2.15(g), and so $v$ receives at least $\frac{1}{12}$ from each of its $6(4^{-})$-neighbours by R3 and R4(a). Thus, $\mu^{*}(v)\geq 2+2\times\frac{1}{12}-4\times\frac{1}{9}-5\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{9}$ to each of its bad $5$-neighbours by R5(b), and $\frac{1}{3}$ to each of its incident $3$-faces by R1.
2.2. The case $\Delta=7$
Recall that $G$ does not admit any $2$-distance $21$-coloring, but any planar graph $G^{\prime}$ obtained from $G$ with smaller $|V(G^{\prime})|+|E(G^{\prime})|$ admits a 2-distance $21$-coloring. We begin with some structural properties of $G$ similar to the case when $\Delta=6$.
The proof of Lemma 2.16 is omitted since it is similar to Lemma 2.1.
Lemma 2.16.
We have $\delta(G)\geq 3$.
Lemma 2.17.
If $v$ is a $3$-vertex, then $m_{3}(v)=0$, $m_{4}(v)\leq 1$ and each neighbour of $v$ is a $7(5^{-})$-vertex.
Proof.
Let $v$ be a $3$-vertex. Assume for a contradiction that $v$ is incident to a $3$-face, say $f_{1}=v_{1}vv_{2}$. It is clear that $d_{2}(v)\leq 19$. If we set $G^{\prime}=G-v+\{v_{1}v_{3}\}$, then $G^{\prime}$ would be proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring with $21$ colors. Since $d_{2}(v)\leq 19$, we can color $v$ with an available color, a contradiction.
Suppose now that $m_{4}(v)\geq 2$, i.e., $v$ is incident to two $4$-faces, say $f_{1}=v_{1}vv_{2}x$ and $f_{2}=v_{2}vv_{3}y$. Clearly, we have $d_{2}(v)\leq 19$. If we set $G^{\prime}=G-v+\{v_{1}v_{3}\}$, then $G^{\prime}$ would be proper with respect to $G$. Similarly as above, we get a contradiction.
Next we claim that each neighbour of $v$ is a $7(5^{-})$-vertex. Note that $v$ cannot have any $7(6^{+})$-neighbour since $m_{3}(v)=0$. Moreover, $v$ cannot have any $6^{-}$-neighbour. Indeed, if $v_{1}$ is such a vertex, then we set $G^{\prime}=G-v+\{v_{1}v_{2},v_{1}v_{3}\}$. In such a case, $d_{2}(v)\leq 20$, and $G^{\prime}$ is proper with respect to $G$. Similarly as above, we get a contradiction. Thus, each neighbour of $v$ is a $7(5^{-})$-vertex. ∎
Lemma 2.18.
If $v$ is a $4$-vertex, then $m_{3}(v)\leq 3$, i.e., $G$ has no 4(4)444(4)4 ( 4 )-vertex.
Proof.
Let $v$ be a $4$-vertex. Assume that $m_{3}(v)=4$, and so we have $d_{2}(v)\leq 20$. If we set $G^{\prime}=G-v$, then $G^{\prime}$ is proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring $f$ with $21$ colors. Since $d_{2}(v)\leq 20$, we can color $v$ with an available color, a contradiction. ∎
Let us define some vertices having special forms: If $v$ is a $4(1)$-, $4(2)$- or $4(3)$-vertices, then we call $v$ as a bad $4$-vertex. Similarly, if $v$ is a $5(4)$- or 5(5)555(5)5 ( 5 )-vertices, then we call $v$ as a bad $5$-vertex.
Lemma 2.19.
Let $v$ be a $4$-vertex.
  • $(a)$
    $v$ has no 5(5)555(5)5 ( 5 )-neighbour.
  • $(b)$
    If $m_{3}(v)\geq 1$, then $v$ has at most one $4$-neighbour.
  • $(c)$
    If $m_{3}(v)\geq 1$, and $v$ has a $4$-neighbour, then $v$ cannot have any $5(4)$-neighbour.
Proof.
$(a)$ By contradiction, assume that $v$ has a 5(5)555(5)5 ( 5 )-neighbour. In such a case, $v$ should be either $4(2)$-vertex or $4(3)$-vertex, so we may assume that $vv_{2}$ is contained in two $3$-faces. Observe that $d_{2}(v)\leq 20$. If we set $G^{\prime}=G-v+\{v_{2}v_{4}\}$, then $G^{\prime}$ is proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring $f$ with $21$ colors. Since $d_{2}(v)\leq 20$, we can color $v$ with an available color, a contradiction.
$(b)$ Let $m_{3}(v)\geq 1$. Suppose that $v$ has two $4$-neighbours, and let $v_{1}$ be one of those $4$-vertices. In this case, we have $d_{2}(v)\leq 20$. If we set $G^{\prime}=G-v+\{v_{1}v_{2},v_{1}v_{3},v_{1}v_{4}\}$, then $G^{\prime}$ is proper with respect to $G$. In a similar way as above, we get a contradiction.
$(c)$ Let $m_{3}(v)\geq 1$. Assume that $v$ has both a $4$-neighbour and a $5(4)$-neighbour. In this case, we have $d_{2}(v)\leq 20$. Similarly as above, we get a contradiction. ∎
Lemma 2.20.
Let $v$ be a $4(1)$-vertex.
  • $(a)$
    If $m_{4}(v)=2$, then $v$ has a $7$-neighbour. In particular, $v$ has either a $7(5^{-})$-neighbour or two $7(6)$-neighbours.
  • $(b)$
    If $m_{4}(v)=3$, then $v$ has two $7(6^{-})$-neighbours.
Proof.
Let $v$ be a $4(1)$-vertex, and denote by $f_{1},f_{2},\ldots,f_{4}$ the faces incident to $v$ such that $f_{1}=v_{1}vv_{2}$.
$(a)$ Let $m_{4}(v)=2$. We first show that $v$ has a $7$-neighbour. Assume for a contradiction that all neighbours of $v$ are $6^{-}$-vertices. In such a case, we have $d_{2}(v)\leq 20$ since $m_{4}(v)=2$. If we set $G^{\prime}=G-v+\{v_{2}v_{3},v_{3}v_{4},v_{1}v_{4}\}$, then $G^{\prime}$ is proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring $f$ with $21$ colors. Since $d_{2}(v)\leq 20$, we can color $v$ with an available color, a contradiction.
Now we assume that $v$ has no $7(5^{-})$-neighbour. Since $v$ has no any 7(7)777(7)7 ( 7 )-neighbour as $m_{3}(v)=1$, the vertex $v$ is incident to a $7(6)$-vertex, say $v_{i}$. Clearly we have $i=1$ or $i=2$, say $i=1$. If the other neighbours of $v$ are $6^{-}$-vertices, then we have $d_{2}(v)\leq 20$. When we set $G^{\prime}=G-v+\{v_{2}v_{3},v_{3}v_{4},v_{1}v_{4}\}$, the graph $G^{\prime}$ is proper with respect to $G$. Similarly as above, we get a contradiction. This infers that $v$ has one more $7(6)$-neighbour.
$(b)$ Let $m_{4}(v)=3$. We first show that $v$ has two $7$-neighbours. Assume for a contradiction that three neighbours of $v$ are $6^{-}$-vertices. In such a case, we have $d_{2}(v)\leq 20$ since $m_{4}(v)=3$. If we set $G^{\prime}=G-v+\{v_{2}v_{3},v_{1}v_{4}\}$, then $G^{\prime}$ is proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring $f$ with $21$ colors. Since $d_{2}(v)\leq 20$, we can color $v$ with an available color, a contradiction. It then follows that $v$ has two $7(6^{-})$-neighbours, since $v$ cannot have any 7(7)777(7)7 ( 7 )-neighbour. ∎
Lemma 2.21.
Let $v$ be a $4(2)$-vertex.
  • $(a)$
    If $m_{4}(v)=0$, then $v$ has either a $7(5^{-})$-neighbour and a $6(4^{-})$-neighbour or two $7(6^{-})$-neighbours.
  • $(b)$
    If $m_{4}(v)=1$, then $v$ has either two $7(6^{-})$-neighbours and two $6(4^{-})$-neighbours or three $7(6^{-})$-neighbours.
  • $(c)$
    If $m_{4}(v)=2$, then $v$ has either two $7(5^{-})$-neighbours and two $7(6)$-neighbours or three $7(5^{-})$-neighbours.
Proof.
Let $v$ be a $4(2)$-vertex, and denote by $f_{1},f_{2},\ldots,f_{4}$ the faces incident to $v$ such that $f_{i}=v_{i}vv_{i+1}$ and $f_{j}=v_{j}vv_{j+1}$ with $i<j$ are $3$-faces (see Figure 3).
We remark that if $d_{2}(v)\leq 20$, then we set $G^{\prime}=G-v+\{v_{2}v_{4}\}$ (when $i=1$, $j=2$) or $G^{\prime}=G-v+\{v_{2}v_{3},v_{1}v_{4}\}$ (when $i=1$, $j=3$), and so the graph $G^{\prime}$ remains proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring $f$ with $21$ colors. Since $d_{2}(v)\leq 20$, we can color $v$ with an available color, a contradiction. We therefore assume that $d_{2}(v)\geq 21$.
$(a)$ Let $m_{4}(v)=0$. Suppose first that $v$ has a 7(7)777(7)7 ( 7 )-neighbour. If $v$ has two $6^{-}$-neighbours, then $d_{2}(v)\leq 20$, a contradiction. This infers that $v$ has two $7(6^{-})$-neighbours. Now we assume that $v$ has no 7(7)777(7)7 ( 7 )-neighbour. Clearly, $v$ has at least one $7$-neighbour, since otherwise, we have $d_{2}(v)\leq 20$ when all neighbours of $v$ are $6^{-}$-vertices, a contradiction. So, $v$ has a $7(6^{-})$-neighbour, say $v_{r}$. Assume that $v$ has exactly one $7(6^{-})$-neighbour, since otherwise, the claim holds. So, the other neighbours of $v$ are $6^{-}$-vertices. If $v_{r}$ is a $7(6)$-vertex or $v$ has a $6(5^{+})$-neighbour, then we have again $d_{2}(v)\leq 20$, a contradiction. Thus, we conclude that $v_{r}$ is a $7(5^{-})$-vertex, as well as $v$ has a $6(4^{-})$-neighbour.
$v$$v_{2}$$v_{3}$$v_{4}$$v_{1}$
(a) $i=1$ and $j=2$
$v$$v_{2}$$v_{3}$$v_{4}$$v_{1}$
(b) $i=1$ and $j=3$
Figure 3. Two possible $3$-faces $f_{i}=v_{i}vv_{i+1}$ and $f_{j}=v_{j}vv_{j+1}$ incident to a $4(2)$-vertex.
$(b)$ Let $m_{4}(v)=1$. We first claim that $v$ has at least two $7$-neighbours. Indeed, if three neighbours of $v$ are $6^{-}$-vertices, then we have $d_{2}(v)\leq 20$ as $m_{4}(v)=2$, a contradiction. Notice that $v$ has at most one 7(7)777(7)7 ( 7 )-neighbour as $m_{3}(v)=2$. We distinguish the remainder of the proof based on the number of edges $v_{i}v_{i+1}$, $v_{j}v_{j+1}$ contained in two $3$-faces.
Suppose that both $v_{i}v_{i+1}$ and $v_{j}v_{j+1}$ are contained in two $3$-faces. In this case, $v$ cannot have a $6^{-}$-neighbour, since otherwise, we have $d_{2}(v)\leq 20$, a contradiction. This infers that $v$ has three $7(6^{-})$-neighbours.
Next we suppose that exactly one of the edges $v_{i}v_{i+1}$, $v_{j}v_{j+1}$ is contained in two $3$-faces. In this case, $v$ cannot have any 7(7)777(7)7 ( 7 )- and 6(6)666(6)6 ( 6 )-neighbours, i.e., $v$ has two $7(6^{-})$-neighbours, say $v_{1},v_{2}$. Consider the vertices $v_{3},v_{4}$, if one of them is a $5^{-}$-vertex or both are $6$-vertices, then we have again $d_{2}(v)\leq 20$, a contradiction. Thus, one of $v_{3},v_{4}$ is a $7(6^{-})$-vertex while the other is a $6^{+}$-vertex. This implies that $v$ has one more $7(6^{-})$-neighbour, i.e., $v$ has totally at least three $7(6^{-})$-neighbours.
Suppose now that none of the edges $v_{i}v_{i+1}$, $v_{j}v_{j+1}$ is contained in two $3$-faces. In this way, $v$ cannot have any 7(7)777(7)7 ( 7 )-, $7(6)$-, 6(6)666(6)6 ( 6 )- and $6(5)$-neighbours. So, $v$ has two $7(5^{-})$-neighbours, say $v_{1},v_{2}$. Consider the vertices $v_{3},v_{4}$, if one of them is a $5^{-}$-vertex, then the other must be $7(5^{-})$-vertex, since otherwise we have $d_{2}(v)\leq 20$, a contradiction. If none of $v_{3},v_{4}$ is a $5^{-}$-vertex, then we deduce that either one of $v_{3},v_{4}$ is a $7(5^{-})$-vertex or both are $6(4^{-})$-vertices.
$(c)$ Let $m_{4}(v)=2$. Notice first that if $v$ has two $6^{-}$-neighbours or one 7(7)777(7)7 ( 7 )-neighbour or one 6(6)666(6)6 ( 6 )-neighbour, then we have $d_{2}(v)\leq 20$ since $m_{4}(v)=2$, a contradiction. Therefore, we deduce that $v$ has three $7(6^{-})$-neighbours. In fact, if $v$ has three $7(6)$-neighbours, then both $v_{i}v_{i+1}$ and $v_{j}v_{j+1}$ must be contained in two $3$-faces, and so we have again $d_{2}(v)\leq 20$, a contradiction. Hence, $v$ has two $7(6^{-})$-neighbours and one $7(5^{-})$-neighbours. Moreover, if $v$ has two $7(6)$-neighbours, the other neighbours of $v$ must be $7(5^{-})$-vertices, since otherwise, we have $d_{2}(v)\leq 20$, a contradiction. Consequently, the claim holds. ∎
Lemma 2.22.
Let $v$ be a $4(3)$-vertex.
  • $(a)$
    If $m_{4}(v)=0$, then $v$ has either four $7(5^{-})$-neighbours or three $7(5^{-})$-neighbours and a $7(6)$-neighbour or three $7(5^{-})$-neighbours and a $6(4^{-})$-neighbour or two $7(5^{-})$-neighbours and two $7(6)$-neighbours.
  • $(b)$
    If $m_{4}(v)=1$, then all neighbours of $v$ are $7(5^{-})$-vertices.
Proof.
Let $v$ be a $4(3)$-vertex, and denote by $f_{1},f_{2},\ldots,f_{4}$ the faces incident to $v$ such that $f_{i}=v_{i}vv_{i+1}$ for each $i\in[3]$ is a $3$-face.
We first claim that at most one edge in $G[N(v)]$ is contained in two $3$-faces. Indeed, if there exist such two edges, then we would have $d_{2}(v)\leq 20$. When we set $G^{\prime}=G-v+\{v_{1}v_{4}\}$, the graph $G^{\prime}$ remains proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring $f$ with $21$ colors. Since $d_{2}(v)\leq 20$, we can color $v$ with an available color, a contradiction. This particularly infers that $v$ has neither 6(6)666(6)6 ( 6 )- nor 7(7)777(7)7 ( 7 )-neighbour. In a similar way, we deduce that $v$ cannot have any $6(5)$-neighbour, two $6(4^{-})$-neighbours, and three $7(6)$-neighbours.
$(a)$ Let $m_{4}(v)=0$. As earlier stated, $v$ has at most one $6(4^{-})$-neighbours. In fact, if $v$ has a $6(4^{-})$-neighbour, then $v$ cannot have any $7(6)$-neighbours, since otherwise, we have $d_{2}(v)\leq 20$, and similarly as above, we get a contradiction. This implies that the neighbours of $v$ other than $6(4^{-})$-vertex are $7(5^{-})$-vertices. Suppose now that $v$ has no $6(4^{-})$-neighbour. Since $v$ has at most two $7(6)$-neighbours, we have either four $7(5^{-})$-neighbours or three $7(5^{-})$-neighbours and a $7(6)$-neighbour or two $7(5^{-})$-neighbours and two $7(6)$-neighbours.
$(b)$ Let $m_{4}(v)=1$. If $v$ has a $7(6^{+})$- or $6^{-}$-neighbour, then $d_{2}(v)\leq 20$, and similarly as above, we get a contradiction. ∎
Lemma 2.23.
Let $v$ be a $5(4)$-vertex.
  • $(a)$
    $v$ cannot have two 5(5)555(5)5 ( 5 )-neighbours.
  • $(b)$
    If $v$ has a 5(5)555(5)5 ( 5 )-neighbour, then $v$ cannot be adjacent to a $4$-vertex.
  • $(c)$
    If $v$ has a $7(6)$-neighbour, then $v$ cannot be adjacent to two $4$-vertices.
  • $(d)$
    If $m_{4}(v)=0$, then $v$ has two neighbours different from $5^{-}$- and 6(6)666(6)6 ( 6 )-vertex.
  • $(e)$
    If $m_{4}(v)=1$, then $v$ has at least three $6^{+}$-neighbours different from 6(6)666(6)6 ( 6 )-vertex.
  • $(f)$
    If $m_{4}(v)=1$, and $v$ has a 7(7)777(7)7 ( 7 )-neighbour, then $v$ has at least four $6^{+}$-neighbours different from 6(6)666(6)6 ( 6 )-vertex.
Proof.
Let $v$ be a $5(4)$-vertex, and denote by $f_{1},f_{2},\ldots,f_{5}$ the faces incident to $v$ such that $f_{i}=v_{i}vv_{i+1}$ for $i\in[4]$. Remark that if $d_{2}(v)\leq 20$, then we set $G^{\prime}=G-v+\{v_{4}v_{5}\}$, and so the graph $G^{\prime}$ remains proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring $f$ with $21$ colors. Since $d_{2}(v)\leq 20$, we can color $v$ with an available color, a contradiction. We therefore assume that $d_{2}(v)\geq 21$.
$(a)$ If $v$ has two 5(5)555(5)5 ( 5 )-neighbours, then there exist at least three edges in $G[N(v)]$ such that each of them is contained in two $3$-faces, and so we have $d_{2}(v)\leq 20$, a contradiction.
$(b)$ Suppose that $v$ has both a 5(5)555(5)5 ( 5 )-neighbour and a $4$-neighbour. Then we have again $d_{2}(v)\leq 20$, a contradiction.
$(c)$ Assume that $v$ has a $7(6)$-neighbour as well as $v$ is adjacent to two $4$-vertices. Since $v$ is adjacent to a $7(6)$-vertex, at least one edge in $G[N(v)]$ is contained in two $3$-faces. Thus, we have $d_{2}(v)\leq 20$, a contradiction.
$(d)$ Let $m_{4}(v)=0$. By contradiction, assume that $v$ has at most one neighbour consisting of $6(5^{-})$- or $7$-vertices, i.e., all neighbours of $v$ but one consist of 6(6)666(6)6 ( 6 )- or $5^{-}$-vertices. Notice that if $v$ has a 6(6)666(6)6 ( 6 )-neighbour, then there exist two edges in $G[N(v)]$ such that each of them is contained in two $3$-faces. It can be easily observed that $v$ has no four 6(6)666(6)6 ( 6 )-neighbours. Consider all the other possibilities; $v$ has either three 6(6)666(6)6 ( 6 )-neighbours or two 6(6)666(6)6 ( 6 )-neighbours and one $5^{-}$-neighbour or one 6(6)666(6)6 ( 6 )-neighbour and two $5^{-}$-neighbours or four $5^{-}$-neighbours. In each case, we have $d_{2}(v)\leq 20$, a contradiction. Thus, $v$ has two neighbours different from $5^{-}$- and 6(6)666(6)6 ( 6 )-vertex.
$(e)$ Let $m_{4}(v)=1$. Assume that $v$ has three neighbours $x,y,z$ consisting of $5^{-}$- or 6(6)666(6)6 ( 6 )-vertices. Recall that if $v$ has a 6(6)666(6)6 ( 6 )-neighbour, then there exist two edges in $G[N(v)]$ such that each of them is contained in two $3$-faces. Consider the possibilities; all $x,y,z$ are $5^{-}$-vertices; two of $x,y,z$ are $5^{-}$-vertices and one is 6(6)666(6)6 ( 6 )-vertex; one of $x,y,z$ is a $5^{-}$-vertex and the others are 6(6)666(6)6 ( 6 )-vertices; all $x,y,z$ are 6(6)666(6)6 ( 6 )-vertices. In each case, we observe that $d_{2}(v)\leq 20$, a contradiction.
$(f)$ Let $m_{4}(v)=1$. Suppose that $v$ has a 7(7)777(7)7 ( 7 )-neighbour. This implies that there exist two edges in $G[N(v)]$ such that each of them is contained in two $3$-faces. In such a case, if $v$ has two neighbours consisting of $5^{-}$- or 6(6)666(6)6 ( 6 )-vertices, then we would have $d_{2}(v)\leq 20$, a contradiction. That is, $v$ has at least four $6^{+}$-neighbours different from 6(6)666(6)6 ( 6 )-vertex. ∎
Lemma 2.24.
Let $v$ be a 5(5)555(5)5 ( 5 )-vertex.
  • $(a)$
    $n_{4}(v)=0$ and $n_{5}(v)\leq 2$.
  • $(b)$
    $v$ cannot have two neighbours consisting of $5(4)$-, 5(5)555(5)5 ( 5 )- or 6(6)666(6)6 ( 6 )-vertices. In particular, $v$ has at least one $7$-neighbour.
  • $(c)$
    If $v$ has four $6$-neighbours, then $v$ has also a $7(5^{-})$-neighbour.
  • $(d)$
    If $n_{5}(v)=0$, and $v$ has both a 6(6)666(6)6 ( 6 )-neighbour and a $6(5^{-})$-neighbour, then $v$ has three $7$-neighbours; at least one of which is a $7(5^{-})$-vertex.
  • $(e)$
    If $n_{5}(v)=1$, then $v$ has no 6(6)666(6)6 ( 6 )-neighbours, in particular, $v$ has at most two $6$-neighbours.
  • $(f)$
    If $n_{5}(v)=1$, and $v$ has exactly one $6(5^{-})$-neighbour, then $v$ has a $7(5^{-})$-neighbour.
  • $(g)$
    If $n_{5}(v)=1$, and $v$ has exactly two $6(5^{-})$-neighbours, then $v$ has two $7(5^{-})$-neighbours.
  • $(h)$
    If $n_{5}(v)=2$, then $v$ has three $7(5^{-})$-neighbours.
Proof.
Let $v$ be a 5(5)555(5)5 ( 5 )-vertex, and denote by $f_{1},f_{2},\ldots,f_{5}$ the $3$-faces incident to $v$ such that $f_{i}=v_{i}vv_{i+1}$ for $i\in[5]$ in a cyclic fashion. Remark that if $d_{2}(v)\leq 20$, then we set $G^{\prime}=G-v$, and so the graph $G^{\prime}$ remains proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring $f$ with $21$ colors. Since $d_{2}(v)\leq 20$, we can color $v$ with an available color, a contradiction. We therefore assume that $d_{2}(v)\geq 21$.
$(a)$ By Lemma 2.19(a), $v$ has no $4^{-}$-neighbours. Moreover, if $v$ has three $5$-neighbours, then we would have $d_{2}(v)\leq 19$, a contradiction.
$(b)$ If $v$ has two neighbours consisting of $5(4)$-, 5(5)555(5)5 ( 5 )- or 6(6)666(6)6 ( 6 )-vertices, then $d_{2}(v)\leq 20$, a contradiction. Besides, if all neighbours of $v$ are $6^{-}$-vertices, then we would have again $d_{2}(v)\leq 20$, a contradiction.
$(c)$ Suppose that $v$ has four $6$-neighbours. By $(b)$, the last neighbour of $v$ must be $7$-vertex, say $v_{1}$. If $v_{1}$ is a $7(6^{+})$-vertex, then there exists at least one edge in $G[N(v)]$ such that it is contained in two $3$-faces, and so we have $d_{2}(v)\leq 20$, a contradiction. Thus, $v_{1}$ is a $7(5^{-})$-vertex.
$(d)$ Let $n_{5}(v)=0$, and suppose that $v$ has both a 6(6)666(6)6 ( 6 )-neighbour and a $6(5^{-})$-neighbour. If $v$ has at most two $7$-neighbours, then we would have $d_{2}(v)\leq 20$, a contradiction. Moreover, at least one of those three $7$-neighbours must be $7(5^{-})$, since otherwise there exist at least three edges in $G[N(v)]$ such that each of them is contained in two $3$-faces, and so we have $d_{2}(v)\leq 20$, a contradiction.
$(e)$ Let $n_{5}(v)=1$. If $v$ has a 6(6)666(6)6 ( 6 )-neighbour or three $6$-neighbours, then we have $d_{2}(v)\leq 20$, a contradiction.
$(f)$ Let $n_{5}(v)=1$, and suppose that $v$ has exactly one $6(5^{-})$-neighbour. It follows from $(e)$ that $v$ has no 6(6)666(6)6 ( 6 )-neighbours, i.e., three neighbours of $v$ are $7$-vertices, where we recall that $v$ has no $4$-neighbour by $(a)$. If all those $7$-vertices are $7(6^{+})$-vertices, then there exist at least two edges in $G[N(v)]$ such that each of them is contained in two $3$-faces, and so we have $d_{2}(v)\leq 20$, a contradiction.
$(g)$ Let $n_{5}(v)=1$, and suppose that $v$ has exactly two $6(5^{-})$-neighbour, i.e., $v$ has two $7$-neighbours. If one of those $7$-vertices is a $7(6^{+})$-vertex, then there exists at least one edge in $G[N(v)]$ such that it is contained in two $3$-faces, and so we have $d_{2}(v)\leq 20$, a contradiction. Thus, $v$ has two $7(5^{-})$-neighbours.
$(h)$ Let $n_{5}(v)=2$. If $v$ has a $6$- or $7(6^{+})$-neighbours, then we have $d_{2}(v)\leq 20$, a contradiction. ∎
Lemma 2.25.
Let $v$ be a $6(5)$-vertex. Then, $v$ has at most four bad $5$-neighbours. In particular, if $m_{4}(v)=1$ and $v$ has four bad $5$-neighbour, then $v$ has two $7(5^{-})$-neighbours.
Proof.
Let $v$ be a $6(5)$-vertex, and denote by $f_{1},f_{2},\ldots,f_{6}$ the faces incident to $v$ such that $f_{i}=v_{i}vv_{i+1}$ for $i\in[5]$.
Assume for a contradiction that $v$ has five bad $5$-neighbours. This implies that there exist at least three edges in $G[N(v)]$ such that each of them is contained in two $3$-faces, and so we have $d_{2}(v)\leq 19$. Since $v$ has five bad $5$-neighbours, one of $v_{1},v_{6}$ is a $5$-vertex, say $v_{1}$. Also, one of $v_{3},v_{5}$ is a $5$-vertex, say $v_{3}$. Let $G^{\prime}=G-v+\{v_{1}v_{6},v_{3}v_{1},v_{3}v_{5}\}$. Clearly, $G^{\prime}$ is proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring $f$ with $21$ colors. Since $d_{2}(v)\leq 19$, we can color $v$ with an available color, a contradiction.
Now we suppose that $m_{4}(v)=1$, and $v$ has four bad $5$-neighbours. Since $v$ has four $5(4^{+})$-neighbours, there exist at least two edges in $G[N(v)]$ such that each of them is contained in two $3$-faces. In this manner, if $v$ has at most one $7(5^{-})$-neighbour, then we would have $d_{2}(v)\leq 20$. We get a contradiction as above. ∎
Let $v$ be a $7$-vertex and let $x$ be a $4$-neighbour of $v$. If $v$ and $x$ have two common neighbours, i.e., the edge $vx$ is contained in two $3$-faces, then $x$ is called a support vertex of $v$.
A vertex $v$ is said to be poor if it is a $4$- or 5(5)555(5)5 ( 5 )-vertex. Recall that $G$ has no 4(4)444(4)4 ( 4 )-vertices, and so poor vertices are $4(0)$-, $4(1)$-, $4(2)$-, $4(3)$- and 5(5)555(5)5 ( 5 )-vertices.
Lemma 2.26.
Let $v$ be a $7$-vertex.
  • $(a)$
    If $v$ has a support vertex, then $d_{2}(v)\geq 21$.
  • $(b)$
    If $4\leq m_{3}(v)\leq 5$, then $v$ has at most six $4$-neighbours.
  • $(c)$
    If $m_{3}(v)=5$ and $n_{4}(v)\leq 2$, then $v$ has at most four poor neighbours.
  • $(d)$
    If $v$ is a $7(6)$-vertex, then $v$ has at most five poor neighbours.
  • $(e)$
    If $n_{4}(v)=3$, then $v$ has at most two 5(5)555(5)5 ( 5 )-neighbours.
  • $(f)$
    If $n_{4}(v)=4$, then $v$ has at most one 5(5)555(5)5 ( 5 )-neighbours.
  • $(g)$
    If $m_{3}(v)=5$, $n_{4}(v)=4$, $v$ has a support vertex, and $v$ has both a $5(4)$- and a 5(5)555(5)5 ( 5 )-neighbours, then $v$ is adjacent to a $7$-vertex.
Proof.
Let $v$ be a $7$-vertex, and denote by $f_{1},f_{2},\ldots,f_{7}$ the faces incident to $v$.
$(a)$ Suppose that $v_{1}$ is a support vertex of $v$. By definition, $v_{1}$ is a $4$-vertex and it is adjacent to both $v_{2}$ and $v_{7}$. Assume for a contradiction that $d_{2}(v)\leq 20$. When we set $G^{\prime}=G-v+\{v_{1}v_{3},v_{1}v_{4},v_{1}v_{5},v_{1}v_{6}\}$, the graph $G^{\prime}$ remains proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring $f$ with $21$ colors. Since $d_{2}(v)\leq 20$, we can color $v$ with an available color, a contradiction.
$(b)$ Let $4\leq m_{3}(v)\leq 5$. Assume for a contradiction that all neighbours of $v$ are $4$-vertices, and so we have $d_{2}(v)\geq 20$. In addition, at least one neighbour of $v$ must be a support vertex due to $m_{3}(v)\geq 4$ and $n_{4}(v)=7$. It then follows from $(a)$ that we get a contradiction.
$(c)$ By contradiction, assume that $v$ has five poor neighbours. Note that a 5(5)555(5)5 ( 5 )-vertex is adjacent to at most one 5(5)555(5)5 ( 5 )-vertex by Lemma 2.24(b). Therefore $v$ has at most three 5(5)555(5)5 ( 5 )-neighbours as $m_{3}(v)=5$. Since $n_{4}(v)\leq 2$, and $v$ has five poor neighbours, we then deduce that $v$ has exactly three 5(5)555(5)5 ( 5 )-neighbours and two $4$-neighbours. However, this is not possible since a 5(5)555(5)5 ( 5 )-vertex has no $4$-neighbour by Lemma 2.24(a).
$(d)$ Let $v$ be a $7(6)$-vertex. By Lemma 2.24(a),(b), a 5(5)555(5)5 ( 5 )-vertex has no $4$-neighbour, also a 5(5)555(5)5 ( 5 )-vertex is adjacent to at most one bad $5$-vertex. Moreover, a $4(1^{+})$-vertex is adjacent to at most one $4$-vertex by Lemma 2.19(b). Those facts imply that $v$ cannot have six poor neighbours.
$(e)$ Let $n_{4}(v)=3$. By Lemma 2.24(a),(b), a 5(5)555(5)5 ( 5 )-vertex has no $4$-neighbour, also a 5(5)555(5)5 ( 5 )-vertex is adjacent to at most one bad $5$-vertex. Thus, one can infer that $v$ has at most two 5(5)555(5)5 ( 5 )-neighbours.
$(f)$ Similarly as $(d)$, the claim holds.
$(g)$ Suppose that $m_{3}(v)=5$, $n_{4}(v)=4$, $v$ has a support vertex, and $v$ has a $5(4)$- and 5(5)555(5)5 ( 5 )-neighbour. If the last neighbour of $v$ is a $6^{-}$-vertex, then we have $d_{2}(v)\leq 20$, a contradiction by $(a)$. ∎
We next apply discharging to show that $G$ does not exist. We use the same initial charges as before, with the following discharging rules.
Discharging Rules
We apply the following discharging rules.
  • R1:
    Every $3$-face receives $\frac{1}{3}$ from each of its incident vertices.
  • R2:
    Every $5^{+}$-face gives $\frac{1}{5}$ to each of its incident vertices.
  • R3:
    Every $6(3^{-})$-vertex gives $\frac{1}{6}$ to each of its neighbours.
  • R4:
    Every $6(4)$-vertex gives $\frac{1}{9}$ to each of its neighbours.
  • R5:
    Every $6(5)$-vertex gives $\frac{1}{9}$ to each of its bad $5$-neighbours.
  • R6:
    Every $7(3^{-})$-vertex gives $\frac{2}{7}$ to each of its neighbours.
  • R7:
    Let $v$ be a $7(4)$- or $7(5)$-vertices. Then,
    • $(a)$
      $v$ gives $\frac{1}{5}$ to each of its $3$-neighbours,
    • $(b)$
      $v$ gives $\frac{1}{4}$ to each of its bad $4$-neighbours,
    • $(c)$
      $v$ gives $\frac{2}{9}$ to each of its 5(5)555(5)5 ( 5 )-neighbours,
    • $(d)$
      $v$ gives $\frac{1}{9}$ to each of its $5(4)$-neighbours,
    • $(e)$
      $v$ gives $\frac{1}{18}$ to each of its $6(5)$-neighbours.
  • R8:
    Let $v$ be a $7(6)$-vertex. Then,
    • $(a)$
      $v$ gives $\frac{1}{6}$ to each of its 5(5)555(5)5 ( 5 )-neighbours,
    • $(b)$
      $v$ gives $\frac{1}{9}$ to each of its $5(4)$-neighbours,
    • $(c)$
      $v$ gives $\frac{1}{6}$ to each of its bad $4$-neighbours.
  • R9:
    Let $v$ be a 7(7)777(7)7 ( 7 )-vertex. Then,
    • $(a)$
      $v$ gives $\frac{1}{6}$ to each of its 5(5)555(5)5 ( 5 )-neighbours,
    • $(b)$
      $v$ gives $\frac{1}{12}$ to each of its $5(4)$-neighbours.
Checking $\mu^{*}(v),\mu^{*}(f)\geq 0$ for $v\in V(G),f\in F(G)$.
First we show that $\mu^{*}(f)\geq 0$ for each $f\in F(G)$. Given a face $f\in F(G)$, recall that the initial charge of a face $f$ is $\mu(f)=\ell(f)-4$. If $f$ is a $3$-face, then it receives $\frac{1}{3}$ from each of its incident vertices by R1, and so $\mu^{*}(f)\geq-1+3\times\frac{1}{3}=0$. If $f$ is a $4$-face, then $\mu(f)=\mu^{*}(f)=0$. Let $f$ be a $5^{+}$-face. By applying R2, $f$ sends $\frac{1}{5}$ to each of its incident vertices. It then follows that $\mu^{*}(f)\geq 1-5\times\frac{1}{5}=0$. Consequently, $\mu^{*}(f)\geq 0$ for each $f\in F(G)$.
We now pick a vertex $v\in V(G)$ with $d(v)=k$. By Lemma 2.16, we have $k\geq 3$.
(1). Let $k=3$. The initial charge of $v$ is $\mu(v)=d(v)-4=-1$. By Lemma 2.17, $m_{3}(v)=0$, $m_{4}(v)\leq 1$ and each neighbour of $v$ is a $7(5^{-})$-vertex. This means that $v$ is incident to at least two $5^{+}$-faces. It then follows that $v$ receives $\frac{1}{5}$ from each of its incident $5^{+}$-faces by R2, and at least $\frac{1}{5}$ from each of its $7(5^{-})$-neighbours by R6 and R7(a). Hence, $\mu^{*}(v)\geq-1+2\times\frac{1}{5}+3\times\frac{1}{5}=0$.
(2). Let $k=4$. The initial charge of $v$ is $\mu(v)=d(v)-4=0$. We have $m_{3}(v)\leq 3$ by Lemma 2.18. If $m_{3}(v)=0$, then $\mu^{*}(v)\geq 0$ since $v$ does not give a charge to its any incident faces. So we may assume that $1\leq m_{3}(v)\leq 3$.
(2.1). Let $m_{3}(v)=1$. If $m_{4}(v)\leq 1$, then $v$ is incident to two $5^{+}$-faces, and by R2, $v$ receives $\frac{1}{5}$ from each of those $5^{+}$-faces. Thus, $\mu^{*}(v)\geq 2\times\frac{1}{5}-\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to its incident $3$-face by R1. If $m_{4}(v)=2$, then $v$ has either a $7(5^{-})$-neighbour or two $7(6)$-neighbours by Lemma 2.20(a). In both cases, $v$ receives at least $\frac{1}{4}=\min\{\frac{1}{4},2\times\frac{1}{6}\}$ from its $7$-neighbours by applying R6-R8. On the other hand, $v$ receives $\frac{1}{5}$ from its incident $5^{+}$-face by R2. Thus, $\mu^{*}(v)\geq\frac{1}{4}+\frac{1}{5}-\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to its incident $3$-face by R1. If $m_{4}(v)=3$, then $v$ has two $7(6^{-})$-neighbours by Lemma 2.20(b). By applying R6-R8, $v$ receives totally at least $\frac{1}{3}$ from its $7$-neighbours. Thus, $\mu^{*}(v)\geq\frac{1}{3}-\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to its incident $3$-face by R1.
(2.2). Let $m_{3}(v)=2$. Since $0\leq m_{4}(v)\leq 2$, we consider the following cases:
If $m_{4}(v)=0$, then $v$ has either a $7(5^{-})$-neighbour and a $6(4^{-})$-neighbour or two $7(6^{-})$-neighbours by Lemma 2.21(a). It follows that $v$ receives totally at least $\frac{1}{3}=\min\{\frac{1}{4}+\frac{1}{9},2\times\frac{1}{6}\}$ from its $6^{+}$-neighbours by R3-R4 and R6-R8. In addition, $v$ is incident to two $5^{+}$-faces, and by R2, $v$ receives $\frac{1}{5}$ from each of those $5^{+}$-faces. Thus, $\mu^{*}(v)\geq\frac{1}{3}+2\times\frac{1}{5}-2\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
If $m_{4}(v)=1$, then $v$ has either two $7(6^{-})$-neighbours and two $6(4^{-})$-neighbours or three $7(6^{-})$-neighbours by Lemma 2.21(b). In both cases, $v$ receives totally at least $\frac{1}{2}=\min\{2\times\frac{1}{6}+2\times\frac{1}{9},3\times\frac{1}{6}\}$ from its $6^{+}$-neighbours by R3-R4 and R6-R8. Also, $v$ receives $\frac{1}{5}$ from its incident $5^{+}$-face by R2. Thus, $\mu^{*}(v)\geq\frac{1}{2}+\frac{1}{5}-2\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
If $m_{4}(v)=2$, then $v$ has either two $7(5^{-})$-neighbours and two $7(6)$-neighbours or three $7(5^{-})$-neighbours by Lemma 2.21(c). In both cases, $v$ receives totally at least $\frac{3}{4}=\min\{2\times\frac{1}{4}+2\times\frac{1}{6},3\times\frac{1}{4}\}$ from its $7^{+}$-neighbours by R6-R8. Thus, $\mu^{*}(v)\geq\frac{3}{4}-2\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
(2.3). Let $m_{3}(v)=3$. Suppose first that $m_{4}(v)=0$. It then follows from Lemma 2.22(a) that then $v$ has either four $7(5^{-})$-neighbours or three $7(5^{-})$-neighbours and a $7(6)$-neighbour or three $7(5^{-})$-neighbours and a $6(4^{-})$-neighbour or two $7(5^{-})$-neighbours and two $7(6)$-neighbours. In each case, $v$ receives at least $\frac{5}{6}=\min\{4\times\frac{1}{4},3\times\frac{1}{4}+\frac{1}{6},3\times\frac{1}{4}+\frac{1}{9},2\times\frac{1}{4}+2\times\frac{1}{6}\}$ from its $6^{+}$-neighbours by R3-R8. In addition, $v$ receives $\frac{1}{5}$ from its incident $5^{+}$-face by R2. Thus, $\mu^{*}(v)\geq\frac{5}{6}+\frac{1}{5}-3\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1. Now, we suppose that $m_{4}(v)=1$. By Lemma 2.22(b), all neighbours of $v$ are $7(5^{-})$-vertices, and so $v$ receives totally $4\times\frac{1}{4}$ from its $7(5^{-})$-neighbours by R6-R7. Thus, $\mu^{*}(v)\geq 4\times\frac{1}{4}-3\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
(3). Let $k=5$. The initial charge of $v$ is $\mu(v)=d(v)-4=1$. We distinguish three cases according to the number of $3$-faces incident to $v$ as follows.
(3.1). Let $m_{3}(v)\leq 3$. Obviously, $\mu^{*}(v)\geq 1-3\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
(3.2). Let $m_{3}(v)=4$. Suppose first that $m_{4}(v)=0$. By Lemma 2.23(d), $v$ has two neighbours different from $5^{-}$- and 6(6)666(6)6 ( 6 )-vertices, i.e., $v$ has two neighbours consisting of $6(5^{-})$- or $7$-vertices. It follows that $v$ receives totally at least $\frac{1}{6}=\min\{2\times\frac{1}{9},\frac{1}{9}+\frac{1}{12},2\times\frac{1}{12}\}$ from its $6^{+}$-neighbours by R3-R9, and $\frac{1}{5}$ from its incident $5^{+}$-face by R2. Thus, $\mu^{*}(v)\geq 1+\frac{1}{6}+\frac{1}{5}-4\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
Next we suppose that $m_{4}(v)=1$. By Lemma 2.23(e), $v$ has at least three $6^{+}$-neighbours different from 6(6)666(6)6 ( 6 )-vertex. If $v$ has no 7(7)777(7)7 ( 7 )-neighbour, then $v$ receives totally at least $3\times\frac{1}{9}$ from its $6^{+}$-neighbours by R3-R8. If $v$ has a 7(7)777(7)7 ( 7 )-neighbours, then $v$ has four $6^{+}$-neighbours different from 6(6)666(6)6 ( 6 )-vertex by Lemma 2.23(f), and so $v$ receives totally at least $4\times\frac{1}{12}$ from its $6^{+}$-neighbours by R3-R9. Thus, $\mu^{*}(v)\geq 1+\frac{1}{3}-4\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
(3.3). Let $m_{3}(v)=5$. Note that $v$ has no $4^{-}$-neighbour, and $v$ has at most two $5$-neighbours by Lemma 2.24(a).
(3.3.1). Let $n_{5}(v)=0$. By Lemma 2.24(b), $v$ has at most one 6(6)666(6)6 ( 6 )-neighbour, i.e., $v$ has at least four $6^{+}$-neighbours different from 6(6)666(6)6 ( 6 )-vertex. First, suppose that $v$ has a 6(6)666(6)6 ( 6 )-neighbour. If $v$ has no $6(5^{-})$-neighbour, then all the neighbours of $v$ different from 6(6)666(6)6 ( 6 )-vertex are $7$-vertices, and by applying R6-R9, $v$ receives totally at least $4\times\frac{1}{6}$ from its $7$-neighbours. So, $\mu^{*}(v)\geq 1+4\times\frac{1}{6}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1. On the other hand, if $v$ has a $6(5^{-})$-neighbour, then $v$ has three $7$-neighbours, at least one of which is a $7(5^{-})$-vertex by Lemma 2.24(d). Then $v$ receives at least $\frac{1}{9}$ from its $6(5^{-})$-neighbour by R3-R5, at least $\frac{2}{9}$ from each of its $7(5^{-})$-neighbours by R6-R7, at least $\frac{1}{6}$ from each of its other $7$-neighbours by R6-R9. So, $v$ receives totally at least $\frac{1}{9}+\frac{2}{9}+2\times\frac{1}{6}$ from its $6^{+}$-neighbours. Thus, $\mu^{*}(v)\geq 1+\frac{2}{3}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
Suppose now that $v$ has no 6(6)666(6)6 ( 6 )-neighbour. That is, all neighbours of $v$ are $6^{+}$-vertices different from 6(6)666(6)6 ( 6 )-vertex. By Lemma 2.24(b), $v$ has at least one $7$-neighbour, say $x$. Note that if $v$ has exactly four $6$-neighbours, then $x$ must be a $7(5^{-})$-vertex by Lemma 2.24(c). This means that $v$ has either four $6$-neighbours and one $7(5^{-})$-neighbour or at most three $6$-neighbours and two $7$-neighbours. In each case, $v$ receives totally at least $\frac{2}{3}=\min\{4\times\frac{1}{9}+\frac{2}{9},3\times\frac{1}{9}+2\times\frac{1}{6}\}$ from its $6^{+}$-neighbours by R3-R9. Thus, $\mu^{*}(v)\geq 1+\frac{2}{3}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
(3.3.2). Let $n_{5}(v)=1$. Notice that $v$ has no 6(6)666(6)6 ( 6 )-neighbours by Lemma 2.24(e), i.e., all neighbours of $v$ but one are $7$- or $6(5^{-})$-vertices. Moreover, $v$ has at most two $6(5^{-})$-neighbours by Lemma 2.24(e). We distinguish three cases according to the number of $6(5^{-})$-vertices adjacent to $v$ as follows.
First suppose that $v$ has no $6(5^{-})$-neighbour. So, $v$ has four $7$-neighbours, and by applying R6-R9, $v$ receives totally at least $4\times\frac{1}{6}$ from its $7$-neighbours. So, $\mu^{*}(v)\geq 1+4\times\frac{1}{6}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
Suppose next that $v$ has exactly one $6(5^{-})$-neighbour. Then $v$ has three $7$-neighbours, one of which is a $7(5^{-})$-vertex by Lemma 2.24(f). Thus, $v$ receives totally at least $\frac{1}{9}+2\times\frac{1}{6}+\frac{2}{9}$ from its $6^{+}$-neighbours by R3-R9. Hence, $\mu^{*}(v)\geq 1+\frac{2}{3}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
Finally, suppose that $v$ has exactly two $6(5^{-})$-neighbours. Then $v$ has two $7(5^{-})$-neighbours by Lemma 2.24(g). So, $v$ receives totally at least $2\times\frac{1}{9}+2\times\frac{2}{9}$ from its $6^{+}$-neighbours by R3-R7. Hence, $\mu^{*}(v)\geq 1+\frac{2}{3}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
(3.3.3). Let $n_{5}(v)=2$. Then $v$ has three $7(5^{-})$-neighbours by Lemma 2.24(h), and so $v$ receives $\frac{2}{9}$ from each of its $7(5^{-})$-neighbours by R6-R7. Thus, $\mu^{*}(v)\geq 1+3\times\frac{2}{9}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
(4). Let $k=6$. The initial charge of $v$ is $\mu(v)=d(v)-4=2$. If $m_{3}(v)\leq 3$, then $\mu^{*}(v)\geq 2-3\times\frac{1}{3}-6\times\frac{1}{6}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and $\frac{1}{6}$ to each of its neighbours by R3. Similarly, if $m_{3}(v)=4$, then $\mu^{*}(v)\geq 2-4\times\frac{1}{3}-6\times\frac{1}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and $\frac{1}{9}$ to each of its neighbours by R4. On the other hand, if $m_{3}(v)=6$, then $\mu^{*}(v)\geq 2-6\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1. Therefore, we further assume that $m_{3}(v)=5$. By Lemma 2.25, $v$ has at most four bad $5$-neighbours. Notice that if $v$ has at most three bad $5$-neighbours, then $\mu^{*}(v)\geq 2-5\times\frac{1}{3}-3\times\frac{1}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and $\frac{1}{9}$ to each of its bad $5$-neighbours by R5. Next we assume that $v$ has exactly four bad $5$-neighbours. If $m_{4}(v)=0$, then $v$ receives $\frac{1}{5}$ from its incident $5^{+}$-face by R2, and so $\mu^{*}(v)\geq 2+\frac{1}{5}-5\times\frac{1}{3}-4\times\frac{1}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-face by R1, and $\frac{1}{9}$ to each of its bad $5$-neighbours by R5. If $m_{4}(v)=1$, then $v$ has two $7(5^{-})$-neighbours by Lemma 2.25, and so $v$ receives at least $\frac{1}{18}$ from each of its $7(5^{-})$-neighbours by R6 and R7(e). Thus, $\mu^{*}(v)\geq 2+2\times\frac{1}{18}-5\times\frac{1}{3}-4\times\frac{1}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-face by R1, and $\frac{1}{9}$ to each of its bad $5$-neighbours by R5.
(5). Let $k=7$. The initial charge of $v$ is $\mu(v)=d(v)-4=3$. Notice first that if $m_{3}(v)\leq 3$, then we have $\mu^{*}(v)\geq 3-3\times\frac{1}{3}-7\times\frac{2}{7}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and $\frac{2}{7}$ to each of its neighbours by R6. Therefore, we may assume that $4\leq m_{3}(v)\leq 7$.
(5.1). Let $m_{3}(v)=4$. By Lemma 2.26(b), $v$ has at most six $4$-neighbours. If $n_{4}(v)\leq 4$, then $v$ sends $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), and at most $\frac{2}{9}$ to each of its other neighbours by R7. Consequently, $\mu^{*}(v)\geq 3-4\times\frac{1}{4}-3\times\frac{2}{9}-4\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1. Suppose now that $5\leq n_{4}(v)\leq 6$. Clearly, $v$ has no 5(5)555(5)5 ( 5 )-neighbour, since a $4$ vertex is not adjacent to any 5(5)555(5)5 ( 5 )-vertex by Lemma 2.19(a). If $v$ has a $3$-neighbour $x$, then the edge $xv$ is contained in at least one $5^{+}$-face by Lemma 2.17. So, $v$ receives at least $\frac{1}{5}$ from its incident $5^{+}$-faces by R2. Thus $\mu^{*}(v)\geq 3+\frac{1}{5}-6\times\frac{1}{4}-\frac{1}{5}-4\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), at most $\frac{1}{5}$ to each of its neighbours other than $4$-vertex by R7, and $\frac{1}{3}$ to each of its incident $3$-faces by R1. On the other hand, if $v$ has no $3$-neighbour, then $\mu^{*}(v)\geq 3-6\times\frac{1}{4}-\frac{1}{7}-4\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), at most $\frac{1}{9}$ to each of its neighbours other than $4$-vertex by R7, and $\frac{1}{3}$ to each of its incident $3$-faces by R1.
(5.2). Let $m_{3}(v)=5$. Notice that $v$ has at most one $3$-neighbour, since a $3$-vertex does not incident to any $3$-faces by Lemma 2.17. First we assume that $v$ has a $3$-neighbour $x$. In such a case, all $3$-faces incident to $f$ must have a consecutive ordering. Moreover, the edge $xv$ is contained in at least one $5^{+}$-face by Lemma 2.17. So, $v$ receives at least $\frac{1}{5}$ from its incident $5^{+}$-faces by R2. Observe first that if $v$ has three $4$-neighbours, then at least one of them must be support vertex of $v$. In such a case, we have $d_{2}(v)\geq 21$ by Lemma 2.26(a). This implies that $v$ has at most four poor neighbours, and recall that a poor vertex receives at most $\frac{1}{4}$ from $v$ by R7(b),(c). It then follows that $\mu^{*}(v)\geq 3+\frac{1}{5}-5\times\frac{1}{3}-\frac{1}{5}-4\times\frac{1}{4}-2\times\frac{1}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{5}$ to its $3$-neighbour $x$ by R7(a), at most $\frac{1}{4}$ to each of its poor neighbours by R7(b),(c), and at most $\frac{1}{9}$ to each of its other neighbours by R7(d),(e). We now suppose that $v$ has at most two $4$-neighbours. It then follows from Lemma 2.26(c) that $v$ has at most four poor neighbours. Thus, similarly as above, we have $\mu^{*}(v)\geq 3+\frac{1}{5}-5\times\frac{1}{3}-\frac{1}{5}-2\times\frac{1}{4}-2\times\frac{2}{9}-2\times\frac{1}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{5}$ to its $3$-neighbour $x$ by R7(a), at most $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), $\frac{2}{9}$ to each of its 5(5)555(5)5 ( 5 )-neighbours by R7(c), and at most $\frac{1}{9}$ to each of its other neighbours by R7(d),(e).
We may further assume that $v$ has no $3$-neighbours. By Lemma 2.26(b), $v$ has at most six $4$-neighbours, i.e., $n_{4}(v)\leq 6$. On the other hand, recall that $m_{3}(v)=5$. Therefore, if $v$ has more than four $4$-neighbours, then there exists $v_{i}\in N(v)$ such that $vv_{i}$ is contained in two $3$-faces, i.e., $v_{i}$ is a support vertex. It follows from Lemma 2.26(a) that we have $d_{2}(v)\geq 21$ when $v$ has a support vertex.
(5.2.1). Let $n_{4}(v)\leq 2$. Note that $v$ has at most four poor neighbours by Lemma 2.26(c). Then $\mu^{*}(v)\geq 3-5\times\frac{1}{3}-2\times\frac{1}{4}-2\times\frac{2}{9}-3\times\frac{1}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), $\frac{2}{9}$ to each of its 5(5)555(5)5 ( 5 )-neighbours by R7(c), at most $\frac{1}{9}$ to each of its other neighbours by R7(d),(e).
(5.2.2). Let $n_{4}(v)=3$. Obviously, $v$ has at most two 5(5)555(5)5 ( 5 )-neighbours by Lemma 2.26(e). Suppose first that $v$ has exactly two 5(5)555(5)5 ( 5 )-neighbours $x,y$. Since a $4$-vertex has no 5(5)555(5)5 ( 5 )-neighbours by Lemma 2.19(a), we deduce that $x$ and $y$ are adjacent. On the other hand, a 5(5)555(5)5 ( 5 )-vertex has at most one bad $5$-neighbour by Lemma 2.24(b), i.e., a 5(5)555(5)5 ( 5 )-vertex cannot have both a 5(5)555(5)5 ( 5 )-neighbour and a $5(4)$-neighbour simultaneously. This implies that the neighbours of $v$ other than $4$- and 5(5)555(5)5 ( 5 )-vertices are $5(3^{-})$- or $6^{+}$-vertices. Thus we have $\mu^{*}(v)\geq 3-5\times\frac{1}{3}-3\times\frac{1}{4}-2\times\frac{2}{9}-2\times\frac{1}{18}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), $\frac{2}{9}$ to each of its 5(5)555(5)5 ( 5 )-neighbours by R7(c), at most $\frac{1}{18}$ to each of its other neighbours by R7(e). Suppose now that $v$ has at most one 5(5)555(5)5 ( 5 )-neighbour. In such a case, we have again $\mu^{*}(v)\geq 3-5\times\frac{1}{3}-3\times\frac{1}{4}-\frac{2}{9}-3\times\frac{1}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), $\frac{2}{9}$ to each of its 5(5)555(5)5 ( 5 )-neighbours by R7(c), at most $\frac{1}{9}$ to each of its other neighbours by R7(d)-(e).
(5.2.3). Let $n_{4}(v)=4$. By Lemma 2.26(f), $v$ has at most one 5(5)555(5)5 ( 5 )-neighbour. First, suppose that $v$ has no 5(5)555(5)5 ( 5 )-neighbours, then $\mu^{*}(v)\geq 3-5\times\frac{1}{3}-4\times\frac{1}{4}-3\times\frac{1}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), at most $\frac{1}{9}$ to each of its other neighbours by R7(d)-(e). We may further assume that $v$ has exactly one 5(5)555(5)5 ( 5 )-neighbour, say $v_{1}$. Denote by $v_{2}$ and $v_{7}$ the common neighbours of $v$ and $v_{1}$. Recall that each of $v_{2},v_{7}$ is a $5^{+}$-vertex by Lemmas 2.17 and 2.19(a). Moreover, for $i\in\{2,7\}$, if $v_{i}$ is adjacent to a $4$-vertex, then $v_{i}$ is different from $5(4)$-vertex by Lemma 2.23(b). Recall also that a $4(1^{+})$-vertex is adjacent to at most one $4$-vertex by Lemma 2.19(b).
Suppose first that $v$ has a support vertex $x$. If $v$ has no $5(4)$-neighbour, then $\mu^{*}(v)\geq 3-5\times\frac{1}{3}-4\times\frac{1}{4}-\frac{2}{9}-2\times\frac{1}{18}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), $\frac{2}{9}$ to its 5(5)555(5)5 ( 5 )-neighbour by R7(c), at most $\frac{1}{18}$ to each of its other neighbours by R7(e). If $v$ has a $5(4)$-neighbour, then $v$ has also a $7$-neighbour by Lemma 2.26(g). Thus, $\mu^{*}(v)\geq 3-5\times\frac{1}{3}-4\times\frac{1}{4}-\frac{2}{9}-\frac{1}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), $\frac{2}{9}$ to its 5(5)555(5)5 ( 5 )-neighbour by R7(c), $\frac{1}{9}$ to its $5(4)$-neighbour by R7(d).
Next, suppose that $v$ has no support vertex. This means that each $4$-neighbour of $v$ is incident to a $4^{+}$-face containing $v$. Obviously, $v$ has one of the two configurations depicted in Figure 4, where $v_{i},v_{j},v_{k},v_{\ell}$ are $4$-vertices. By Lemma 2.19(a), a $4$-vertex has no 5(5)555(5)5 ( 5 )-neighbours, so the configuration in Figure 4(a) is not possible for $v$. So the neighbours of $v$ can only form as depicted in Figure 4(b). We then infer that $v$ cannot have any $5(4)$-neighbour by Lemma 2.23(b). Thus $\mu^{*}(v)\geq 3-5\times\frac{1}{3}-4\times\frac{1}{4}-\frac{2}{9}-2\times\frac{1}{18}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), $\frac{2}{9}$ to its 5(5)555(5)5 ( 5 )-neighbour by R7(c), at most $\frac{1}{18}$ to each of its other neighbours by R7(e).
$v$$v_{j}$$v_{k}$$v_{\ell}$$v_{i}$
$v$$v_{j}$$v_{k}$$v_{\ell}$$v_{i}$
Figure 4. Two possible configurations of a $7(5)$-vertex $v$.
(5.2.4). Let $n_{4}(v)=5$. Clearly, $v$ has a support vertex, and so we have $d_{2}(v)\geq 21$ by Lemma 2.26(a). This particularly implies that $v$ has no 5(5)555(5)5 ( 5 )-neighbours by Lemma 2.19(a). Also, $v$ has at most one $5(4)$-neighbour since $d_{2}(v)\geq 21$. Suppose first that $v$ has a $5(4)$-neighbour or two $6(5)$-neighbours. Then we conclude that $v$ is incident to two $5^{+}$-face since $d_{2}(v)\geq 21$. By applying R2, $v$ receives $\frac{1}{5}$ from each of its incident $5^{+}$-faces. In such a case, we have $\mu^{*}(v)\geq 3+2\times\frac{1}{5}-5\times\frac{1}{3}-5\times\frac{1}{4}-\frac{1}{9}-\frac{1}{18}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), $\frac{1}{9}$ to its $5(4)$-neighbour by R7(d), $\frac{1}{18}$ to each of its $6(5)$-neighbours by R7(e). Now we suppose that $v$ has no $5(4)$-neighbour, and $v$ has at most one $6(5)$-neighbour. Then $\mu^{*}(v)\geq 3-5\times\frac{1}{3}-5\times\frac{1}{4}-\frac{1}{18}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), $\frac{1}{18}$ to its $6(5)$-neighbour by R7(e).
(5.2.5). Let $n_{4}(v)=6$. Clearly, $v$ has a support vertex, and so we have $d_{2}(v)\geq 21$ by Lemma 2.26(a), which implies that $v$ is incident to two $5^{+}$-faces. In addition, $v$ has no $5(4)$- and 5(5)555(5)5 ( 5 )-neighbours since $d_{2}(v)\geq 21$. By R2, $v$ receives $\frac{1}{5}$ from each of its incident $5^{+}$-faces. Thus we have $\mu^{*}(v)\geq 3+2\times\frac{1}{5}-5\times\frac{1}{3}-6\times\frac{1}{4}-\frac{1}{18}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), $\frac{1}{18}$ to its $6(5)$-neighbour by R7(e).
(5.3). Let $m_{3}(v)=6$. By Lemma 2.26(d), $v$ has at most five poor neighbours. If $v$ has at most four poor neighbours, then $\mu^{*}(v)\geq 3-6\times\frac{1}{3}-4\times\frac{1}{6}-3\times\frac{1}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{6}$ to each of its poor neighbours by R8(a),(c), $\frac{1}{9}$ to each of its $5(4)$-neighbours by R8(b). Suppose further that $v$ has exactly five poor neighbours. First, observe that $v$ has no $4(0)$-neighbours as $m_{3}(v)=6$, i.e., all $4$-neighbours of $v$ are bad $4$-vertices. Note that a $5(4)$-neighbour of $v$ is not adjacent to two poor vertices by Lemma 2.23(a),(b),(c). Moreover, a $4$-vertex has no 5(5)555(5)5 ( 5 )-neighbours by Lemma 2.19(a). Furthermore, a bad $4$-vertex has at most one $4$-neighbour by Lemma 2.19(b). In addition, by Lemma 2.24(b), a 5(5)555(5)5 ( 5 )-vertex is adjacent to at most one 5(5)555(5)5 ( 5 )-vertex. All those facts imply that $v$ has at most one $5(4)$-neighbour. Hence, $\mu^{*}(v)\geq 3-6\times\frac{1}{3}-5\times\frac{1}{6}-\frac{1}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{6}$ to each of its poor neighbours by R8(a),(c), $\frac{1}{9}$ to its $5(4)$-neighbour by R8(b).
(5.4). Let $m_{3}(v)=7$. By Lemma 2.24(b), a 5(5)555(5)5 ( 5 )-vertex is adjacent to at most one 5(5)555(5)5 ( 5 )-vertex. So, we deduce that $v$ has at most four 5(5)555(5)5 ( 5 )-neighbours. In particular, if $v$ has four such neighbours, then $v$ cannot have any $5(4)$-neighbours by Lemmas 2.23(a) and 2.24(b). Thus $\mu^{*}(v)\geq 3-7\times\frac{1}{3}-4\times\frac{1}{6}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1 and $\frac{1}{6}$ to each of its 5(5)555(5)5 ( 5 )-neighbours by R9(a). If $v$ has three 5(5)555(5)5 ( 5 )-neighbours, then $v$ has at most two $5(4)$-neighbours, and so $\mu^{*}(v)\geq 3-7\times\frac{1}{3}-3\times\frac{1}{6}-2\times\frac{1}{12}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-face by R1, $\frac{1}{6}$ to each of its 5(5)555(5)5 ( 5 )-neighbour by R9(a), and $\frac{1}{12}$ to each of its $5(4)$-neighbours by R9(b). Suppose now that $v$ has at most two 5(5)555(5)5 ( 5 )-neighbours. By Lemma 2.24(b), a 5(5)555(5)5 ( 5 )-vertex has at most one bad $5$-neighbour. We then deduce that $v$ has at most six bad $5$-neighbours when $v$ has at least one 5(5)555(5)5 ( 5 )-neighbour. Thus, $\mu^{*}(v)\geq 3-7\times\frac{1}{3}-2\times\frac{1}{6}-4\times\frac{1}{12}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-face by R1, $\frac{1}{6}$ to each of its 5(5)555(5)5 ( 5 )-neighbour by R9(a), and $\frac{1}{12}$ to each of its $5(4)$-neighbours by R9(b).
2.3. The case $\Delta=8$
Recall that $G$ does not admit any $2$-distance $23$-coloring, but any planar graph $G^{\prime}$ obtained from $G$ with smaller $|V(G^{\prime})|+|E(G^{\prime})|$ admits a 2-distance $23$-coloring.
Starting with some structural properties of $G$, we address similar results as in the case $\Delta=7$. The proof of the following lemma is omitted since it is similar to Lemma 2.1.
Lemma 2.27.
We have $\delta(G)\geq 3$.
Lemma 2.28.
Let $v$ be a $3$-vertex. Then $m_{3}(v)=0$ and $m_{4}(v)\leq 1$. In particular,
  • $(a)$
    $v$ has no $6^{-}$-neighbour,
  • $(b)$
    if $m_{4}(v)=0$, then $v$ has two $8(6^{-})$-neighbours,
  • $(c)$
    if $m_{4}(v)=1$, then $v$ has three $8(6^{-})$-neighbours.
Proof.
Let $v$ be a $3$-vertex. If $v$ is incident to a $3$-face, say $f_{1}=v_{1}vv_{2}$, then we set $G^{\prime}=G-v+\{v_{1}v_{3}\}$. If $v$ is incident to two $4$-faces, say $f_{1}=v_{1}vv_{2}x$ and $f_{2}=v_{2}vv_{3}y$, then we set $G^{\prime}=G-v+\{v_{1}v_{3}\}$. If $v$ is adjacent to a $6^{-}$-vertex, say $v_{1}$, then we set $G^{\prime}=G-v+\{v_{1}v_{2},v_{1}v_{3}\}$. In each case, $G^{\prime}$ remains proper with respect to $G$ as well as $d_{2}(v)\leq 22$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring with $23$ colors. Since $d_{2}(v)\leq 22$, we can color $v$ with an available color, a contradiction. This particularly implies that $v$ cannot have two (resp. one) $7$-neighbours when $m_{4}(v)=0$ (resp. $m_{4}(v)=1$). ∎
Recall that a $3$-vertex cannot have any $6^{-}$-neighbour by Lemma 2.28(a). That is, a vertex of degree $4$, $5$, or $6$ has no $3$-neighbour. This fact will be assumed in the sequel.
Lemma 2.29.
Let $v$ be a $4$-vertex with $m_{3}(v)\geq 1$.
  • $(a)$
    If $m_{3}(v)=1$, then $v$ has no 5(5)555(5)5 ( 5 )-neighbour. In particular, $v$ has at most one $4$-neighbour.
  • $(b)$
    If $m_{3}(v)=2$, then $v$ has at most one neighbour consisting of $4$- or $5$-vertices.
  • $(c)$
    If $m_{3}(v)\geq 3$, then $v$ has neither a $4$-neighbour nor a $5(4^{+})$-neighbour.
Proof.
Let $v$ be a $4$-vertex.
$(a)$ Let $m_{3}(v)=1$. Obviously, $v$ cannot have any 5(5)555(5)5 ( 5 )-neighbour as $m_{3}(v)=1$. Now, we assume that $v$ has two $4$-neighbours, and so $d_{2}(v)\leq 22$. If we set $G^{\prime}=G-v+\{v_{1}v_{2},v_{1}v_{3},v_{1}v_{4}\}$ for a $4$-neighbour $v_{1}$ of $v$, then $G^{\prime}$ is proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring with $23$ colors. Since $d_{2}(v)\leq 22$, we can color $v$ with an available color, a contradiction.
$(b)$ Let $m_{3}(v)=2$. Suppose that $v$ has two neighbours consisting of $4$- or $5$-vertices. So, $d_{2}(v)\leq 22$, and we get a contradiction as above.
$(c)$ Let $m_{3}(v)\geq 3$. If $v$ has a $4$- or $5(4^{+})$-neighbour, then we have $d_{2}(v)\leq 22$. In a similar way as above, we get a contradiction. ∎
Lemma 2.30.
Let $v$ be a $4(1)$-vertex. If $m_{4}(v)\geq 2$, then $v$ has two $7^{+}$-neighbours.
Proof.
Let $v$ be a $4(1)$-vertex. As $m_{3}(v)=1$, the vertex $v$ is incident to a $3$-face $f_{1}$, say $f_{1}=v_{1}vv_{2}$. Suppose that $m_{4}(v)\geq 2$. By contradiction, assume that $v$ has at most one $7^{+}$-neighbour. This implies that at least one of $v_{1},v_{2}$ is a $6^{-}$-vertex, say $v_{1}$. In particular, we have $d_{2}(v)\leq 22$. If we set $G^{\prime}=G-v+\{v_{1}v_{3},v_{1}v_{4}\}$, then $G^{\prime}$ is proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring with $23$ colors. Since $d_{2}(v)\leq 22$, we can color $v$ with an available color, a contradiction. ∎
Lemma 2.31.
Let $v$ be a $4(2)$-vertex.
  • $(a)$
    If $m_{4}(v)=0$, then $v$ has two $7^{+}$-neighbours different from 7(7)777(7)7 ( 7 )- and 8(8)888(8)8 ( 8 )-vertices.
  • $(b)$
    If $m_{4}(v)=1$, then $v$ has either two $8(7^{-})$- or three $7^{+}$-neighbours different from 7(7)777(7)7 ( 7 )- and 8(8)888(8)8 ( 8 )-vertices.
  • $(c)$
    If $m_{4}(v)=2$, then $v$ has either one $8(7^{-})$- and three $7(6^{-})$-neighbours or two $8(7^{-})$- and one $7(6^{-})$-neighbours or three $8(7^{-})$-neighbours.
Proof.
Let $v$ be a $4(2)$-vertex, and denote by $f_{1},f_{2},\ldots,f_{4}$ the faces incident to $v$ such that $f_{i}=v_{i}vv_{i+1}$ and $f_{j}=v_{j}vv_{j+1}$ with $i<j$ are $3$-faces. Notice first that if $v$ has a 7(7)777(7)7 ( 7 )- or a 8(8)888(8)8 ( 8 )-neighbours, then there exist two edges in $G[N(v)]$ such that each of them is contained in two $3$-faces.
$(a)$ Let $m_{4}(v)=0$. Assume for a contradiction that $v$ has at most one $7^{+}$-neighbour different from 7(7)777(7)7 ( 7 )- and 8(8)888(8)8 ( 8 )-vertices, i.e., $v$ has three neighbours consisting of $6^{-}$-, 7(7)777(7)7 ( 7 )- or 8(8)888(8)8 ( 8 )-vertices. In such a case, we clearly have $d_{2}(v)\leq 22$. If we set $G^{\prime}=G-v+\{v_{2}v_{4}\}$ (when $i=1$, $j=2$) or $G^{\prime}=G-v+\{v_{2}v_{3},v_{1}v_{4}\}$ (when $i=1$, $j=3$), then $G^{\prime}$ is proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring with $23$ colors. Since $d_{2}(v)\leq 22$, we can color $v$ with an available color, a contradiction.
$(b)$ Let $m_{4}(v)=1$. Notice that $v$ cannot have two neighbours consisting of 7(7)777(7)7 ( 7 )- or 8(8)888(8)8 ( 8 )-vertices as $m_{3}(v)=2$. Suppose first that $v$ has no 7(7)777(7)7 ( 7 )- and 8(8)888(8)8 ( 8 )-vertices. If $v$ has either two $6^{-}$- and one $7$-neighbours or three $6^{-}$-neighbours, then we would have $d_{2}(v)\leq 22$, and similarly as above, we get a contradiction. Thus, we conclude that $v$ has either two $8(7^{-})$-neighbours or three $7^{+}$-neighbours. Now we suppose that $v$ has a 7(7)777(7)7 ( 7 )- or 8(8)888(8)8 ( 8 )-neighbours. If $v$ has two $6^{-}$-neighbours, then we have $d_{2}(v)\leq 21$. Moreover, if $v$ has both a $6^{-}$-neighbour and a $7$-neighbour different from 7(7)777(7)7 ( 7 )-vertex, then we have $d_{2}(v)\leq 22$. Similarly as above, we get a contradiction. Thus, the claim holds.
$(c)$ Let $m_{4}(v)=2$. We first show that $v$ has at least one $8(7^{-})$-neighbours. Indeed, if all neighbours of $v$ are $7^{-}$- or 8(8)888(8)8 ( 8 )-vertices, then we have $d_{2}(v)\leq 22$ since $m_{4}(v)=2$. In a similar way as above, we get a contradiction. Therefore, $v$ has at least one $8(7^{-})$-neighbours. If $v$ has exactly one $8(7^{-})$-neighbours, and exactly two $7(6^{-})$-neighbours, or if $v$ has exactly two $8(7^{-})$-neighbours, and no $7(6^{-})$-neighbours, then we have $d_{2}(v)\leq 22$. Similarly as above, we get again a contradiction. ∎
Lemma 2.32.
Let $v$ be a $4(3)$-vertex.
  • $(a)$
    If $m_{4}(v)=0$, then $v$ has either one $8(7^{-})$- and three $7(6^{-})$-neighbours or two $8(7^{-})$- and one $7(6^{-})$-neighbours or two $8(7^{-})$- and two 8(8)888(8)8 ( 8 )-neighbours or three $8(7^{-})$-neighbours.
  • $(b)$
    If $m_{4}(v)=1$, then $v$ has either two $8(7^{-})$- and two $7(6^{-})$-neighbours or three $8(7^{-})$-neighbours.
Proof.
Let $v$ be a $4(3)$-vertex, and denote by $f_{1},f_{2},\ldots,f_{4}$ the faces incident to $v$ such that $f_{i}=v_{i}vv_{i+1}$ for each $i\in[3]$ is a $3$-face.
$(a)$ Let $m_{4}(v)=0$. We first show that $v$ has at least one $8(7^{-})$-neighbours. Indeed, if all neighbours of $v$ are $7^{-}$- or 8(8)888(8)8 ( 8 )-vertices, then we would have $d_{2}(v)\leq 22$. When we set $G^{\prime}=G-v+\{v_{1}v_{4}\}$, the graph $G^{\prime}$ remains proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring with $23$ colors. Since $d_{2}(v)\leq 22$, we can color $v$ with an available color, a contradiction. Thus $v$ has at least one $8(7^{-})$-neighbours. First suppose that $v$ has exactly one $8(7^{-})$-neighbour. If $v$ has no $7(6^{-})$-neighbours, then we have $d_{2}(v)\leq 22$. Moreover, if $v$ has one $7(6^{-})$-neighbour, and two neighbours consisting of $6^{-}$-, 7(7)777(7)7 ( 7 )- or 8(8)888(8)8 ( 8 )-vertices, then we would have $d_{2}(v)\leq 22$. Furthermore, if $v$ has two $7(6^{-})$-neighbours and one neighbour consisting of $6^{-}$-, 7(7)777(7)7 ( 7 )- or 8(8)888(8)8 ( 8 )-vertices, then we would have $d_{2}(v)\leq 22$. Similarly as above, we get a contradiction. We then conclude that $v$ has three $7(6^{-})$-neighbours. Now, we suppose that $v$ has exactly two $8(7^{-})$-neighbours. If $v$ has no $7(6^{-})$-neighbour, then $v$ is adjacent to two 8(8)888(8)8 ( 8 )-neighbours, since otherwise, we have $d_{2}(v)\leq 22$, and similarly as above, we get a contradiction. Thus, $v$ has either one $7(6^{-})$-neighbours or two 8(8)888(8)8 ( 8 )-neighbours.
$(b)$ Let $m_{4}(v)=1$. If $v$ has at most one $8(7^{-})$-neighbours, then we would have $d_{2}(v)\leq 22$, and similarly as above, we get a contradiction. Thus, $v$ has at least two $8(7^{-})$-neighbours. Moreover, if $v$ has two neighbours consisting of $6^{-}$-, 7(7)777(7)7 ( 7 )- or 8(8)888(8)8 ( 8 )-vertices, then we would have $d_{2}(v)\leq 22$. Furthermore, if $v$ has two $8(7^{-})$-neighbours, one $7(6^{-})$-neighbour, and one neighbour consisting of $6^{-}$-, 7(7)777(7)7 ( 7 )- or 8(8)888(8)8 ( 8 )-vertices, then we have again $d_{2}(v)\leq 22$. Similarly as above, we get a contradiction. ∎
Lemma 2.33.
Let $v$ be a 4(4)444(4)4 ( 4 )-vertex. Then, $v$ has either three $8(7^{-})$- and one $7(5^{-})$-neighbours or four $8(7^{-})$-neighbours.
Proof.
If $v$ has a $6^{-}$-, $7(6^{+})$- or 8(8)888(8)8 ( 8 )-neighbours, then we would have $d_{2}(v)\leq 22$. In addition, if $v$ has two $7^{-}$-neighbours, then we would have again $d_{2}(v)\leq 22$. When we set $G^{\prime}=G-v$, the graph $G^{\prime}$ remains proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring with $23$ colors. Since $d_{2}(v)\leq 22$, we can color $v$ with an available color, a contradiction. This implies that $v$ has either three $8(7^{-})$- and one $7(5^{-})$-neighbours or four $8(7^{-})$-neighbours. ∎
Lemma 2.34.
Let $v$ be a $5$-vertex.
  • $(a)$
    If $v$ is a 5(5)555(5)5 ( 5 )-vertex, and adjacent to a 5(5)555(5)5 ( 5 )-vertex, then $v$ has no $4$- and $5(4)$-neighbours.
  • $(b)$
    If $v$ is a 5(5)555(5)5 ( 5 )-vertex, then $v$ has at most one neighbour consisting of $4$- or 5(5)555(5)5 ( 5 )-vertices.
Proof.
$(a)$ Suppose that $v$ is a 5(5)555(5)5 ( 5 )-vertex adjacent to a 5(5)555(5)5 ( 5 )-vertex. If $v$ is also adjacent to a $4$- or $5(4)$-vertices, then we have $d_{2}(v)\leq 22$. When we set $G^{\prime}=G-v$, the graph $G^{\prime}$ remains proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring $f$ with $23$ colors. Since $d_{2}(v)\leq 22$, we can color $v$ with an available color, a contradiction.
$(b)$ Let $v$ be a 5(5)555(5)5 ( 5 )-vertex. Suppose that $v$ has two neighbours consisting of $4$- or 5(5)555(5)5 ( 5 )-vertices. If $v$ has a 5(5)555(5)5 ( 5 )-neighbour, then we get a contradiction by $(a)$. Else, if $v$ has two $4$-neighbours, then we have $d_{2}(v)\leq 22$, and similarly as above, we get a contradiction. ∎
Lemma 2.35.
Let $v$ be a $5(4)$-vertex.
  • $(a)$
    If $m_{4}(v)=0$, then $v$ has two $6^{+}$-neighbours different from 6(6)666(6)6 ( 6 )-vertex.
  • $(b)$
    If $m_{4}(v)=1$, then $v$ has three $6^{+}$-neighbours different from 6(6)666(6)6 ( 6 )-vertex.
Proof.
Let $v$ be a $5(4)$-vertex, and denote by $f_{1},f_{2},\ldots,f_{5}$ the faces incident to $v$ such that $f_{i}=v_{i}vv_{i+1}$ for $i\in[4]$.
$(a)$ Let $m_{4}(v)=0$. By contradiction, assume that $v$ has at most one $6^{+}$-neighbour different from 6(6)666(6)6 ( 6 )-vertex, i.e., all neighbours of $v$ but one consist of 6(6)666(6)6 ( 6 )- and $5^{-}$-vertices. Notice that if $v$ has a 6(6)666(6)6 ( 6 )-neighbour, then there exist two edges in $G[N(v)]$ such that each of them is contained in two $3$-faces. It can be easily observed that $v$ has no four 6(6)666(6)6 ( 6 )-neighbours. Consider all the other possibilities; $v$ has either three 6(6)666(6)6 ( 6 )-neighbours or two 6(6)666(6)6 ( 6 )-neighbour and one $5^{-}$-neighbours or one 6(6)666(6)6 ( 6 )-neighbour and two $5^{-}$-neighbours or four $5^{-}$-neighbours. In each case, we have $d_{2}(v)\leq 22$. When we set $G^{\prime}=G-v+\{v_{1}v_{5}\}$, the graph $G^{\prime}$ remains proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring $f$ with $23$ colors. Since $d_{2}(v)\leq 22$, we can color $v$ with an available color, a contradiction.
$(b)$ Let $m_{4}(v)=1$. Assume that $v$ has three neighbours $x,y,z$ consisting of $5^{-}$- or 6(6)666(6)6 ( 6 )-vertices. Note that if $v$ has a 6(6)666(6)6 ( 6 )-neighbour, then there exist two edges in $G[N(v)]$ such that each of them is contained in two $3$-faces. Consider all the possibilities; all $x,y,z$ are $5^{-}$-vertices; two of $x,y,z$ are $5^{-}$-vertices and the last is 6(6)666(6)6 ( 6 )-vertex; one of $x,y,z$ is a $5^{-}$-vertex and the others are 6(6)666(6)6 ( 6 )-vertices; all $x,y,z$ are 6(6)666(6)6 ( 6 )-vertices. In each case, we have $d_{2}(v)\leq 22$, and similarly as above, we get a contradiction. ∎
Lemma 2.36.
Let $v$ be a 5(5)555(5)5 ( 5 )-vertex.
  • $(a)$
    If $n_{4}(v)=1$, then $v$ has three $7^{+}$-neighbours different from 7(7)777(7)7 ( 7 )-vertex.
  • $(b)$
    If $n_{4}(v)=n_{5}(v)=0$, then $v$ has either three $8(7^{-})$-neighbours or four $6^{+}$-neighbours different from 6(6)666(6)6 ( 6 )-vertex such that two of which are $7^{+}$-vertices different from 7(7)777(7)7 ( 7 )-vertex.
  • $(c)$
    If $n_{4}(v)=0$ and $n_{5}(v)=1$, then $v$ has either three $7^{+}$-neighbours different from 7(7)777(7)7 ( 7 )-vertex or two $6(5^{-})$- and two $7^{+}$-neighbours different from 7(7)777(7)7 ( 7 )-vertex.
  • $(d)$
    If $n_{4}(v)=0$ and $n_{5}(v)=2$, then $v$ has three $7^{+}$-neighbours different from 7(7)777(7)7 ( 7 )-vertex.
Proof.
Let $v$ be a 5(5)555(5)5 ( 5 )-vertex, and denote by $f_{1},f_{2},\ldots,f_{5}$ the $3$-faces incident to $v$ such that $f_{i}=v_{i}vv_{i+1}$ for $i\in[5]$ in a cyclic fashion.
$(a)$ Let $n_{4}(v)=1$, and assume that $v$ has at most two $7^{+}$-neighbours different from 7(7)777(7)7 ( 7 )-vertex, i.e., $v$ has two neighbours consisting of $5$-, $6$- or 7(7)777(7)7 ( 7 )-vertices. It is clear that $d_{2}(v)\leq 22$. When we set $G^{\prime}=G-v$, the graph $G^{\prime}$ remains proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring $f$ with $23$ colors. Since $d_{2}(v)\leq 22$, we can color $v$ with an available color, a contradiction.
$(b)$ Let $n_{4}(v)=n_{5}(v)=0$. Assume that $v$ has at most two $8(7^{-})$-neighbours. If $v$ has at most three $6^{+}$-neighbours different from 6(6)666(6)6 ( 6 )-vertex, then $v$ has two 6(6)666(6)6 ( 6 )-neighbours as $n_{4}(v)=n_{5}(v)=0$. In particular, $v$ has a neighbour forming $7$- or 8(8)888(8)8 ( 8 )-vertices. In each case, we have $d_{2}(v)\leq 22$, and similarly as above, we get a contradiction. Thus, $v$ has four $6^{+}$-neighbours different from 6(6)666(6)6 ( 6 )-vertex. If only one of those neighbours is a $7^{+}$-vertex different from 7(7)777(7)7 ( 7 )-vertex, then we again have $d_{2}(v)\leq 22$, and similarly as above, we get a contradiction.
$(c)$ Let $n_{4}(v)=0$ and $n_{5}(v)=1$. If $v$ has at most one $7^{+}$-neighbour different from 7(7)777(7)7 ( 7 )-vertex, then we have $d_{2}(v)\leq 22$, and similarly as above, we get a contradiction. Thus, $v$ has at least two $7^{+}$-neighbours different from 7(7)777(7)7 ( 7 )-vertex. Suppose that $v$ has exactly two $7^{+}$-neighbours different from 7(7)777(7)7 ( 7 )-vertex. If $v$ has at most one $6(5^{-})$-neighbour, i.e., $v$ has one $5$-neighbour, at most one $6(5^{-})$ and one neighbour consisting of 6(6)666(6)6 ( 6 )- or 7(7)777(7)7 ( 7 )-vertices, then we have again $d_{2}(v)\leq 22$. Similarly as above, we get a contradiction.
$(d)$ Let $n_{4}(v)=0$ and $n_{5}(v)=2$. If $v$ has at most two $7^{+}$-neighbours different from 7(7)777(7)7 ( 7 )-vertex, then we have $d_{2}(v)\leq 22$, and similarly as above, we get a contradiction. ∎
Lemma 2.37.
Let $v$ be a $6(5)$-vertex. Then, $v$ has at most four $5(4^{+})$-neighbours. In particular, if $m_{4}(v)=1$ and $v$ has four $5(4^{+})$-neighbours, then $v$ has two $8(6^{-})$-neighbours.
Proof.
Let $v$ be a $6(5)$-vertex, and denote by $f_{1},f_{2},\ldots,f_{6}$ the faces incident to $v$ such that $f_{i}=v_{i}vv_{i+1}$ for $i\in[5]$.
Assume for a contradiction that $v$ has five $5(4^{+})$-neighbours. This implies that there exist at least three edges in $G[N(v)]$ such that each of them is contained in two $3$-faces, and so we have $d_{2}(v)\leq 20$. If we set $G^{\prime}=G-v$, the graph $G^{\prime}$ is proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring $f$ with $23$ colors. Since $d_{2}(v)\leq 20$, we can color $v$ with an available color, a contradiction.
Now we suppose that $m_{4}(v)=1$, and $v$ has four $5(4^{+})$-neighbours. Since $v$ has four $5(4^{+})$-neighbours, there exist at least two edges in $G[N(v)]$ such that each of them is contained in two $3$-faces. If $v$ has at most one $8(6^{-})$-neighbour, then we would have $d_{2}(v)\leq 22$. Similarly as above, we get a contradiction. ∎
Let $v$ be a $7^{+}$-vertex and let $x$ be a $5^{-}$-neighbour of $v$. If $v$ and $x$ have two common neighbours, i.e., the edge $vx$ is contained in two $3$-faces, then $x$ is called a support vertex of $v$.
Lemma 2.38.
Let $v$ be a $7^{+}$-vertex. If $v$ has a support vertex, then $d_{2}(v)\geq 23$.
Proof.
Let $v$ be a $7^{+}$-vertex. Suppose that $v_{1}$ is a support vertex of $v$. By definition, $v_{1}$ is a $5^{-}$-vertex and it is adjacent to both $v_{2}$ and $v_{7}$. By contradiction, assume that $d_{2}(v)\leq 22$. When we set $G^{\prime}=G-v+\{v_{1}v_{3},v_{1}v_{4},v_{1}v_{5},v_{1}v_{6}\}$, the graph $G^{\prime}$ remains proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring $f$ with $23$ colors. Since $d_{2}(v)\leq 22$, we can color $v$ with an available color, a contradiction. ∎
We next apply discharging to show that $G$ does not exist. We use the same initial charges as before, with the following discharging rules.
Discharging Rules
We apply the following discharging rules.
  • R1:
    Every $3$-face receives $\frac{1}{3}$ from each of its incident vertices.
  • R2:
    Every $5^{+}$-face gives $\frac{1}{5}$ to each of its incident vertices.
  • R3:
    Every $3$-vertex receives $\frac{1}{5}$ from each of its $8(6^{-})$-neighbour.
  • R4:
    Every $4(1)$-vertex receives $\frac{1}{6}$ from each of its $7(6^{-})$- and $8(7^{-})$-neighbours.
  • R5:
    Let $v$ be a $4(2)$-vertex. Then,
    • $(a)$
      $v$ receives $\frac{1}{6}$ from each of its $7(6^{-})$-neighbours,
    • $(b)$
      $v$ receives $\frac{1}{4}$ from each of its $8(7^{-})$-neighbours.
  • R6:
    Let $v$ be a $4(3)$-vertex. Then,
    • $(a)$
      $v$ receives $\frac{1}{6}$ from each of its $7(6^{-})$-neighbours,
    • $(b)$
      $v$ receives $\frac{1}{3}$ from each of its $8(7^{-})$-neighbours,
    • $(c)$
      $v$ receives $\frac{1}{9}$ from each of its 8(8)888(8)8 ( 8 )-neighbours.
  • R7:
    Every 4(4)444(4)4 ( 4 )-vertex receives $\frac{1}{3}$ from each of its $7(5^{-})$- and $8(7^{-})$-neighbours.
  • R8:
    Every $5(4)$-vertex receives $\frac{1}{9}$ from each of its $6^{+}$-neighbours different from 6(6)666(6)6 ( 6 )-vertex.
  • R9:
    Let $v$ be a 5(5)555(5)5 ( 5 )-vertex. Then,
    • $(a)$
      $v$ receives $\frac{1}{9}$ from each of its $6(5^{-})$- and 7(7)777(7)7 ( 7 )-neighbours,
    • $(b)$
      $v$ receives $\frac{2}{9}$ from each of its $7(6^{-})$- and $8$-neighbours.
  • R10:
    Every $6(5)$-vertex receives $\frac{1}{18}$ from each of its $8(6^{-})$-neighbours.
Checking $\mu^{*}(v),\mu^{*}(f)\geq 0$ for $v\in V(G),f\in F(G)$.
First we show that $\mu^{*}(f)\geq 0$ for each $f\in F(G)$. Given a face $f\in F(G)$, if $f$ is a $3$-face, then it receives $\frac{1}{3}$ from each of its incident vertices by R1, and so $\mu^{*}(f)\geq\ell(f)-4+3\times\frac{1}{3}=0$. If $f$ is a $4$-face, then $\mu(f)=\mu^{*}(f)=0$. Let $f$ be a $5^{+}$-face. By applying R2, $f$ sends $\frac{1}{5}$ to each of its incident vertices. It then follows that $\mu^{*}(f)\geq\ell(f)-4-5\times\frac{1}{5}\geq 0$. Consequently, $\mu^{*}(f)\geq 0$ for each $f\in F(G)$.
We now pick a vertex $v\in V(G)$ with $d(v)=k$. By Lemma 2.27, we have $k\geq 3$.
(1). Let $k=3$. The initial charge of $v$ is $\mu(v)=d(v)-4=-1$. By Lemma 2.28, $m_{3}(v)=0$ and $m_{4}(v)\leq 1$. If $m_{4}(v)=0$, then $v$ is adjacent to two $8(6^{-})$-vertices by Lemma 2.28(b), and so $\mu^{*}(v)\geq-1+3\times\frac{1}{5}+2\times\frac{1}{5}=0$ after $v$ receives $\frac{1}{5}$ from each of its incident $5^{+}$-faces by R2, and $\frac{1}{5}$ from each of its $8(6^{-})$-neighbours by R3. On the other hand, if $m_{4}(v)=1$, then $v$ is adjacent to three $8(6^{-})$-neighbours by Lemma 2.28(c), and so $\mu^{*}(v)\geq-1+2\times\frac{1}{5}+3\times\frac{1}{5}=0$ after $v$ receives $\frac{1}{5}$ from each of its incident $5^{+}$-faces by R2, and $\frac{1}{5}$ from each of its $8(6^{-})$-neighbours by R3.
(2). Let $k=4$. The initial charge of $v$ is $\mu(v)=d(v)-4=0$. If $m_{3}(v)=0$, then $\mu^{*}(v)\geq 0$, since $v$ does not give any charge to its incident faces. Therefore, we may assume that $1\leq m_{3}(v)\leq 4$.
(2.1). Let $m_{3}(v)=1$. If $m_{4}(v)\leq 1$, then $v$ is incident to two $5^{+}$-faces, and by R2, $v$ receives $\frac{1}{5}$ from each of those $5^{+}$-faces. Thus, $\mu^{*}(v)\geq 2\times\frac{1}{5}-\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to its incident $3$-face by R1. If $m_{4}(v)\geq 2$, then $v$ has two $7^{+}$-neighbours by Lemma 2.30, which are clearly different from 7(7)777(7)7 ( 7 )- and 8(8)888(8)8 ( 8 )-vertices as $m_{3}(v)=1$. It follows that $v$ receives $\frac{1}{6}$ from each of its $7^{+}$-neighbours by R4. Thus, $\mu^{*}(v)\geq 2\times\frac{1}{6}-\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to its incident $3$-face by R1.
(2.2). Let $m_{3}(v)=2$. Since $0\leq m_{4}(v)\leq 2$, we consider the following cases:
If $m_{4}(v)=0$, then $v$ has two $7^{+}$-neighbours different from 7(7)777(7)7 ( 7 )- and 8(8)888(8)8 ( 8 )-vertices by Lemma 2.31(a). It follows that $v$ receives totally at least $2\times\frac{1}{6}$ from its $7^{+}$-neighbours by R5. In addition, $v$ is incident to two $5^{+}$-faces, and by R2, $v$ receives $\frac{1}{5}$ from each of those $5^{+}$-faces. Thus, $\mu^{*}(v)\geq 2\times\frac{1}{6}+2\times\frac{1}{5}-2\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
If $m_{4}(v)=1$, then $v$ has either two $8(7^{-})$-neighbours or three $7^{+}$-neighbours different from 7(7)777(7)7 ( 7 )- and 8(8)888(8)8 ( 8 )-vertices by Lemma 2.31(b). In each case, $v$ receives totally at least $\frac{1}{2}$ from its $7^{+}$-neighbours by R5. Also, $v$ receives $\frac{1}{5}$ from its incident $5^{+}$-face by R2. Thus, $\mu^{*}(v)\geq\frac{1}{2}+\frac{1}{5}-2\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
If $m_{4}(v)=2$, then $v$ has either one $8(7^{-})$- and three $7(6^{-})$-neighbours or two $8(7^{-})$- and one $7(6^{-})$-neighbours or three $8(7^{-})$-neighbours by Lemma 2.31(c). In each case, $v$ receives totally at least $\frac{2}{3}=\min\{\frac{1}{4}+3\times\frac{1}{6},2\times\frac{1}{4}+\frac{1}{6},3\times\frac{1}{4}\}$ from its $7^{+}$-neighbours by R5. Thus, $\mu^{*}(v)\geq\frac{2}{3}-2\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
(2.3). Let $m_{3}(v)=3$. Suppose first that $m_{4}(v)=0$. It then follows from Lemma 2.32(a) that $v$ has either one $8(7^{-})$- and three $7(6^{-})$-neighbours or two $8(7^{-})$- and one $7(6^{-})$-neighbours or two $8(7^{-})$- and two 8(8)888(8)8 ( 8 )-neighbours or three $8(7^{-})$-neighbours. In each case, $v$ receives totally at least $\frac{5}{6}=\min\{\frac{1}{3}+3\times\frac{1}{6},2\times\frac{1}{3}+\frac{1}{6},2\times\frac{1}{3}+2\times\frac{1}{9},3\times\frac{1}{3}\}$ from its $7^{+}$-neighbours by R6. In addition, $v$ receives $\frac{1}{5}$ from its incident $5^{+}$-face by R2. Thus, $\mu^{*}(v)\geq\frac{5}{6}+\frac{1}{5}-3\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1. Now, we suppose that $m_{4}(v)=1$. By Lemma 2.32(b), $v$ has either two $8(7^{-})$- and two $7(6^{-})$-neighbours or three $8(7^{-})$-neighbours. In both cases, $v$ receives totally at least $1=\min\{2\times\frac{1}{3}+2\times\frac{1}{6},3\times\frac{1}{3}\}$ from its $7^{+}$-neighbours by R6. Thus, $\mu^{*}(v)\geq 1-3\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
(2.4). Let $m_{3}(v)=4$. By Lemma 2.33, $v$ has either three $8(7^{-})$- and one $7(5^{-})$-neighbour or four $8(7^{-})$-neighbours. By applying R7, $v$ receives $\frac{1}{3}$ from each of its $7^{+}$-neighbours. Thus, $\mu^{*}(v)\geq 4\times\frac{1}{3}-4\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
(3). Let $k=5$. The initial charge of $v$ is $\mu(v)=d(v)-4=1$. If $m_{3}(v)\leq 3$, then $\mu^{*}(v)\geq 1-3\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1. Therefore, we may assume that $4\leq m_{3}(v)\leq 5$.
(3.1). Let $m_{3}(v)=4$. Suppose first that $m_{4}(v)=0$. By Lemma 2.35(a), $v$ has two $6^{+}$-neighbours different from 6(6)666(6)6 ( 6 )-vertex. It follows that $v$ receives $\frac{1}{9}$ from each of its $6^{+}$-neighbours different from 6(6)666(6)6 ( 6 )-vertex by R8, and $\frac{1}{5}$ from its incident $5^{+}$-face by R2. Thus, $\mu^{*}(v)\geq 1+2\times\frac{1}{9}+\frac{1}{5}-4\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1. Next we suppose that $m_{4}(v)=1$. By Lemma 2.35(b), $v$ has three $6^{+}$-neighbours different from 6(6)666(6)6 ( 6 )-vertex. So, $v$ receives $\frac{1}{9}$ from each of its $6^{+}$-neighbours different from 6(6)666(6)6 ( 6 )-vertex by R8. Thus, $\mu^{*}(v)\geq 1+3\times\frac{1}{9}-4\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
(3.2). Let $m_{3}(v)=5$. We distinguish two subcases according to the number of $4$-vertices adjacent to $v$ as follows. Recall that $0\leq n_{4}(v)\leq 1$ by Lemma 2.34(b).
(3.2.1). Let $n_{4}(v)=0$. If $n_{5}(v)=0$, then, by Lemma 2.36(b), $v$ has either three $8(7^{-})$-neighbours or four $6^{+}$-neighbours different from 6(6)666(6)6 ( 6 )-vertex, two of which are $7^{+}$-vertices different from 7(7)777(7)7 ( 7 )-vertex. It follows from applying R9 that $v$ receives totally at least $\frac{2}{3}=\min\{3\times\frac{2}{9},2\times\frac{1}{9}+2\times\frac{2}{9}\}$ from its $6^{+}$-neighbours. Thus, $\mu^{*}(v)\geq 1+\frac{2}{3}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
If $n_{5}(v)=1$, then, by Lemma 2.36(c), $v$ has either three $7^{+}$-neighbours different from 7(7)777(7)7 ( 7 )-vertex or two $6(5^{-})$- and two $7^{+}$-neighbours different from 7(7)777(7)7 ( 7 )-vertex. So, $v$ receives totally at least $\frac{2}{3}=\min\{3\times\frac{2}{9},2\times\frac{1}{9}+2\times\frac{2}{9}\}$ from its $6^{+}$-neighbours by R9. Thus, $\mu^{*}(v)\geq 1+\frac{2}{3}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
If $n_{5}(v)=2$, then $v$ has three $7^{+}$-neighbours different from 7(7)777(7)7 ( 7 )-vertex by Lemma 2.36(d). By R9(b), $v$ receives $\frac{2}{9}$ from each of its $7^{+}$-neighbours different from 7(7)777(7)7 ( 7 )-vertex. Thus, $\mu^{*}(v)\geq 1+3\times\frac{2}{9}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
(3.2.2). Let $n_{4}(v)=1$. By Lemma 2.36(a), $v$ has three $7^{+}$-neighbours different from 7(7)777(7)7 ( 7 )-vertex. It follows that $v$ receives $\frac{2}{9}$ from each of its $7^{+}$-neighbours different from 7(7)777(7)7 ( 7 )-vertex by R9(b). So, $\mu^{*}(v)\geq 1+3\times\frac{2}{9}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
(4). Let $k=6$. The initial charge of $v$ is $\mu(v)=d(v)-4=2$. If $m_{3}(v)\leq 4$, then $\mu^{*}(v)\geq 2-4\times\frac{1}{3}-6\times\frac{1}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and $\frac{1}{9}$ to each of its $5(4^{+})$-neighbours by R8 and R9(a). On the other hand, if $m_{3}(v)=6$, then $\mu^{*}(v)\geq 2-6\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1. Therefore, we may assume that $m_{3}(v)=5$. By Lemma 2.37, $v$ has at most four $5(4^{+})$-neighbours. If $v$ has at most three $5(4^{+})$-neighbours, then $\mu^{*}(v)\geq 2-5\times\frac{1}{3}-3\times\frac{1}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and $\frac{1}{9}$ to each of its $5(4^{+})$-neighbours by R8 and R9(a). Next we assume that $v$ has exactly four $5(4^{+})$-neighbours. If $m_{4}(v)=0$, then $v$ receives $\frac{1}{5}$ from its incident $5^{+}$-face by R2, and so $\mu^{*}(v)\geq 2+\frac{1}{5}-5\times\frac{1}{3}-4\times\frac{1}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and $\frac{1}{9}$ to each of its $5(4^{+})$-neighbours by R8 and R9(a). If $m_{4}(v)=1$, then $v$ has two $8(6^{-})$-neighbours by Lemma 2.37, and so $v$ receives $\frac{1}{18}$ from each of its $8(6^{-})$-neighbours by R10. Thus, $\mu^{*}(v)\geq 2+2\times\frac{1}{18}-5\times\frac{1}{3}-4\times\frac{1}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and $\frac{1}{9}$ to each of its $5(4^{+})$-neighbours by R8 and R9(a).
(5). Let $k=7$. The initial charge of $v$ is $\mu(v)=d(v)-4=3$. Notice first that if $m_{3}(v)\leq 3$, then we have $\mu^{*}(v)\geq 3-3\times\frac{1}{3}-2\times\frac{1}{3}-5\times\frac{2}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its 4(4)444(4)4 ( 4 )-neighbours by R7, and at most $\frac{2}{9}$ to each of its other neighbours by R4-R6 and R8-R9, where we note that $v$ can have at most two 4(4)444(4)4 ( 4 )-neighbours due to $m_{3}(v)\leq 3$. Therefore, we may assume that $4\leq m_{3}(v)\leq 7$.
(5.1). Let $m_{3}(v)=4$. Observe that $v$ can have at most two neighbours consisting of 4(4)444(4)4 ( 4 )- or 5(5)555(5)5 ( 5 )-vertices by Lemmas 2.29(c) and 2.34(b). If $v$ has at most one 4(4)444(4)4 ( 4 )-neighbour, then $\mu^{*}(v)\geq 3-4\times\frac{1}{3}-\frac{1}{3}-6\times\frac{2}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to its 4(4)444(4)4 ( 4 )-neighbour by R7, and at most $\frac{2}{9}$ to each of its other neighbours by R4-R6 and R8-R9. If $v$ has exactly two 4(4)444(4)4 ( 4 )-neighbours, then $v$ is not adjacent to any 5(5)555(5)5 ( 5 )-vertex as stated earlier, and so $\mu^{*}(v)\geq 3-4\times\frac{1}{3}-2\times\frac{1}{3}-5\times\frac{1}{6}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its 4(4)444(4)4 ( 4 )-neighbours by R7, and at most $\frac{1}{6}$ to each of its other neighbours by R4-R6 and R8.
(5.2). Let $m_{3}(v)=5$. By Lemmas 2.29(c) and 2.34(b), $v$ has at most two 4(4)444(4)4 ( 4 )-neighbours. If $v$ has no 4(4)444(4)4 ( 4 )-neighbour, then $v$ is adjacent to at most three 5(5)555(5)5 ( 5 )-neighbour by Lemma 2.34(b), and so $\mu^{*}(v)\geq 3-5\times\frac{1}{3}-3\times\frac{2}{9}-4\times\frac{1}{6}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{2}{9}$ to each of its 5(5)555(5)5 ( 5 )-neighbours by R9(b), and at most $\frac{1}{6}$ to each of its other neighbours by R4-R6 and R8. If $v$ has exactly one 4(4)444(4)4 ( 4 )-neighbour, then $v$ is adjacent to at most two 5(5)555(5)5 ( 5 )-neighbours by Lemmas 2.29(c) and 2.34(b). Moreover, $v$ has at least two neighbours $v_{i}$ and $v_{j}$ that are neither $4$-vertices nor $5(4^{+})$-vertices by Lemma 2.29(c), since $v$ has a 4(4)444(4)4 ( 4 )-neighbour. Clearly, neither $v_{i}$ nor $v_{j}$ receives any charge from $v$. Therefore, $\mu^{*}(v)\geq 3-5\times\frac{1}{3}-\frac{1}{3}-2\times\frac{2}{9}-2\times\frac{1}{6}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to its 4(4)444(4)4 ( 4 )-neighbour by R7, $\frac{2}{9}$ to each of its 5(5)555(5)5 ( 5 )-neighbours by R9(b), and at most $\frac{1}{6}$ to each of its $4(1^{+})$- and $5(4)$-neighbours different from 4(4)444(4)4 ( 4 )-vertex by R4-R6 and R8. If $v$ has exactly two 4(4)444(4)4 ( 4 )-neighbours, then $v$ is not adjacent to any 5(5)555(5)5 ( 5 )-vertices by Lemma 2.29(c). Moreover, $v$ has at least three neighbors $v_{i}$, $v_{j}$, and $v_{k}$ that are neither $4$-vertices nor $5(4^{+})$-vertices by Lemma 2.29(c), since $v$ has two 4(4)444(4)4 ( 4 )-neighbors. Clearly, none of $v_{i}$, $v_{j}$, or $v_{k}$ receives any charge from $v$. Thus, $\mu^{*}(v)\geq 3-5\times\frac{1}{3}-2\times\frac{1}{3}-2\times\frac{1}{6}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its 4(4)444(4)4 ( 4 )-neighbours by R7, and at most $\frac{1}{6}$ to each of its $4(1^{+})$- and $5(4)$-neighbours different from 4(4)444(4)4 ( 4 )-vertex by R4-R6 and R8.
(5.3). Let $m_{3}(v)=6$. Notice first that if $v$ has at most four neighbours forming $4$- or $5$-vertices, then we have $\mu^{*}(v)\geq 3-6\times\frac{1}{3}-4\times\frac{2}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and at most $\frac{2}{9}$ to each of its $5^{-}$-neighbours by R4-R6 and R8-R9. Therefore we may further assume that $v$ has at least five neighbours forming $4$- or $5$-vertices, which clearly infers that $v$ has at least one support vertex. Thus we have $d_{2}(v)\geq 23$ by Lemma 2.38. On the other hand, by Lemma 2.29, we conclude that a $4(1^{+})$-vertex has at most one $4$-neighbour. In addition, a 5(5)555(5)5 ( 5 )-vertex has at most one neighbour consisting of $4$- or 5(5)555(5)5 ( 5 )-vertices by Lemma 2.34(b). This implies that $v$ has at most four 5(5)555(5)5 ( 5 )-neighbours, and at most five $4(1^{+})$-vertices.
Suppose first that $v$ has no 5(5)555(5)5 ( 5 )-neighbours. If $v$ has at most four $4(1^{+})$-neighbours, then $\mu^{*}(v)\geq 3-6\times\frac{1}{3}-4\times\frac{1}{6}-3\times\frac{1}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{6}$ to each of its $4(1^{+})$-neighbours by R4-R6, and $\frac{1}{9}$ to each of its $5(4)$-neighbours by R8. If $v$ has five $4(1^{+})$-neighbours, then $v$ cannot have any $5(4)$-neighbours as $d_{2}(v)\geq 23$. So, $\mu^{*}(v)\geq 3-6\times\frac{1}{3}-5\times\frac{1}{6}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and $\frac{1}{6}$ to each of its $4(1^{+})$-neighbours by R4-R6.
Next we suppose that $v$ has exactly one 5(5)555(5)5 ( 5 )-neighbour. Since $d_{2}(v)\geq 23$, we can say that $v$ has at most four $4$-neighbours. If $v$ has at most two $4$-neighbours, then $\mu^{*}(v)\geq 3-6\times\frac{1}{3}-\frac{2}{9}-2\times\frac{1}{6}-4\times\frac{1}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{2}{9}$ to its 5(5)555(5)5 ( 5 )-neighbour by R9(b), $\frac{1}{6}$ to each of its $4(1^{+})$-neighbours by R4-R6, and $\frac{1}{9}$ to each of its $5(4)$-neighbours by R8. If $v$ has at least three $4$-neighbours, then $v$ is adjacent to at most one $5(4)$-vertices as $d_{2}(v)\geq 23$, and so $\mu^{*}(v)\geq 3-6\times\frac{1}{3}-\frac{2}{9}-4\times\frac{1}{6}-\frac{1}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{2}{9}$ to each of its 5(5)555(5)5 ( 5 )-neighbours by R9(b), $\frac{1}{6}$ to each of its $4(1^{+})$-neighbours by R4-R6, and $\frac{1}{9}$ to its $5(4)$-neighbour by R8.
Suppose now that $v$ has exactly two 5(5)555(5)5 ( 5 )-neighbours. Then we deduce that $v$ has at most three neighbours consisting of $4$- or $5(4)$-vertices as $d_{2}(v)\geq 23$, and so $\mu^{*}(v)\geq 3-6\times\frac{1}{3}-2\times\frac{2}{9}-3\times\frac{1}{6}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{2}{9}$ to each of its 5(5)555(5)5 ( 5 )-neighbours by R9(b), at most $\frac{1}{6}$ to each of its $4(1^{+})$ and $5(4)$-neighbours by R4-R6 and R8.
Suppose that $v$ has exactly three 5(5)555(5)5 ( 5 )-neighbours. It follows that $v$ has at most one $4$-neighbour as $d_{2}(v)\geq 23$ together with Lemma 2.34(a). By the same reason, $v$ has at most two neighbours consisting of $4$- or $5(4)$-vertices. Thus, we have $\mu^{*}(v)\geq 3-6\times\frac{1}{3}-3\times\frac{2}{9}-\frac{1}{6}-\frac{1}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{2}{9}$ to each of its 5(5)555(5)5 ( 5 )-neighbours by R9(b), $\frac{1}{6}$ to each of its $4(1^{+})$-neighbours by R4-R6, and $\frac{1}{9}$ to each of its $5(4)$-neighbours by R8.
Finally, suppose that $v$ has four 5(5)555(5)5 ( 5 )-neighbours. Then $v$ has no $4$- and $5(4)$-neighbours by Lemma 2.34(a),(b), and so $\mu^{*}(v)\geq 3-6\times\frac{1}{3}-4\times\frac{2}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{2}{9}$ to each of its 5(5)555(5)5 ( 5 )-neighbours by R9(b).
(5.4). Let $m_{3}(v)=7$. By Lemma 2.38, we deduce that $v$ has at most five $5(4^{+})$-neighbours, and so $\mu^{*}(v)\geq 3-7\times\frac{1}{3}-5\times\frac{1}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-face by R1, $\frac{1}{9}$ to each of its $5(4^{+})$-neighbour by R8-R9.
(6). Let $k=8$. The initial charge of $v$ is $\mu(v)=d(v)-4=4$. Notice first that if $m_{3}(v)\leq 4$, then we have $\mu^{*}(v)\geq 4-4\times\frac{1}{3}-8\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and at most $\frac{1}{3}$ to each of its neighbours by R3-R10. Therefore, we may assume that $5\leq m_{3}(v)\leq 8$.
(6.1). Let $m_{3}(v)=5$. By Lemma 2.29(c), a $4(3^{+})$-vertex cannot be adjacent to any $4$-vertex. It follows that $v$ has at most five $4(3^{+})$-neighbours. If $v$ has at most four $4(3^{+})$-neighbours, then $\mu^{*}(v)\geq 4-5\times\frac{1}{3}-4\times\frac{1}{3}-4\times\frac{1}{4}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its $4(3^{+})$-neighbours by R6(b), R7, and at most $\frac{1}{4}$ to each of its other neighbours by R3-R5 and R8-R10. If $v$ has five $4(3^{+})$-neighbours, then $v$ cannot have any $3$-, $4(2^{-})$- or $5(4^{+})$-neighbour by Lemmas 2.28(a) and 2.29(c). Thus $\mu^{*}(v)\geq 4-5\times\frac{1}{3}-5\times\frac{1}{3}-3\times\frac{1}{18}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its $4(3^{+})$-neighbours by R6(b) and R7, $\frac{1}{18}$ to each of its $6(5)$-neighbours by R10.
(6.2). Let $m_{3}(v)=6$. By Lemma 2.29(c), $v$ has at most five $4(3^{+})$-neighbours. If $v$ has five such neighbours, then $v$ is not adjacent to any $3$-, $4(2^{-})$- or $5(4^{+})$-vertex by Lemmas 2.28(a) and 2.29(c). Thus $\mu^{*}(v)\geq 4-6\times\frac{1}{3}-5\times\frac{1}{3}-3\times\frac{1}{18}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its $4(3^{+})$-neighbours by R6(b) and R7, $\frac{1}{18}$ to each of its $6(5)$-neighbours by R10. We may further assume that $v$ has at most four $4(3^{+})$-neighbours.
Suppose first that $v$ has no $4(3^{+})$-neighbour. Then $\mu^{*}(v)\geq 4-6\times\frac{1}{3}-8\times\frac{1}{4}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and at most $\frac{1}{4}$ to each of its neighbours by R3-R5 and R8-R10.
Next suppose that $v$ has exactly one $4(3^{+})$-neighbour. It follows from Lemmas 2.28(a) and 2.29(c) that $v$ has at most six neighbours consisting of $3$-, $4(2^{-})$- or $5(4^{+})$-vertices. Thus $\mu^{*}(v)\geq 4-6\times\frac{1}{3}-\frac{1}{3}-6\times\frac{1}{4}-\frac{1}{18}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its $4(3^{+})$-neighbours by R6(b) and R7, at most $\frac{1}{4}$ to each of its $3$-, $4(1)$-, $4(2)$- and $5(4^{+})$-neighbours by R3-R5 and R8-R9, $\frac{1}{18}$ to each of its $6(5)$-neighbours by R10.
Suppose now that $v$ has exactly two $4(3^{+})$-neighbours. In such a case, $v$ has at most five neighbours consisting of $3$-, $4(2^{-})$- or $5(4^{+})$-vertices by Lemmas 2.28(a) and 2.29(c). Thus $\mu^{*}(v)\geq 4-6\times\frac{1}{3}-2\times\frac{1}{3}-5\times\frac{1}{4}-\frac{1}{18}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its $4(3^{+})$-neighbours by R6(b) and R7, at most $\frac{1}{4}$ to each of its $3$-, $4(1)$-, $4(2)$- and $5(4^{+})$-neighbours by R3-R5 and R8-R9, $\frac{1}{18}$ to each of its $6(5)$-neighbours by R10.
Suppose that $v$ has exactly three $4(3^{+})$-neighbours. In such a case, $v$ can have at most three neighbours consisting of $3$-, $4(2^{-})$- or $5(4^{+})$-vertices by Lemmas 2.28(a) and 2.29(c). Thus $\mu^{*}(v)\geq 4-6\times\frac{1}{3}-3\times\frac{1}{3}-3\times\frac{1}{4}-2\times\frac{1}{18}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its $4(3^{+})$-neighbours by R6(b) and R7, at most $\frac{1}{4}$ to each of its $3$-, $4(1)$-, $4(2)$- and $5(4^{+})$-neighbours by R3-R5 and R8-R9, $\frac{1}{18}$ to each of its $6(5)$-neighbours by R10.
Finally suppose that $v$ has exactly four $4(3^{+})$-neighbours. In such a case, $v$ can have at most one neighbour consisting of $3$-, $4(2^{-})$- or $5(4^{+})$-vertices by Lemmas 2.28(a) and 2.29(c). Thus $\mu^{*}(v)\geq 4-6\times\frac{1}{3}-4\times\frac{1}{3}-\frac{1}{4}-3\times\frac{1}{18}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its $4(3^{+})$-neighbours by R6(b) and R7, and at most $\frac{1}{4}$ to its $3$-, $4(1)$-, $4(2)$- and $5(4^{+})$-neighbours by R3-R5 and R8-R9, $\frac{1}{18}$ to each of its $6(5)$-neighbours by R10.
(6.3). Let $m_{3}(v)=7$. Notice first that if $v$ has at most five $5^{-}$-neighbours, then we have $\mu^{*}(v)\geq 4-7\times\frac{1}{3}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and at most $\frac{1}{3}$ to each of its $5^{-}$-neighbours by R4-R9. Therefore we may assume that $v$ has at least six $5^{-}$-neighbours, which clearly infers that $v$ has at least one support vertex. Thus we have $d_{2}(v)\geq 23$ by Lemma 2.38. Observe that $v$ has at most six $4$-neighbours as $d_{2}(v)\geq 23$. In particular, by Lemma 2.29(c), a $4(3^{+})$-vertex has no $4$- and $5(4^{+})$-neighbours. So, we deduce that $v$ has at most four $4(3^{+})$-neighbours. If $v$ has four such neighbours, then $v$ is not adjacent to any $4(2^{-})$- or $5(4^{+})$-vertices by Lemma 2.29(c). Thus $\mu^{*}(v)\geq 4-7\times\frac{1}{3}-4\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its $4(3^{+})$-neighbours by R6(b) and R7. We may further assume that $v$ has at most three $4(3^{+})$-neighbours.
Suppose first that $v$ has no $4(3^{+})$-neighbour. Note that $v$ has at most six neighbours consisting of $4(2^{-})$- or 5(5)555(5)5 ( 5 )-vertices as $d_{2}(v)\geq 23$. If $v$ has six such neighbours, then $v$ cannot be adjacent to any $5(4)$-vertex as $d_{2}(v)\geq 23$, and so $\mu^{*}(v)\geq 4-7\times\frac{1}{3}-6\times\frac{1}{4}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, at most $\frac{1}{4}$ to each of its $4(1)$-, $4(2)$- and 5(5)555(5)5 ( 5 )-neighbours by R4-R5 and R9(b). If $v$ has at most five neighbours consisting of $4(2^{-})$- or 5(5)555(5)5 ( 5 )-vertices, then $\mu^{*}(v)\geq 4-7\times\frac{1}{3}-5\times\frac{1}{4}-3\times\frac{1}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, at most $\frac{1}{4}$ to each of its $4(1)$-, $4(2)$- and 5(5)555(5)5 ( 5 )-neighbours by R4-R5 and R9(b), $\frac{1}{9}$ to each of its $5(4)$-neighbours by R8.
Next we suppose that $v$ has exactly one $4(3^{+})$-neighbour. By Lemma 2.29(c), $v$ has at least one neighbour different from $4$- and $5(4^{+})$-vertices. On the other hand, a $4(t)$-vertex with $1\leq t\leq 2$ has at most one neighbour consisting of $4$- or 5(5)555(5)5 ( 5 )-vertices by Lemma 2.29(a),(b). It then follows from Lemma 2.29(c) and Lemma 2.34(b) that $v$ has at most four neighbours consisting of $4(1)$-, $4(2)$- or 5(5)555(5)5 ( 5 )-vertices. Thus $\mu^{*}(v)\geq 4-7\times\frac{1}{3}-\frac{1}{3}-4\times\frac{1}{4}-2\times\frac{1}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its $4(3^{+})$-neighbours by R6(b) and R7, at most $\frac{1}{4}$ to each of its $4(1)$-, $4(2)$- and 5(5)555(5)5 ( 5 )-neighbours by R4-R5 and R9(b), $\frac{1}{9}$ to each of its $5(4)$-neighbours by R8.
Suppose that $v$ has exactly two $4(3^{+})$-neighbours. By Lemma 2.29(c), a $4(3^{+})$-vertex has neither $4$-neighbour nor $5(4^{+})$-neighbour. This means that $v$ has at least two neighbours different from $4$- and $5(4^{+})$-vertices. Thus $\mu^{*}(v)\geq 4-7\times\frac{1}{3}-2\times\frac{1}{3}-4\times\frac{1}{4}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its $4(3^{+})$-neighbours by R6(b) and R7, and at most $\frac{1}{4}$ to each of its $4(1)$-, $4(2)$- and $5(4^{+})$-neighbours by R4-R5 and R8-R9.
Finally suppose that $v$ has exactly three $4(3^{+})$-neighbours. It follows from Lemma 2.29(c) that $v$ has at least three neighbours different from $4$- and $5(4^{+})$-vertices. Thus $\mu^{*}(v)\geq 4-7\times\frac{1}{3}-3\times\frac{1}{3}-2\times\frac{1}{4}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its $4(3^{+})$-neighbours by R6(b) and R7, at most $\frac{1}{4}$ to each of its $4(1)$-, $4(2)$- and $5(4^{+})$-neighbours by R4-R5 and R8-R9.
(6.4). Let $m_{3}(v)=8$. By Lemma 2.34(b), $v$ has at most five 5(5)555(5)5 ( 5 )-neighbours. If $v$ has at most four such neighbours, then $\mu^{*}(v)\geq 4-8\times\frac{1}{3}-4\times\frac{2}{9}-4\times\frac{1}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{2}{9}$ to each of its 5(5)555(5)5 ( 5 )-neighbours by R9(b), and at most $\frac{1}{9}$ to each of its $4(3)$- and $5(4)$-neighbours by R6(c) and R8. If $v$ has five 5(5)555(5)5 ( 5 )-neighbours, then $v$ has neither $4$-neighbour nor $5(4)$-neighbour by Lemma 2.34(a). Thus, $\mu^{*}(v)\geq 4-8\times\frac{1}{3}-5\times\frac{2}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and $\frac{2}{9}$ to each of its 5(5)555(5)5 ( 5 )-neighbours by R9(b).
Acknowledgement
The author extends their gratitude to Kengo Aoki for pointing out a significant error in a previous proof of Lemma 2.1.