### v3

# A $2$-distance $(2\Delta+7)$-coloring of planar graphs

Abstract.
A vertex coloring of a graph $G$ is called a $2$-distance coloring if any two vertices at a distance at most $2$ from each other receive different colors. Recently, Bousquet et al. (Discrete Mathematics, 346(4), 113288, 2023) proved that $2\Delta+7$ colors are sufficient for the $2$-distance coloring of planar graphs with maximum degree $\Delta\geq 9$. In this paper, we strengthen their result by removing the maximum degree constraint and show that all planar graphs admit a 2-distance $(2\Delta+7)$-coloring. This particularly improves the result of Van den Heuvel and McGuinness (Journal of Graph Theory, 42(2), 110-124, 2003).
Key words and phrases:
Coloring, 2-distance coloring, girth, planar graph.
1. Introduction
Among the numerous problems and concepts associated with graphs, coloring stands out as a fundamental and extensively studied topic. A vertex coloring of a graph assigns colors to its vertices so that adjacent vertices receive distinct colors. One particularly interesting variant of vertex coloring is 2-distance coloring, where any two vertices at a distance of $2$ or less have distinct colors. Its motivation arises from the necessity in some real-life problems to assign different colors not only to adjacent vertices but also to those in close proximity [14, 16]. The concept of 2-distance coloring was first introduced in [12, 13], and it has received considerable attention due to the fact that certain problems, such as the Total Coloring Conjecture (see [5] for details), can be formulated as a 2-distance coloring of specific graphs. A comprehensive survey on 2-distance coloring and its related coloring concepts has been presented by Cranston [5].
The smallest number of colors for which graph $G$ admits a 2-distance coloring is known as the 2-distance chromatic number $\chi_{2}(G)$ of $G$. In 1977, Wegner [19] proposed the following conjecture.
Conjecture 1.1.
For every planar graph $G$ with maximum degree $\Delta$, $\chi_{2}(G)\leq 7$ if $\Delta=3$, $\chi_{2}(G)\leq\Delta+5$ if $4\leq\Delta\leq 7$, and $\chi_{2}(G)\leq\big\lfloor\frac{3\Delta}{2}\big\rfloor+1$ if $\Delta\geq 8$.
Wegner’s Conjecture is one of the most popular problems in graph coloring, and it has remained open for many years, except for the case when $\Delta=3$, which was solved by Thomassen [18] (independently by Hartke et al. [9]). The conjecture is also known to be asymptotically true for large values of $\Delta$, as shown in [1, 10]. For general planar graphs, Van den Heuvel and McGuinness [11] showed that $\chi_{2}(G)\leq 2\Delta+25$, while the bound $\chi_{2}(G)\leq\big\lceil\frac{5\Delta}{3}\big\rceil+78$ was proved by Molloy and Salavatipour [17]. On the other hand, some improved results are presented in [3, 4, 6, 7, 8] with a certain degree or girth restrictions. In particular, Bousquet et al. [2] recently showed that $\chi_{2}(G)\leq 2\Delta+7$ if $\Delta\geq 9$, which significantly improves upon the work in [15].
In this paper, we strengthen the result of Bousquet et al. [2] by removing the maximum degree constraint and show that all planar graphs admit a 2-distance $(2\Delta+7)$-coloring, which particularly improves the best-known bound of $2\Delta+25$ provided by Van den Heuvel and McGuinness [11].
Theorem 1.2.
For every planar graph $G$, we have $\chi_{2}(G)\leq 2\Delta+7$.
All graphs in this paper are assumed to be simple, and we refer to [20] for terminology and notation not defined here. When $G$ is a graph, we use $V(G),E(G),F(G),$ and $\Delta(G)$ to denote the vertex, edge and face set, and the maximum degree of $G$, respectively. If there is no confusion in the context, we abbreviate $\Delta(G)$ to $\Delta$. For a positive integer $t$, we denote by $[t]$ the set of integers between $1$ and $t$. Given a planar graph $G$, we denote by $\ell(f)$ the length of a face $f$ and by $d(v)$ the degree of a vertex $v$. A $k$-vertex is a vertex of degree $k$. A $k^{-}$-vertex is a vertex of degree at most $k$ while a $k^{+}$-vertex is a vertex of degree at least $k$. A $k$ ($k^{-}$ or $k^{+}$)-face is defined analogously as for the vertices, where the degree of a face is the number of edges incident to it. A vertex $u\in N(v)$ is called $k$-neighbour (resp. $k^{-}$-neighbour, $k^{+}$-neighbour) of $v$ if $d(u)=k$ (resp. $d(u)\leq k$, $d(u)\geq k$).
For a vertex $v\in V(G)$, we use $n_{i}(v)$ to denote the number of $i$-vertices adjacent to $v$. We denote by $d(u,v)$ the distance between $u$ and $v$ for a pair $u,v\in V(G)$. Also, we set $N_{i}(v):=\{u\in V(G)\ |\ 1\leq d(u,v)\leq i\}$ for $i\geq 1$, so $N_{1}(v)=N(v)$ and let $d_{2}(v)=|N_{2}(v)|$. For $v\in V(G)$, we use $m_{k}(v)$ to denote the number of $k$-faces incident with $v$. A $k$-vertex $v$ with $m_{3}(v)=d$ is called $k(d)$-vertex. In particular, a $k$-vertex $v$ is called $k(d^{-})$-vertex (resp. $k(d^{+})$-vertex) if $m_{3}(v)\leq d$ (resp. $m_{3}(v)\geq d$). Two faces $f_{1}$ and $f_{2}$ are said to be adjacent if they share a common edge.
2. The Proof of Theorem 1.2
Let $G$ be a minimal counterexample to Theorem 1.2 such that $|V(G)|+|E(G)|$ is minimum. Thus, $G$ does not admit any $2$-distance $(2\Delta+7)$-coloring, whereas every planar graph $G^{\prime}$ obtained from $G$ with $|V(G^{\prime})|+|E(G^{\prime})|<|V(G)|+|E(G)|$ admits a $2$-distance $(2\Delta+7)$-coloring. By the results of [2, 6], it suffices to consider the case $6\leq\Delta\leq 8$. Clearly, $G$ is connected.
We call a graph $H$ proper with respect to $G$ if $H$ is obtained from $G$ by deleting some edges or vertices and adding some edges, in such a way that every pair of vertices $x_{1},x_{2}\in V(G)\cap V(H)$ that have distance at most $2$ in $G$ also have distance at most $2$ in $H$. If $f$ is a $2$-distance coloring of such a graph $H$, then $f$ can be extended to a $2$-distance coloring of $G$, provided that each remaining uncolored vertex has an available color. By the minimality of $G$, we obtain the following.
Remark 2.1.
If $H$ is proper with respect to $G$ such that $V(G)\setminus V(H)=\{v\}$, then $d_{2}(v)\geq 2\Delta+7$.
Proof.
Assume, for a contradiction, that there exists a vertex $v\in V(G)\setminus V(H)$ such that $d_{2}(v)\leq 2\Delta+6$. Since $H$ is proper with respect to $G$, by the minimality of $G$, the graph $H$ admits a $2$-distance coloring using $2\Delta+7$ colors. As $d_{2}(v)\leq 2\Delta+6$, there is at least one available color for $v$, and thus the coloring of $H$ extends to a $2$-distance coloring of $G$, a contradiction to the minimality of $G$. ∎
For a vertex $v$ with $d(v)=k$, let $v_{1},v_{2},\ldots,v_{k}$ denote the neighbors of $v$ in clockwise order, and let $f_{1},f_{2},\ldots,f_{k}$ be the faces incident to $v$. Define $E(v)=\{v_{i}v_{i+1}\in E(G):i\in[k]\,\}$, where indices are taken modulo $k$. Let $t(v)$ denote the number of edges in $E(v)$ that are contained in two $3$-faces. We begin with a useful lemma that provides an upper bound on $d_{2}(v)$.
Lemma 2.2.
For any vertex $v$, we have $d_{2}(v)\leq\left(\sum_{u\in N(v)}d(u)\right)-2m_{3}(v)-m_{4}(v)-t(v)$.
Proof.
Recall that $d_{2}(v)$ counts the number of vertices at distance at most $2$ from $v$, that is, $d_{2}(v)=|N_{2}(v)|$. Let $p=\sum_{u\in N(v)}d(u)$. From $p$, we subtract $2$ for each $3$-face incident to $v$ and $1$ for each $4$-face incident to $v$. Furthermore, for each edge $v_{i}v_{i+1}$, if $v_{i}$ and $v_{i+1}$ have a common neighbor other than $v$, we subtract an additional $1$ from $p$. Consequently, we obtain $d_{2}(v)\leq p-2m_{3}(v)-m_{4}(v)-t(v).$
In the remainder of the paper, we divide the proof into three cases according to $\Delta\in\{6,7,8\}$. In each case, we employ the same overall strategy: we first derive structural properties from the minimality of $G$, and then introduce an appropriate set of discharging rules and apply the discharging method to reach a contradiction, thereby showing that $G$ cannot exist.
2.1. The case $\Delta=6$
Recall that $G$ does not admit a $2$-distance 1919-coloring, whereas every planar graph $G^{\prime}$ obtained from $G$ with a smaller value of $|V(G^{\prime})|+|E(G^{\prime})|$ admits one.
Lemma 2.3.
$\delta(G)\geq 3$.
Proof.
If $v$ is a vertex of degree at most $2$, then $d_{2}(v)\leq 12$. Let $G^{\prime}$ be the graph obtained from GvG-v by adding an edge between the vertices in $N(v)$. Notice that $G^{\prime}$ is proper with respect to $G$. By Remark 2.1, this yields a contradiction. ∎
Lemma 2.4.
Let $v$ be a $3$-vertex. Then $m_{3}(v)=0$, $m_{4}(v)\leq 1$, and every neighbor of $v$ is a $6(4^{-})$-vertex.
Proof.
Assume first, for a contradiction, that $v$ is incident to a $3$-face, say $f_{1}=v_{1}vv_{2}$. By Lemma 2.2, we have $d_{2}(v)\leq 16$. If we set $G^{\prime}=G-v+\{v_{1}v_{3}\}$, then $G^{\prime}$ is proper with respect to $G$, yielding a contradiction to Remark 2.1.
Suppose next that $m_{4}(v)\geq 2$, that is, $v$ is incident to two $4$-faces, say $f_{1}=v_{1}vv_{2}x$ and $f_{2}=v_{2}vv_{3}y$. Again, $d_{2}(v)\leq 16$. If we set $G^{\prime}=G-v+\{v_{1}v_{3}\}$, then $G^{\prime}$ is proper with respect to $G$, giving the same contradiction as above.
We now show that every neighbor of $v$ is a $6(4^{-})$-vertex. Since $m_{3}(v)=0$, the vertex $v$ cannot have a $6(5^{+})$-neighbor. Moreover, $v$ cannot have a $5^{-}$-neighbor. Indeed, suppose $v_{1}$ is such a neighbor. Let $G^{\prime}=G-v+\{v_{1}v_{2},v_{1}v_{3}\}$. In this case $d_{2}(v)\leq 17$, and $G^{\prime}$ is proper with respect to $G$, again contradicting Remark 2.1. Hence every neighbor of $v$ must be a $6(4^{-})$-vertex. ∎
We now introduce some terminology for vertices of special types. A $4(1)$- or $4(2)$-vertex is called a bad $4$-vertex. Similarly, a $5(4)$- or $5(5)$-vertex is called a bad $5$-vertex.
Lemma 2.5.
If $v$ is a $4$-vertex, then $m_{3}(v)\leq 2$, and $v$ has no $k(k)$-neighbour for any $k\leq 6$.
Proof.
Let $v$ be a $4$-vertex. Assume, for a contradiction, that $m_{3}(v)\geq 3$. Let $f_{i}=v_{i}vv_{i+1}$ for $i\in[3]$. By Lemma 2.2, we have $d_{2}(v)\leq 18$. If we define $G^{\prime}=G-v+\{v_{1}v_{4}\}$ (assuming $v_{1}v_{4}\notin E(G)$), then $G^{\prime}$ is proper with respect to $G$. This contradicts Remark 2.1.
Now suppose that $v$ has a $k(k)$-neighbour for some $k\leq 6$. Without loss of generality, assume that $v_{2}$ is a $k(k)$-vertex. Then $v$ must be incident with two consecutive $3$-faces $f_{1}$ and $f_{2}$, say $f_{1}=v_{1}vv_{2}$ and $f_{2}=v_{2}vv_{3}$. Moreover, each of the edges $v_{1}v_{2}$ and $v_{2}v_{3}$ must be contained in two $3$-faces. Hence $d_{2}(v)\leq 18$ by Lemma 2.2. If we set $G^{\prime}=G-v+\{v_{2}v_{4}\}$, then $G^{\prime}$ would again be proper with respect to $G$, giving the same contradiction. ∎
Lemma 2.6.
Let $v$ be a $4(1)$-vertex. If $v$ has a $4$-neighbour, then the other neighbours of $v$ are neither $4$-vertices nor bad $5$-vertices.
Proof.
Suppose that $v$ has a $4$-neighbour, say $v_{1}$, and assume for a contradiction that $v$ has another neighbour that is either a $4$-vertex or a $5(4^{+})$-vertex. In this case, Lemma 2.2 implies $d_{2}(v)\leq 18$. If we form $G^{\prime}=G-v+\{v_{1}v_{2},\,v_{1}v_{3},\,v_{1}v_{4}\}$, then $G^{\prime}$ is proper with respect to $G$. By Remark 2.1, this yields a contradiction. ∎
Lemma 2.7.
Let $v$ be a $4(1)$-vertex. If $m_{4}(v)=r$ for $0\leq r\leq 3$, then $n_{6}(v)\geq r+1$.
Proof.
Let $f_{1}=v_{1}vv_{2}$ be the unique $3$-face incident to $v$. Assume $m_{4}(v)=r$ for some $0\leq r\leq 3$. Suppose, for a contradiction, that $n_{6}(v)\leq r$. Then $v$ has 4r4-r neighbours of degree at most $5$, and is incident to $r$ 44-faces. By Lemma 2.2, this implies $d_{2}(v)\leq 18$. In particular, either $v$ has two $5^{-}$-neighbours, or $v$ is incident to three $4$-faces. We show that in each case, we can construct a graph $G^{\prime}$ proper with respect to $G$, which gives a contradiction.
If $v$ has two $5^{-}$-neighbours $v_{i},v_{j}$ such that $\{i,j\}\cap\{1,2\}\neq\emptyset$, then we set $G^{\prime}=G-v+\{v_{k}v_{3},v_{k}v_{4}\}$ for $k\in\{i,j\}\cap\{1,2\}$. If $v$ has two $5^{-}$-neighbours $v_{i},v_{j}$ with $i<j$ such that $\{i,j\}\cap\{1,2\}=\emptyset$, then we set $G^{\prime}=G-v+\{v_{i}v_{j},v_{i}v_{2},v_{j}v_{1}\}$. If $v$ is incident to exactly three $4$-faces, then we set $G^{\prime}=G-v+\{v_{2}v_{3},v_{1}v_{4}\}$. In each case, $G^{\prime}$ is proper with respect to $G$. By Remark 2.1, this yields a contradiction. Hence, $n_{6}(v)\geq r+1$, as desired. ∎
Lemma 2.8.
Let $v$ be a $4(2)$-vertex. Then $m_{4}(v)\leq 1$. In particular, if $m_{4}(v)=r$ for $0\leq r\leq 1$, then $n_{6}(v)\geq r+3$.
Proof.
Suppose that $f_{1}$ and $f_{2}$ are $3$-faces incident to $v$, where $f_{1}=v_{i}vv_{i+1}$ and $f_{2}=v_{j}vv_{j+1}$ with $i<j$, and $i,j\in[4]$ taken cyclically. Assume, for a contradiction, that $v$ is incident to two $4$-faces. Then, by Lemma 2.2, we have $d_{2}(v)\leq 18$. If $f_{1}$ and $f_{2}$ are adjacent, choose $v_{k}\in N(v)\setminus\{v_{i},v_{i+1},v_{j},v_{j+1}\}$ and set $G^{\prime}=G-v+\{v_{j}v_{k}\}$. Otherwise, set $G^{\prime}=G-v+\{v_{i}v_{j+1},\,v_{i+1}v_{j}\}$. In both cases, the resulting graph $G^{\prime}$ is proper with respect to $G$. By Remark 2.1, this yields a contradiction. Thus $m_{4}(v)\leq 1$.
Now assume $m_{4}(v)=0$ and $n_{6}(v)\leq 2$. Then $v$ has two $5^{-}$-neighbours, and so Lemma 2.2 gives $d_{2}(v)\leq 18$. Similarly, if $m_{4}(v)=1$ and $n_{6}(v)\leq 3$, then $v$ has a $5^{-}$-neighbour, and again $d_{2}(v)\leq 18$. In each case, applying the same modification of $G$ as above yields a proper graph $G^{\prime}$, contradicting Remark 2.1. Thus $n_{6}(v)\geq r+3$ for $r=m_{4}(v)$. ∎
Lemma 2.9.
Let $v$ be a $4(2)$-vertex. Then, $v$ has neither a $4$-neighbour nor a bad $5$-neighbour.
Proof.
Suppose first that one of $v_{i}$’s is a $4$-vertex, say $v_{1}$. Obviously, we have $d_{2}(v)\leq 18$, and let $G^{\prime}=G-v+\{v_{1}v_{2},v_{1}v_{3},v_{1}v_{4}\}$, assuming these edges are not already present. Observe that $G^{\prime}$ is proper with respect to $G$. This contradicts Remark 2.1.
Suppose now that $v$ has a bad $5$-neighbour, say $v_{i}$. Then, there exists an edge $v_{i}v_{j}$ for $v_{j}\in N(v)$ such that $v_{i}v_{j}$ is contained in two $3$-faces. By Lemma 2.2, this implies $d_{2}(v)\leq 18$. If we set $G^{\prime}=G-v+\{v_{i}v_{p},v_{i}v_{t}\}$ for $v_{p},v_{t}\in N(v)\setminus\{v_{i},v_{j}\}$, then $G^{\prime}$ would be proper with respect to $G$. Similarly as above, we get a contradiction. ∎
Lemma 2.10.
Let $v$ be a $5(4)$-vertex. If $v$ has a $4$-neighbour, then the other neighbours of $v$ are neither $4$-vertices nor $5(4)$-vertices.
Proof.
Let $f_{i}=v_{i}vv_{i+1}$ for $i\in[4]$. Suppose that $v$ has a $4$-neighbour, and assume for a contradiction that $v$ has another neighbour that is either a $4$-vertex or a $5(4)$-vertex. Then, by Lemma 2.2, we obtain $d_{2}(v)\leq 18$. If we define $G^{\prime}=G-v+\{v_{1}v_{5}\}$, then $G^{\prime}$ is proper with respect to $G$. By Remark 2.1, this yields a contradiction. ∎
The following is an immediate consequence of Lemma 2.10, since a $5(4)$-vertex cannot simultaneously have a $4$-neighbour and a $5(4)$-neighbour.
Corollary 2.11.
If $v$ is a $4(1)$-vertex, then $v$ has no two $5(4)$-neighbours.
Lemma 2.12.
Let $v$ be a $5(4)$-vertex.
  • $(a)$
    If $m_{4}(v)=0$, then $v$ has at least two $6(5^{-})$-neighbours.
  • $(b)$
    If $m_{4}(v)=1$, then $v$ has at least three $6(5^{-})$-neighbours.
Proof.
Let $f_{i}=v_{i}vv_{i+1}$ for $i\in[4]$.
$(a)$ Let $m_{4}(v)=0$ and assume, for a contradiction, that $v$ has at most one $6(5^{-})$-neighbour. Then $v$ has four neighbours that are all $6(6)$- or $5^{-}$-vertices, and hence $d_{2}(v)\leq 18$ by Lemma 2.2. If we define $G^{\prime}=G-v+\{v_{1}v_{5}\}$, then $G^{\prime}$ is proper with respect to $G$. By Remark 2.1, this yields a contradiction.
$(b)$ Let $m_{4}(v)=1$ and assume, again for a contradiction, that $v$ has at most two $6(5^{-})$-neighbours. Then $v$ has three neighbours that are $6(6)$- or $5^{-}$-vertices, implying $d_{2}(v)\leq 18$ by Lemma 2.2. As in the previous case, this leads to a contradiction. ∎
Proposition 2.13.
A $5(4)$-vertex cannot have two non-adjacent $5(4)$-neighbours.
Proof.
Let $v$ be a $5(4)$-vertex, and let $f_{i}=v_{i}vv_{i+1}$ for $i\in[4]$. Assume that $v$ has two $5(4)$-neighbours $v_{i}$ and $v_{j}$ such that $|i-j|\geq 2$. Then either $v_{i}v_{i-1}$ or $v_{i}v_{i+1}$ (in cyclic order) is contained in two $3$-faces. Similarly, either $v_{j}v_{j-1}$ or $v_{j}v_{j+1}$ is contained in two $3$-faces. Thus $d_{2}(v)\leq 18$ by Lemma 2.2. Setting $G^{\prime}=G-v+\{v_{1}v_{5}\}$ yields a graph $G^{\prime}$ that is proper with respect to $G$, giving a contradiction by Remark 2.1. ∎
Lemma 2.14.
Let $v$ be a $5(5)$-vertex. Then $v$ has neither a $4$-neighbour nor a bad $5$-neighbour nor a $6(6)$-neighbour. In particular, $n_{5}(v)\leq 1$.
Proof.
Let $f_{i}=v_{i}vv_{i+1}$ for $i\in[5]$ in cyclic order. By Lemma 2.5, $v$ has no $4$-neighbour. Now suppose that $v$ has either a bad $5$-neighbour, or a $6(6)$-neighbour, or two $5$-neighbours. In each of these cases, Lemma 2.2 gives $d_{2}(v)\leq 18$. If we set $G^{\prime}=G-v$, then $G^{\prime}$ remains proper with respect to $G$. By Remark 2.1, this contradicts the minimality of $G$. ∎
The following is an immediate consequence of Lemma 2.14.
Corollary 2.15.
If $v$ is a $5(5)$-vertex, then $v$ has four $6(5^{-})$-neighbours.
An edge $uv$ is said to be special if $v$ is a $5(5)$-vertex and $uv$ is contained in two $3$-faces $f_{1},f_{2}$, each of which is adjacent to a $4^{+}$-face (see Figure 1). Note that $u$ is a $5^{+}$-vertex by Lemmas 2.4 and 2.5,
[…]
Figure 1. A special edge uvuv with a $5(5)$-vertex $v$ and a $5^{+}$-vertex $u$.
Proposition 2.16.
Let $v$ be a $5(5)$-vertex, and let $v_{i}v_{i+1}\in E(v)$.
  • $(a)$
    If $v_{i}v_{i+1}$ is contained in two $3$-faces, then all neighbours of $v$ are $6(5^{-})$-vertices. In particular, for every vertex $v_{j}\in N(v)\setminus\{v_{i},v_{i+1}\}$, the edge $vv_{j}$ is a special edge.
  • $(b)$
    If no edge in $E(v)$ is contained in two $3$-faces, then each edge $vv_{i}$ is a special edge for every $v_{i}\in N(v)$, and $v$ has four $6(4^{-})$-neighbours.
Proof.
Let $f_{i}=v_{i}vv_{i+1}$ for $i\in[5]$ in cyclic order.
$(a)$ Suppose that the edge $v_{i}v_{i+1}$ is contained in two $3$-faces. If $v$ has either a $5^{-}$-neighbour or a $6(6)$-neighbour, then Lemma 2.2 implies $d_{2}(v)\leq 18$. Removing $v$ (i.e., taking $G^{\prime}=G-v$) yields a graph proper with respect to $G$, contradicting Remark 2.1. Hence every neighbour of $v$ must be a $6(5^{-})$-vertex.
Now suppose that there exists another neighbour $v_{j}\in N(v)\setminus\{v_{i},v_{i+1}\}$ such that either $v_{j}v_{j-1}$ or $v_{j}v_{j+1}$ is contained in two $3$-faces. Then Lemma 2.2 again gives $d_{2}(v)\leq 18$, producing the same contradiction. Therefore, for every $v_{j}\in N(v)\setminus\{v_{i},v_{i+1}\}$, the edge $vv_{j}$ is a special edge.
$(b)$ Suppose no edge in $E(v)$ is contained in two $3$-faces. Then, by definition of a special edge, each $vv_{i}$ is a special edge for all $v_{i}\in N(v)$. Since a $5(5)$-vertex has no bad $5$-neighbour and no $6(6)$-neighbour by Lemma 2.14, it follows from Corollary 2.15 that $v$ has four $6(5^{-})$-neighbours. ∎
Lemma 2.17.
Let $v$ be a $6(5)$-vertex. Then $n_{4}(v)\leq 3$. In particular:
  • $(a)$
    If $n_{4}(v)=3$, then $v$ has no bad $5$-neighbour.
  • $(b)$
    If $n_{4}(v)=2$, then $v$ has at most two bad $5$-neighbours.
  • $(c)$
    If $n_{4}(v)=2$ and $v$ has two bad $5$-neighbours, then $v$ has two $6(4^{-})$-neighbours.
  • $(d)$
    If $n_{4}(v)=1$, then $v$ has at most three bad $5$-neighbours.
  • $(e)$
    If $n_{4}(v)=1$ and $v$ has three bad $5$-neighbours, then $v$ has a $6(4^{-})$-neighbour.
  • $(f)$
    If $n_{4}(v)=0$, then $v$ has at most four bad $5$-neighbours.
  • $(g)$
    If $n_{4}(v)=0$, $m_{4}(v)=1$, and $v$ has four bad $5$-neighbours, then $v$ has two $6(4^{-})$-neighbours.
Proof.
Let $f_{i}=v_{i}vv_{i+1}$ for $i\in[5]$. Recall that $v$ has no $3$-neighbour by Lemma 2.4.
We claim that $n_{4}(v)\leq 3$. Indeed, if $v$ had four $4$-neighbours, then Lemma 2.2 would imply $d_{2}(v)\leq 18$. Among $v_{2},v_{3},v_{4},v_{5}$, at least one must be a $4$-vertex; assume $v_{2}$ is such. Then setting $G^{\prime}=G-v+\{v_{2}v_{4},v_{2}v_{5},v_{2}v_{6}\}$ yields a graph that is proper with respect to $G$, contradicting Remark 2.1.
$(a)$ Let $n_{4}(v)=3$. If $v$ had a bad $5$-neighbour $v_{i}$, then by Lemma 2.2 we would have $d_{2}(v)\leq 18$. Similarly as above, this leads to a contradiction.
$(b)$ Let $n_{4}(v)=2$. If $v$ had three bad $5$-neighbours, then by Lemma 2.2 we would have $d_{2}(v)\leq 17$. If one of $v_{2},v_{3},v_{4},v_{5}$ is a $4$-vertex (say $v_{2}$), take $G^{\prime}=G-v+\{v_{2}v_{4},v_{2}v_{5},v_{2}v_{6}\}$. Otherwise, both $v_{1}$ and $v_{6}$ are $4$-vertices; with a $5$-neighbour $v_{2}$, we may take $G^{\prime}=G-v+\{v_{1}v_{6},v_{2}v_{4},v_{2}v_{6}\}$. In both cases, $G^{\prime}$ is proper with respect to $G$, a contradiction.
$(c)$ Let $n_{4}(v)=2$, and suppose that $v$ has two bad $5$-neighbours. Each bad $5$-neighbour $v_{i}$ forces either $v_{i}v_{i-1}$ or $v_{i}v_{i+1}$ to lie in two $3$-faces. If $v$ had at most one $6(4^{-})$-neighbour, then the last neighbour of $v$ would be either a $5$- or $6(5)$- or $6(6)$-vertex, giving $d_{2}(v)\leq 18$. Applying the same reductions as in part (b) yields a contradiction. Thus $v$ has at least two $6(4^{-})$-neighbours.
$(d)$ Let $n_{4}(v)=1$. If $v$ had four bad $5$-neighbours, then by Lemma 2.2 we would have $d_{2}(v)\leq 18$. If one of $v_{2},v_{3},v_{4},v_{5}$ is a $4$-vertex (say $v_{2}$), use $G^{\prime}=G-v+\{v_{2}v_{4},v_{2}v_{5},v_{2}v_{6}\}$. Otherwise the $4$-vertex is $v_{1}$ or $v_{6}$; assume $v_{1}$. Since $v$ has four $5$-neighbours, one of $v_{3},v_{5}$ is $5$; assume $v_{3}$. Then $G^{\prime}=G-v+\{v_{1}v_{6},v_{3}v_{1},v_{3}v_{5}\}$ is proper with respect to $G$, a contradiction.
$(e)$ Let $n_{4}(v)=1$ and suppose that $v$ has three bad $5$-neighbours. Observe that two distinct edges among $v_{i}v_{i+1}$ and $v_{j}v_{j+1}$ are contained in two $3$-faces. If $v$ had no $6(4^{-})$-neighbour, then its remaining neighbours would be $5$-, $6(5)$-, or $6(6)$-vertices, giving $d_{2}(v)\leq 18$ by Lemma 2.2. If one of $v_{2},v_{3},v_{4},v_{5}$ is a $4$-vertex (say $v_{2}$), then define $G^{\prime}=G-v+\{v_{2}v_{4},v_{2}v_{5},v_{2}v_{6}\}$. Otherwise, the $4$-vertex must be $v_{1}$ or $v_{6}$; assume $v_{1}$. If $v_{2},v_{4},v_{6}$ are $5$-vertices, then we set $G^{\prime}=G-v+\{v_{1}v_{6},v_{2}v_{4},v_{2}v_{6}\}$. Else, one of $v_{3},v_{5}$ is a $5$-vertex; assume $v_{3}$. Then $G^{\prime}=G-v+\{v_{1}v_{6},v_{3}v_{1},v_{3}v_{5}\}$. In each case, $G^{\prime}$ is proper with respect to $G$, contradicting Remark 2.1. Therefore, $v$ must have a $6(4^{-})$-neighbour.
$(f)$ Let $n_{4}(v)=0$. If $v$ had five bad $5$-neighbours, then there would exist three distinct edges of $E(v)$ such that each of them is contained in two $3$-faces, giving $d_{2}(v)\leq 18$ by Lemma 2.2. Among $v_{1},v_{6}$, one is $5$-vertex; assume $v_{1}$. Among $v_{3},v_{5}$, one is $5$-vertex; assume $v_{3}$. Then $G^{\prime}=G-v+\{v_{1}v_{6},v_{3}v_{1},v_{3}v_{5}\}$ is proper with respect to $G$, a contradiction. Thus $v$ has at most four bad $5$-neighbours.
$(g)$ Let $n_{4}(v)=0$, $m_{4}(v)=1$, and suppose that $v$ has four bad $5$-neighbours. Observe that two non-adjacent edges among $v_{i}v_{i+1}$ is contained in two $3$-faces. If $v$ had at most one $6(4^{-})$-neighbour, then its remaining neighbour would be a $5(3^{-})$-vertex or a $6(5^{+})$-vertex, giving $d_{2}(v)\leq 18$ by Lemma 2.2. This implies either $v$ has five $5$-neighbours or three edges around $v$ lie in two $3$-faces.
If $v$ has five $5$-neighbours, the argument from part (f) applies, giving a contradiction. Thus $v$ has exactly four $5$-neighbours. If $v_{1}$ (resp. $v_{6}$) is a $5$-vertex and one of $v_{3},v_{5}$ (resp. $v_{2},v_{4}$) is a $5$-vertex, we again use the same reduction as in (f). Therefore we may assume $v_{1},v_{6}$ are $6$-vertices and $v_{2},v_{3},v_{4},v_{5}$ are bad $5$-vertices. Since a $5(5)$-vertex cannot have any bad $5$-neighbour by Lemma 2.14, each of $v_{2},v_{3},v_{4},v_{5}$ must be a $5(4)$-vertex. However, by Proposition 2.13, a $5(4)$-vertex cannot have two non-adjacent $5(4)$-neighbours. This contradicts planarity of $G$, and therefore $v$ must have at least two $6(4^{-})$-neighbours. ∎
We now apply discharging method to show that $G$ cannot exist. First, we assign to each vertex $v$ a charge $\mu(v)=d(v)-4$ and to each face $f$ a charge $\mu(f)=\ell(f)-4$. By Euler’s formula, we have
$\sum_{v\in V(G)}\left(d(v)-4\right)+\sum_{f\in F(G)}(\ell(f)-4)=-8.$
We next present some rules and redistribute accordingly. Once the discharging finishes, we check the final charge $\mu^{*}(v)$ and $\mu^{*}(f)$. If $\mu^{*}(v)\geq 0$ and $\mu^{*}(f)\geq 0$, we get a contradiction that no such a counterexample can exist.
Discharging Rules
We apply the following discharging rules.
  • R1:
    Every $3$-face receives $\frac{1}{3}$ from each of its incident vertices.
  • R2:
    Let $f$ be a $5^{+}$-face. Then $f$ gives
    • $(a)$
      $\frac{1}{3}$ to each of its incident $3$-vertices and bad $4$-vertices,
    • $(b)$
      $\frac{1}{5}$ to each of its incident $5(4)$-vertices,
    • $(c)$
      $\frac{1}{5}$ to each of its incident $6(5)$-vertex having no $4$-neighbour.
  • R3:
    Every $6(3^{-})$-vertex gives $\frac{1}{6}$ to each of its neighbours.
  • R4:
    Let $v$ be a $6(4)$-vertex. Then $v$ gives
    • $(a)$
      $\frac{1}{12}$ to each of its $4$- and $6(5)$-neighbours,
    • $(b)$
      $\frac{1}{9}$ to each of its $3$- and $5(4)$-neighbours,
    • $(c)$
      $\frac{1}{6}$ to each of its $5(5)$-neighbour $u$ if $uv$ is a special edge,
    • $(d)$
      $\frac{1}{9}$ to each of its $5(5)$-neighbour $u$ if $uv$ is not a special edge.
  • R5:
    Let $v$ be a $6(5)$-vertex. Then $v$ gives
    • $(a)$
      $\frac{1}{12}$ to each of its bad $4$-neighbours,
    • $(b)$
      $\frac{1}{9}$ to each of its bad $5$-neighbours.
Checking $\mu^{*}(v),\mu^{*}(f)\geq 0$ for $v\in V(G),f\in F(G)$.
First we show that $\mu^{*}(f)\geq 0$ for each $f\in F(G)$. Given a face $f\in F(G)$, if $f$ is a $3$-face, then it receives $\frac{1}{3}$ from each of incident vertices by R1, and so $\mu^{*}(f)=-1+3\times\frac{1}{3}=0$. If $f$ is a $4$-face, then $\mu(f)=\mu^{*}(f)=0$.
Let $f$ be a $5$-face. Suppose first that $f$ is not incident to any $3$-vertex. Note that if $f$ is not incident to any bad $4$-vertices, then $\mu^{*}(v)\geq 1-5\times\frac{1}{5}=0$ after $f$ sends $\frac{1}{5}$ to each of its incident vertices by R2(b),(c). Therefore we further suppose that $f$ is incident to at least one bad $4$-vertex, let $u$ be such a bad $4$-vertex incident to $f$. If $u$ is a $4(1)$-vertex, then it has at most one neighbour consisting of $4$- or $5(4)$-vertex by Lemma 2.6 and Corollary 2.11. Besides, if $u$ is a $4(2)$-vertex, then the neighbours of $u$ are neither $4$-vertex nor $5(4)$-vertex by Lemma 2.9. As a result, $f$ is incident to at most three bad $4$-vertices. On the other hand, a $5(4)$-vertex having a $4$-neighbour cannot have any other $4$- or $5(4)$-neighbour by Lemma 2.10. Thus, we conclude the following.
  • If $f$ is incident to three bad $4$-vertices, then the remaining vertices incident to $f$ are neither a $4$-vertex, nor a $5(4)$-vertex, nor a $6(5)$-vertex having no $4$-neighbour.
  • If $f$ is incident to exactly two bad $4$-vertices, then $f$ is incident to at most one vertex that is either a $5(4)$-vertex or a $6(5)$-vertex having no $4$-neighbour.
  • If $f$ is incident to exactly one bad $4$-vertex, say $w$, then one of the neighbors of $w$ lying on $f$ is neither a $4$-vertex, nor a $5(4)$-vertex, nor a $6(5)$-vertex having no $4$-neighbour.
In each case, we deduce that $\mu^{*}(f)\geq 0$, since the face $f$ sends $\frac{1}{3}$ to each of its incident bad $4$-vertices by R2(a), $\frac{1}{5}$ to each of its incident $5(4)$-vertices by R2(b) and $\frac{1}{5}$ to each of its incident $6(5)$-vertices having no $4$-neighbour by R2(c).
Now we suppose that $f$ is incident to a $3$-vertex, say $v_{1}$. Then the neighbours of $v_{1}$ lying on $f$, denote by $v_{2},v_{5}$, are different from $5^{-}$- and $6(5^{+})$-vertex by Lemma 2.4. Thus, $\mu^{*}(v)\geq 1-3\times\frac{1}{3}=0$ after $f$ sends at most $\frac{1}{3}$ to each of its incident vertices other than $v_{2},v_{5}$ by R2.
On the other hand, if $f$ is a $6^{+}$-face, then we have $\mu^{*}(v)\geq\ell(f)-4-\ell(f)\times\frac{1}{3}\geq 0$ after $v$ sends at most $\frac{1}{3}$ to each of its incident vertices by R2.
We now pick a vertex $v\in V(G)$ with $d(v)=k$. By Lemma 2.3, we have $k\geq 3$.
(1). Let $k=3$. By Lemma 2.4, $m_{3}(v)=0$, $m_{4}(v)\leq 1$, and each neighbour of $v$ is a $6(4^{-})$-vertex. Note that $v$ is incident to at least two $5^{+}$-faces, and each of them gives $\frac{1}{3}$ to $v$ by R2(a). On the other hand, $v$ is incident to three $6(4^{-})$-vertices, and each of them gives at least $\frac{1}{9}$ to $v$ by R3 and R4(b). Thus, we have $\mu^{*}(v)\geq-1+2\times\frac{1}{3}+3\times\frac{1}{9}=0$.
(2). Let $k=4$. By Lemma 2.5, $v$ is incident to at most two $3$-faces, in particular, $v$ has no $6(6)$-neighbour. Clearly, we have $\mu^{*}(v)\geq 0$ when $v$ is not incident to any $3$-face, since $v$ does not give a charge to any vertex or face. We may therefore assume that $1\leq m_{3}(v)\leq 2$; namely, $v$ is a bad $4$-vertex.
First suppose that $m_{3}(v)=1$. Let $m_{4}(v)=r$ for $0\leq r\leq 3$. By Lemma 2.7, we have $n_{6}(v)\geq r+1$, i.e., $v$ has at least $r+1$ many $6(5^{-})$-neighbours, and each of them gives at least $\frac{1}{12}$ to $v$ by R3, R4(a) and R5(a). On the other hand, $v$ is incident to 3r3-r many $5^{+}$-faces, and each of them gives $\frac{1}{3}$ to $v$ by R2(a). Consequently, $v$ receives totally $(3-r)\times\frac{1}{3}$ from its incident $5^{+}$-faces, and totally $(r+1)\times\frac{1}{12}$ from its $6(5^{-})$-neighbours. Thus, $\mu^{*}(v)\geq(3-r)\times\frac{1}{3}+(r+1)\times\frac{1}{12}-\frac{1}{3}\geq 0$ after $v$ sends $\frac{1}{3}$ to its incident $3$-face by R1.
We now suppose that $m_{3}(v)=2$. Then, by Lemma 2.8, we have $m_{4}(v)=r\leq 1$ and $n_{6}(v)\geq r+3$. Since $v$ is not adjacent to any $6(6)$-vertex, $v$ has at least $r+3$ many $6(5^{-})$-neighbours, and each of them gives at least $\frac{1}{12}$ to $v$ by R3, R4(a) and R5(a). On the other hand, $v$ is incident to 2r2-r many $5^{+}$-faces, and each of them gives $\frac{1}{3}$ to $v$ by R2(a). Consequently, $v$ receives totally $(2-r)\times\frac{1}{3}$ from its incident $5^{+}$-faces, and totally $(r+3)\times\frac{1}{12}$ from its $6$-neighbours. Thus, $\mu^{*}(v)\geq(2-r)\times\frac{1}{3}+(r+3)\times\frac{1}{12}-2\times\frac{1}{3}\geq 0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
(3). Let $k=5$. We first note that if $m_{3}(v)\leq 3$, then $\mu^{*}(v)\geq 1-3\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1. Therefore, we may assume that $m_{3}(v)\geq 4$; namely, $v$ is a bad $5$-vertex.
Let $m_{3}(v)=4$. If $m_{4}(v)=0$, then $v$ has at least two $6(5^{-})$-neighbours by Lemma 2.12(a), and $v$ receives at least $\frac{1}{9}$ from each of its $6(5^{-})$-neighbours by R3, R4(b) and R5(b). Also, $v$ receives $\frac{1}{5}$ from its incident $5^{+}$-face by R2(b). Thus, $\mu^{*}(v)\geq 1+2\times\frac{1}{9}+\frac{1}{5}-4\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1. Suppose now that $m_{4}(v)=1$. Then $v$ has at least three $6(5^{-})$-neighbours by Lemma 2.12(b). It follows from applying R3, R4(b) and R5(b) that $v$ receives at least $\frac{1}{9}$ from each of its $6(5^{-})$-neighbours. Thus, $\mu^{*}(v)\geq 1+3\times\frac{1}{9}-4\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
Let $m_{3}(v)=5$. Recall that $v$ has no $6(6)$-neighbour by Lemma 2.14. Suppose first that there exists $v_{i}v_{i+1}\in E(v)$ such that $v_{i}v_{i+1}$ is contained in two $3$-faces. It then follows from Proposition 2.16(a) that all neighbours of $v$ are $6(5^{-})$-vertices, and each $vv_{j}$ for $v_{j}\in N(v)\setminus\{v_{i},v_{i+1}\}$ is a special edge. This particularly implies that each $v_{j}\in N(v)\setminus\{v_{i},v_{i+1}\}$ is a $6(4^{-})$-vertex. Then, $v$ receives $\frac{1}{6}$ from each $v_{j}$ for $j\in[5]\setminus\{i,i+1\}$ by R3 and R4(c), and at least $\frac{1}{9}$ from each of $v_{i},v_{i+1}$ by R3, R4(d) and R5(b). Thus, $v$ totally receives at least $\frac{2}{3}$ from its $6$-neighbours. Next we suppose that no edge $v_{i}v_{i+1}\in E(v)$ is contained in two $3$-faces. By Proposition 2.16(b), there exist four $6(4^{-})$-vertices $v_{1},v_{2},v_{3},v_{4}\in N(v)$ such that each $vv_{i}$ is a special edge. Then, $v$ receives $\frac{1}{6}$ from each $v_{i}$ for $i\in[4]$ by R3 and R4(c). Clearly, $v$ receives totally at least $\frac{2}{3}$ from its $6$-neighbours. Hence, $\mu^{*}(v)\geq 1+\frac{2}{3}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
(4). Let $k=6$. Notice first that if $m_{3}(v)\leq 3$, then $\mu^{*}(v)\geq 2-3\times\frac{1}{3}-6\times\frac{1}{6}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and at most $\frac{1}{6}$ to each of its neighbours by R3. In addition, if $m_{3}(v)=6$, then $\mu^{*}(v)\geq 2-6\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1. Therefore we may assume that $4\leq m_{3}(v)\leq 5$.
(4.1). Let $m_{3}(v)=4$. Obviously, $v$ has one of the three configurations depicted in Figure 2. Notice that $v$ only gives charge to its incident $3$-faces, and $4^{-}$-, $5(4^{+})$-, $6(5)$-neighbours by R1, R4. We first note that if $v$ has no $5(5)$-neighbour $v_{i}$ such that $vv_{i}$ is a special edge, then $v$ gives at most $\frac{1}{9}$ to each of its neighbours by R4(a),(b),(d), and so $\mu^{*}(v)\geq 2-6\times\frac{1}{9}-4\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1. We may further assume that $v$ has at least one $5(5)$-neighbour $v_{i}$ such that $vv_{i}$ is a special edge. In fact, we will determine the final charge of $v$ based on the number of $5(5)$-neighbors $v_{i}$ of $v$ for which each edge $vv_{i}$ is a special edge. It can be easily observed that $v$ cannot have three $5(5)$-neighbours by Lemma 2.14.
[…]
Figure 2. Three configurations of $6(4)$-vertices.
Suppose first that $v$ has two $5(5)$-neighbours $v_{i}$ and $v_{j}$ such that both $vv_{i}$ and $vv_{j}$ are special edges. By Lemma 2.14, $v_{i}$ and $v_{j}$ are non-adjacent, so the configuration in Figure 2(b) is not possible for $v$. Moreover, $v$ cannot have the form in Figure 2(c) as well, since $vv_{i}$ and $vv_{j}$ are special edges. Hence, we deduce that the neighbours of $v$ can only form as depicted in Figure 2(a), where $v_{i}=v_{2}$ and $v_{j}=v_{5}$. By Lemmas 2.4 and 2.14, $v_{2}$ and $v_{5}$ cannot have a $4^{-}$-neighbour, bad $5$-neighbour and $6(6)$-neighbour, so the other neighbours of $v$ are $6(5^{-})$-vertices or $5(3^{-})$-vertices. Thus, $v$ gives $\frac{1}{6}$ to each of $v_{2},v_{5}$ by R4(c), and at most $\frac{1}{12}$ to each of the other neighbours of $v$ by R4(a). Consequently, $\mu^{*}(v)\geq 2-2\times\frac{1}{6}-4\times\frac{1}{12}-4\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
Now we suppose that $v$ has exactly one $5(5)$-neighbour $v_{i}$ such that $vv_{i}$ is a special edge. Then, similarly as above, we deduce that the neighbours of $v$ can only form as depicted in Figure 2(a) where $v_{i}=v_{2}$ (or $v_{i}=v_{5}$). By Lemmas 2.4 and 2.14, $v_{2}$ cannot have a $4^{-}$-neighbour, bad $5$-neighbour and $6(6)$-neighbour, so the neighbours of $v_{2}$ other than $v$ are $6(5^{-})$-vertices or $5(3^{-})$-vertices. Thus, $v$ gives $\frac{1}{6}$ to $v_{2}$, at most $\frac{1}{12}$ to each of $v_{1},v_{3}$ by R4(a), at most $\frac{1}{9}$ to each of $v_{4},v_{5},v_{6}$ by R4(a),(b),(d). Hence, $\mu^{*}(v)\geq 2-\frac{1}{6}-2\times\frac{1}{12}-3\times\frac{1}{9}-4\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
(4.2). Let $m_{3}(v)=5$. By Lemma 2.17, we have $n_{4}(v)\leq 3$.
Let $n_{4}(v)=3$. Note that $v$ has no bad $5$-neighbours by Lemma 2.17(a). So, $v$ only gives charge to its bad $4$-neighbours and its incident $3$-faces. Thus, $\mu^{*}(v)\geq 2-5\times\frac{1}{3}-3\times\frac{1}{12}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and $\frac{1}{12}$ to each of its bad $4$-neighbours by R5(a).
Let $n_{4}(v)=2$. By Lemma 2.17(b), $v$ has at most two bad $5$-neighbours. If $v$ has such two neighbours, then the remaining neighbours of $v$ are $6(4^{-})$-vertices by Lemma 2.17(c), i.e., $v$ has two $6(4^{-})$-neighbours. In such a case, $v$ gives $\frac{1}{12}$ to each of its bad $4$-neighbours by R5(a), and $\frac{1}{9}$ to each of its bad $5$-neighbours by R5(b). On the other hand, $v$ receives at least $\frac{1}{12}$ from each of its $6(4^{-})$-neighbours by R3 and R4(a). Thus, $\mu^{*}(v)\geq 2-2\times\frac{1}{12}-2\times\frac{1}{9}+2\times\frac{1}{12}-5\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1. Suppose now that $v$ has at most one bad $5$-neighbour. Then, $v$ gives $\frac{1}{12}$ to each of its bad $4$-neighbours by R5(a), and $\frac{1}{9}$ to its bad $5$-neighbour R5(b) (if exists). Thus, $\mu^{*}(v)\geq 2-2\times\frac{1}{12}-\frac{1}{9}-5\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
Let $n_{4}(v)=1$. By Lemma 2.17(d), $v$ has at most three bad $5$-neighbours. If $v$ has such three neighbours, then $v$ has a $6(4^{-})$-neighbour by Lemma 2.17(e), and so $v$ receives at least $\frac{1}{12}$ from its $6(4^{-})$-neighbour by R3 and R4(a). On the other hand, $v$ gives $\frac{1}{12}$ to its bad $4$-neighbour by R5(a), and $\frac{1}{9}$ to each of its bad $5$-neighbours by R5(b). Thus, $\mu^{*}(v)\geq 2+\frac{1}{12}-\frac{1}{12}-3\times\frac{1}{9}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1. Suppose now that $v$ has at most two bad $5$-vertices. Then, $\mu^{*}(v)\geq 2-\frac{1}{12}-2\times\frac{1}{9}-5\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{12}$ to its bad $4$-neighbour R5(a), $\frac{1}{9}$ to each of its bad $5$-neighbours R5(b), and $\frac{1}{3}$ to each of its incident $3$-faces by R1.
Let $n_{4}(v)=0$. Notice first that $v$ has at most four bad $5$-neighbours by Lemma 2.17(f). If $v$ has at most three bad $5$-neighbours, then $v$ gives $\frac{1}{9}$ to each of its bad $5$-neighbours by R5(b), and so $\mu^{*}(v)\geq 2-3\times\frac{1}{9}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1. Assume further that $v$ has exactly four bad $5$-neighbours. If $m_{4}(v)=0$, i.e, $v$ is incident to a $5^{+}$-face, then $v$ receives $\frac{1}{5}$ from its incident $5^{+}$-face by R2(c). Thus, $\mu^{*}(v)\geq 2+\frac{1}{5}-4\times\frac{1}{9}-5\times\frac{1}{3}>0$ after $v$ gives $\frac{1}{9}$ to each of its bad $5$-neighbours by R5(b), and $\frac{1}{3}$ to each of its incident $3$-faces by R1. If $m_{4}(v)=1$, then $v$ has two $6(4^{-})$-neighbours by Lemma 2.17(g), and so $v$ receives at least $\frac{1}{12}$ from each of its $6(4^{-})$-neighbours by R3 and R4(a). Thus, $\mu^{*}(v)\geq 2+2\times\frac{1}{12}-4\times\frac{1}{9}-5\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{9}$ to each of its bad $5$-neighbours by R5(b), and $\frac{1}{3}$ to each of its incident $3$-faces by R1.
2.2. The case $\Delta=7$
Recall that $G$ does not admit any $2$-distance 2121-coloring, whereas any planar graph $G^{\prime}$ obtained from $G$ with a smaller value of $|V(G^{\prime})|+|E(G^{\prime})|$ admits a $2$-distance 2121-coloring. We begin by establishing several structural properties of $G$, similar to those in the case $\Delta=6$.
The proof of Lemma 2.18 is omitted, as it is similar to the proof of Lemma 2.3.
Lemma 2.18.
$\delta(G)\geq 3$.
Lemma 2.19.
If $v$ is a $3$-vertex, then $m_{3}(v)=0$, $m_{4}(v)\leq 1$, and each neighbour of $v$ is a $7(5^{-})$-vertex.
Proof.
Let $v$ be a $3$-vertex. Assume first that $v$ is incident to a $3$-face, say $f_{1}=v_{1}vv_{2}$. By Lemma 2.2, we have $d_{2}(v)\leq 19$. Setting $G^{\prime}=G-v+\{v_{1}v_{3}\}$ yields a graph that is proper with respect to $G$, contradicting Remark 2.1. Thus $m_{3}(v)=0$.
Next, suppose that $m_{4}(v)\geq 2$, i.e., $v$ is incident to two $4$-faces, say $f_{1}=v_{1}vv_{2}x$ and $f_{2}=v_{2}vv_{3}y$. Again, Lemma 2.2 gives $d_{2}(v)\leq 19$. The graph $G^{\prime}=G-v+\{v_{1}v_{3}\}$ is proper with respect to $G$, giving the same contradiction. Hence $m_{4}(v)\leq 1$.
Finally, we show that each neighbour of $v$ must be a $7(5^{-})$-vertex. Since $m_{3}(v)=0$, the vertex $v$ cannot have a $7(6^{+})$-neighbour. Moreover, $v$ cannot have any $6^{-}$-neighbour; otherwise, if $v_{1}$ is such a neighbour, then considering $G^{\prime}=G-v+\{v_{1}v_{2},v_{1}v_{3}\}$ gives $d_{2}(v)\leq 20$ by Lemma 2.2, and $G^{\prime}$ is proper with respect to $G$, again a contradiction. Thus all neighbours of $v$ are $7(5^{-})$-vertices. ∎
Lemma 2.20.
If $v$ is a $4$-vertex, then $m_{3}(v)\leq 3$. In particular, $G$ has no $4(4)$-vertex.
Proof.
Let $v$ be a $4$-vertex. Assume that $m_{3}(v)=4$. Then $d_{2}(v)\leq 20$ by Lemma 2.2. If we take $G^{\prime}=G-v$, then $G^{\prime}$ is proper with respect to $G$, contradicting Remark 2.1. Thus $m_{3}(v)\leq 3$. ∎
We now introduce some vertices of special types: a $4(1)$-, $4(2)$-, or $4(3)$-vertex is called a bad $4$-vertex, and a $5(4)$- or $5(5)$-vertex is called a bad $5$-vertex.
Lemma 2.21.
Let $v$ be a $4$-vertex.
  • $(a)$
    $v$ has no $5(5)$-neighbour.
  • $(b)$
    If $m_{3}(v)\geq 1$, then $v$ has at most one $4$-neighbour.
  • $(c)$
    If $m_{3}(v)\geq 1$ and $v$ has a $4$-neighbour, then $v$ cannot have any $5(4)$-neighbour.
Proof.
$(a)$ Suppose, for a contradiction, that $v$ has a $5(5)$-neighbour, say $v_{2}$. We may therefore assume that each of $v_{1}v_{2},v_{2}v_{3}$ is contained in two $3$-faces. By Lemma 2.2, we have $d_{2}(v)\leq 20$. Setting $G^{\prime}=G-v+\{v_{2}v_{4}\}$ yields a graph that is proper with respect to $G$, contradicting Remark 2.1.
$(b)$ Let $m_{3}(v)\geq 1$. Suppose that $v$ has two $4$-neighbours, and let $v_{1}$ be one of them. Then $d_{2}(v)\leq 20$ by Lemma 2.2. Taking $G^{\prime}=G-v+\{v_{1}v_{2},v_{1}v_{3},v_{1}v_{4}\}$ gives a graph that is proper with respect to $G$, a contradiction.
$(c)$ Let $m_{3}(v)\geq 1$. Suppose that $v$ has both a $4$-neighbour and a $5(4)$-neighbour. Then Lemma 2.2 again gives $d_{2}(v)\leq 20$, and the same reduction as in part (b) contradicts Remark 2.1. ∎
Lemma 2.22.
Let $v$ be a $4(1)$-vertex.
  • $(a)$
    If $m_{4}(v)=2$, then $v$ has a $7$-neighbour. In particular, $v$ has either a $7(5^{-})$-neighbour or two $7(6)$-neighbours.
  • $(b)$
    If $m_{4}(v)=3$, then $v$ has two $7(6^{-})$-neighbours.
Proof.
Let $f_{1}=v_{1}vv_{2}$.
$(a)$ Suppose $m_{4}(v)=2$. We first show that $v$ has a $7$-neighbour. Assume all neighbours of $v$ are $6^{-}$-vertices. Then $d_{2}(v)\leq 20$ by Lemma 2.2. Setting $G^{\prime}=G-v+\{v_{2}v_{3},v_{3}v_{4},v_{1}v_{4}\}$ gives a graph that is proper with respect to $G$, contradicting Remark 2.1. Thus $v$ has at least one $7$-neighbour.
Now assume $v$ has no $7(5^{-})$-neighbour. Since $m_{3}(v)=1$, $v$ cannot have any $7(7)$-neighbour, so it must have a $7(6)$-neighbour. If the remaining neighbours of $v$ are $6^{-}$-vertices, then again $d_{2}(v)\leq 20$ by Lemma 2.2, and the same replacement as above yields a graph that is proper with respect to $G$. Thus $v$ must have a second $7(6)$-neighbour.
$(b)$ Suppose $m_{4}(v)=3$. We first show that $v$ has two $7$-neighbours. Assume three neighbours of $v$ are $6^{-}$-vertices. Then $d_{2}(v)\leq 20$ by Lemma 2.2. Setting $G^{\prime}=G-v+\{v_{2}v_{3},v_{1}v_{4}\}$ produces a graph that is proper with respect to $G$, contradicting Remark 2.1. Thus $v$ has two $7$-neighbours, and since it cannot have any $7(7)$-neighbour, these neighbours must be $7(6^{-})$-vertices. ∎
Lemma 2.23.
Let $v$ be a $4(2)$-vertex.
  • $(a)$
    If $m_{4}(v)=0$, then $v$ has either a $7(5^{-})$-neighbour and a $6(4^{-})$-neighbour, or two $7(6^{-})$-neighbours.
  • $(b)$
    If $m_{4}(v)=1$, then $v$ has either two $7(6^{-})$-neighbours and two $6(4^{-})$-neighbours, or three $7(6^{-})$-neighbours.
  • $(c)$
    If $m_{4}(v)=2$, then $v$ has either two $7(5^{-})$-neighbours and two $7(6)$-neighbours, or three $7(5^{-})$-neighbours.
Proof.
Let $f_{i}=v_{i}vv_{i+1}$ and $f_{j}=v_{j}vv_{j+1}$ be the two $3$-faces incident to $v$, with $i<j$ (see Figure 3).
Whenever $d_{2}(v)\leq 20$, we reduce $G$ as follows: If $(i,j)=(1,2)$, we set $G^{\prime}=G-v+\{v_{2}v_{4}\}$. If $(i,j)=(1,3)$, we set $G^{\prime}=G-v+\{v_{2}v_{3},v_{1}v_{4}\}$. In each case, $G^{\prime}$ is proper with respect to $G$, contradicting Remark 2.1. Therefore, throughout the proof we may assume $d_{2}(v)\geq 21$.
$(a)$ Let $m_{4}(v)=0$. Suppose first that $v$ has a $7(7)$-neighbour. Then $v$ cannot have two $6^{-}$-neighbours; otherwise $d_{2}(v)\leq 20$ by Lemma 2.2, a contradiction. Thus $v$ has two $7(6^{-})$-neighbours.
Now suppose $v$ has no $7(7)$-neighbour. Since $v$ cannot have all neighbours of degree at most $6$, it follows that $v$ has at least one $7$-neighbour; let it be $v_{r}$. Assume that $v$ has exactly one $7(6^{-})$-neighbour (otherwise the conclusion already holds). Then all other neighbours of $v$ are $6^{-}$-vertices. If $v_{r}$ were a $7(6)$-vertex, or if $v$ had any $6(5^{+})$-neighbour, then again $d_{2}(v)\leq 20$ by Lemma 2.2, a contradiction. Hence $v_{r}$ must be a $7(5^{-})$-vertex, and $v$ must have a $6(4^{-})$-neighbour.
[…]
(a) $i=1$ and $j=2$
[…]
(b) $i=1$ and $j=3$
Figure 3. Two possible 33-faces $f_{i}=v_{i}vv_{i+1}$ and $f_{j}=v_{j}vv_{j+1}$ incident to a $4(2)$-vertex.
$(b)$ Let $m_{4}(v)=1$. We first claim that $v$ has at least two $7$-neighbours. Indeed, if three neighbours of $v$ were $6^{-}$-vertices, then $d_{2}(v)\leq 20$ (since $m_{4}(v)=1$), a contradiction. Note also that $v$ has at most one $7(7)$-neighbour, because $m_{3}(v)=2$. We now distinguish cases depending on how many of the edges $v_{i}v_{i+1}$ and $v_{j}v_{j+1}$ lie in two $3$-faces.
Suppose that both $v_{i}v_{i+1}$ and $v_{j}v_{j+1}$ lie in two $3$-faces. Then $v$ cannot have any $6^{-}$-neighbour (otherwise $d_{2}(v)\leq 20$). Hence all its neighbours are $7$-vertices, and thus $v$ has three $7(6^{-})$-neighbours.
Suppose next that exactly one of $v_{i}v_{i+1}$, $v_{j}v_{j+1}$ lies in two $3$-faces. Then $v$ cannot have any $7(7)$- or $6(6)$-neighbour, so it has exactly two $7(6^{-})$-neighbours, say $v_{1}$ and $v_{2}$. Consider $v_{3}$ and $v_{4}$. If one is a $5^{-}$-vertex, or if both are $6$-vertices, then $d_{2}(v)\leq 20$, a contradiction. Thus one of $v_{3},v_{4}$ is a $7(6^{-})$-vertex and the other is a $6^{+}$-vertex. Hence $v$ has at least three $7(6^{-})$-neighbours.
Finally suppose that neither $v_{i}v_{i+1}$ nor $v_{j}v_{j+1}$ lies in two $3$-faces. Then $v$ has no $7(7)$-, $7(6)$-, $6(6)$-, or $6(5)$-neighbour. Thus $v$ has two $7(5^{-})$-neighbours, say $v_{1}$ and $v_{2}$. For $v_{3},v_{4}$: If one is a $5^{-}$-vertex, then the other must be a $7(5^{-})$-vertex (else $d_{2}(v)\leq 20$). If neither is a $5^{-}$-vertex, then either one is a $7(5^{-})$-vertex or both are $6(4^{-})$-vertices. This completes part (b).
$(c)$ Let $m_{4}(v)=2$. If $v$ had two $6^{-}$-neighbours, or a $7(7)$-neighbour, or a $6(6)$-neighbour, then $d_{2}(v)\leq 20$ (since $m_{4}(v)=2$), a contradiction. Thus $v$ has exactly three $7(6^{-})$-neighbours.
If $v$ had three $7(6)$-neighbours, then both $v_{i}v_{i+1}$ and $v_{j}v_{j+1}$ would have to lie in two $3$-faces, forcing again $d_{2}(v)\leq 20$, a contradiction. Thus $v$ has at most two $7(6)$-neighbours.
If $v$ has a $7(6)$-neighbour, then at least one of $v_{i}v_{i+1}$ or $v_{j}v_{j+1}$ lies in two $3$-faces, say $v_{i}v_{i+1}$. This forces that $v_{i+1}$ is a $7(6^{-})$-vertex, and all neighbours of $v$ other than $v_{i},v_{i+1}$ are $7(5^{-})$-vertices (otherwise $d_{2}(v)\leq 20$). Hence $v$ has either: two $7(6)$-neighbours and two $7(5^{-})$-neighbours, or three $7(5^{-})$-neighbours. ∎
Lemma 2.24.
Let $v$ be a $4(3)$-vertex.
  • $(a)$
    If $m_{4}(v)=0$, then $v$ has one of the following: four $7(5^{-})$-neighbours, or three $7(5^{-})$- and one $7(6)$-neighbours, or three $7(5^{-})$- and one $6(4^{-})$-neighbours, or two $7(5^{-})$- and two $7(6)$-neighbours.
  • $(b)$
    If $m_{4}(v)=1$, then all neighbours of $v$ are $7(5^{-})$-vertices.
Proof.
Let $f_{i}=v_{i}vv_{i+1}$ for $i\in[3]$. If $d_{2}(v)\leq 20$, then $G^{\prime}=G-v+\{v_{1}v_{4}\}$ yields a graph that is proper with respect to $G$, contradicting Remark 2.1. Therefore, throughout the proof we may assume $d_{2}(v)\geq 21$. Note that at most one of the edges $v_{1}v_{2}$, $v_{2}v_{3}$, $v_{3}v_{4}$ is contained in two $3$-faces, and so $v$ has no $6(6)$- or $7(7)$-neighbour. Observe also that $v$ has no $5^{-}$- or $6(5)$-neighbour; $v$ has at most one $6$-neighbour; and $v$ has at most two $7(6)$-neighbours.
$(a)$ Let $m_{4}(v)=0$. As noted above, $v$ has at most one $6(4^{-})$-neighbour. If $v$ has such a $6(4^{-})$-neighbour, then it cannot have a $7(6)$-neighbour, since otherwise $d_{2}(v)\leq 20$, a contradiction. Thus all other neighbours of $v$ are $7(5^{-})$-vertices. If $v$ has no $6(4^{-})$-neighbour, then the only possible neighbourhood types are exactly those listed in the statement, since $v$ has at most two $7(6)$-neighbours.
$(b)$ Let $m_{4}(v)=1$. If $v$ has a $7(6^{+})$- or a $6$-neighbour, then $d_{2}(v)\leq 20$ by Lemma 2.2, a contradiction. Hence every neighbour of $v$ must be a $7(5^{-})$-vertex. ∎
Lemma 2.25.
Let $v$ be a $5(4)$-vertex.
  • $(a)$
    $v$ cannot have two $5(5)$-neighbours.
  • $(b)$
    If $v$ has a $5(5)$-neighbour, then $v$ has no $4$-neighbour.
  • $(c)$
    If $v$ has a $7(6)$-neighbour, then $v$ is adjacent to at most one $4$-vertex.
  • $(d)$
    If $m_{4}(v)=0$, then $v$ has at least two neighbours that are neither $5^{-}$- nor $6(6)$-vertices.
  • $(e)$
    If $m_{4}(v)=1$, then $v$ has at least three $6^{+}$-neighbours different from $6(6)$-vertex.
  • $(f)$
    If $m_{4}(v)=1$ and $v$ has a $7(7)$-neighbour, then $v$ has at least four $6^{+}$-neighbours different from $6(6)$-vertex.
Proof.
Let $f_{i}=v_{i}vv_{i+1}$ for $i\in[4]$. If $d_{2}(v)\leq 20$, then setting $G^{\prime}=G-v+\{v_{4}v_{5}\}$ produces a graph $G^{\prime}$ that is proper with respect to $G$, contradicting Remark 2.1. Thus throughout the proof we assume $d_{2}(v)\geq 21$.
$(a)$ If $v$ had two $5(5)$-neighbours, then there would exist at least three edges in $E(v)$ such that each of them lies on two $3$-faces, which forces $d_{2}(v)\leq 20$ by Lemma 2.2, a contradiction.
$(b)$ If $v$ had both a $5(5)$-neighbour and a $4$-neighbour, then again $d_{2}(v)\leq 20$ by Lemma 2.2, a contradiction.
$(c)$ If $v$ had a $7(6)$-neighbour and two $4$-neighbours, then some edge of $E(v)$ would be contained in two $3$-faces, which gives $d_{2}(v)\leq 20$ by Lemma 2.2, a contradiction.
$(d)$ Let $m_{4}(v)=0$. Suppose that $v$ has at most one neighbour that is a $6(5^{-})$- or $7$-vertex. Then all remaining neighbours of $v$ are $6(6)$- or $5^{-}$-vertices. If $v$ has a $6(6)$-neighbour, then two edges of $E(v)$ lie in two $3$-faces. Considering all possible distributions of $5^{-}$- and $6(6)$-neighbours, one always obtains $d_{2}(v)\leq 20$ by Lemma 2.2, a contradiction. Thus $v$ must have at least two neighbours that are neither $5^{-}$- nor $6(6)$-vertices.
$(e)$ Let $m_{4}(v)=1$. Assume that $v$ has three neighbours forming $5^{-}$- or $6(6)$-vertices. Note that if $v$ has a $6(6)$-neighbour, then two edges of $E(v)$ lie on two $3$-faces. By checking all possible distributions of three neighbours among $5^{-}$-vertices and $6(6)$-vertices, we always obtain $d_{2}(v)\leq 20$ by Lemma 2.2, a contradiction.
$(f)$ Let $m_{4}(v)=1$ and suppose $v$ has a $7(7)$-neighbour. Then at least two edges of $E(v)$ are contained in two $3$-faces. If $v$ had two neighbours forming $5^{-}$- or $6(6)$-vertices, then $d_{2}(v)\leq 20$ by Lemma 2.2, a contradiction. ∎
Lemma 2.26.
Let $v$ be a $5(5)$-vertex.
  • $(a)$
    $n_{4}(v)=0$ and $n_{5}(v)\leq 2$.
  • $(b)$
    $v$ cannot have two neighbours that are $5(4^{+})$-, or $6(6)$-vertices. In particular, $v$ has at least one $7$-neighbour.
  • $(c)$
    If $v$ has four $6$-neighbours, then $v$ also has a $7(5^{-})$-neighbour.
  • $(d)$
    If $n_{5}(v)=0$ and $v$ has both a $6(5^{-})$- and a $6(6)$-neighbour, then $v$ has three $7$-neighbours, at least one of which is $7(5^{-})$-vertex.
  • $(e)$
    If $n_{5}(v)=1$, then $v$ has no $6(6)$-neighbours; in particular, $v$ has at most two $6$-neighbours.
  • $(f)$
    If $n_{5}(v)=1$ and $v$ has exactly one $6(5^{-})$-neighbour, then $v$ has a $7(5^{-})$-neighbour.
  • $(g)$
    If $n_{5}(v)=1$ and $v$ has exactly two $6(5^{-})$-neighbours, then $v$ has two $7(5^{-})$-neighbours.
  • $(h)$
    If $n_{5}(v)=2$, then $v$ has three $7(5^{-})$-neighbours.
Proof.
Let $f_{i}=v_{i}vv_{i+1}$ for $i\in[5]$ in cyclic order. If $d_{2}(v)\leq 20$, then taking $G^{\prime}=G-v$ gives a graph that is proper with respect to $G$, contradicting Remark 2.1. We therefore assume $d_{2}(v)\geq 21$.
$(a)$ By Lemma 2.21(a), $v$ has no $4$-neighbour. If $v$ had three $5$-neighbours, then $d_{2}(v)\leq 19$ by Lemma 2.2, a contradiction.
$(b)$ If $v$ had two neighbours consisting of $5(4^{+})$-, or $6(6)$-vertex, then $d_{2}(v)\leq 20$ by Lemma 2.2, a contradiction. Similarly, if all neighbours of $v$ were $6^{-}$-vertices, then again $d_{2}(v)\leq 20$, a contradiction. Hence $v$ must have at least one $7$-neighbour.
$(c)$ Suppose that $v$ has four $6$-neighbours. By (b), the fifth neighbour must be a $7$-vertex, say $v_{1}$. If $v_{1}$ were a $7(6^{+})$-vertex, then some edge in $E(v)$ would be contained in two $3$-faces, implying $d_{2}(v)\leq 20$ by Lemma 2.2, a contradiction. Hence $v_{1}$ is a $7(5^{-})$-vertex.
$(d)$ Let $n_{5}(v)=0$, and suppose that $v$ has both a $6(5^{-})$- and a $6(6)$-neighbour. If $v$ had only two $7$-neighbours, then $d_{2}(v)\leq 20$ by Lemma 2.2, a contradiction. Thus $v$ has three $7$-neighbours. At least one of them must be $7(5^{-})$-vertex; otherwise there would exist three edges of $E(v)$ such that each of them lies in two $3$-faces, again forcing $d_{2}(v)\leq 20$ by Lemma 2.2.
$(e)$ Let $n_{5}(v)=1$. If $v$ had a $6(6)$-neighbour or three $6$-neighbours, then $d_{2}(v)\leq 20$ by Lemma 2.2, a contradiction.
$(f)$ Let $n_{5}(v)=1$, and suppose that $v$ has exactly one $6(5^{-})$-neighbour. By (e), $v$ has no $6(6)$-neighbour, so the other three neighbours are $7$-vertices. If all were $7(6^{+})$-vertices, then at least two edges of $E(v)$ would lie in two $3$-faces, giving $d_{2}(v)\leq 20$ by Lemma 2.2, a contradiction. Thus $v$ has a $7(5^{-})$-neighbour.
$(g)$ Let $n_{5}(v)=1$, and suppose that $v$ has exactly two $6(5^{-})$-neighbours. By (a) and (e), $v$ has two $7$-neighbours. If one were a $7(6^{+})$-vertex, some edge of $E(v)$ would lie in two $3$-faces, giving $d_{2}(v)\leq 20$ by Lemma 2.2, a contradiction. Thus both $7$-neighbours are $7(5^{-})$-vertices.
$(h)$ Let $n_{5}(v)=2$. If $v$ had a $6$-neighbour or a $7(6^{+})$-neighbour, then $d_{2}(v)\leq 20$ by Lemma 2.2, a contradiction. ∎
Lemma 2.27.
Let $v$ be a $6(5)$-vertex. Then $v$ has at most four bad $5$-neighbours. In particular, if $m_{4}(v)=1$ and $v$ has four bad $5$-neighbours, then $v$ has two $7(5^{-})$-neighbours.
Proof.
Let $f_{i}=v_{i}vv_{i+1}$ for $i\in[5]$. First, we claim that $v$ cannot have five bad $5$-neighbours. Indeed, if all five neighbours were bad $5$-vertices, then at least three edges in $E(v)$ would each lie in two $3$-faces, forcing $d_{2}(v)\leq 19$ by Lemma 2.2. Since $v_{1}$ or $v_{6}$ must be a $5$-vertex (say $v_{1}$) and similarly one of $v_{3},v_{5}$ must be a $5$-vertex (say $v_{3}$), we may take $G^{\prime}=G-v+\{v_{1}v_{6},\,v_{3}v_{1},\,v_{3}v_{5}\},$ and $G^{\prime}$ is proper with respect to $G$, contradicting Remark 2.1. Thus $v$ has at most four bad $5$-neighbours.
Now assume $m_{4}(v)=1$ and $v$ has four bad $5$-neighbours. Then at least two edges in $E(v)$ lie in two $3$-faces. If $v$ had at most one $7(5^{-})$-neighbour, then $d_{2}(v)\leq 20$ by Lemma 2.2, a contradiction. ∎
Let $v$ be a $7$-vertex, and let $x$ be a $4$-neighbour of $v$. If $v$ and $x$ have two common neighbours (equivalently, the edge $vx$ is contained in two $3$-faces), then $x$ is called a support neighbour of $v$.
A vertex is said to be poor if it is a $4$-vertex or a $5(5)$-vertex. Since $G$ has no $4(4)$-vertices, the poor vertices are exactly $4(0)$-, $4(1)$-, $4(2)$-, $4(3)$-, and $5(5)$-vertices.
Lemma 2.28.
Let $v$ be a $7$-vertex. Then:
  • $(a)$
    If $v$ has a support neighbour, then $d_{2}(v)\geq 21$.
  • $(b)$
    If $4\leq m_{3}(v)\leq 5$, then $v$ has at most six $4$-neighbours.
  • $(c)$
    If $m_{3}(v)=5$ and $n_{4}(v)\leq 2$, then $v$ has at most four poor neighbours.
  • $(d)$
    If $v$ is a $7(6)$-vertex, then $v$ has at most five poor neighbours.
  • $(e)$
    If $n_{4}(v)=3$, then $v$ has at most two $5(5)$-neighbours.
  • $(f)$
    If $n_{4}(v)=4$, then $v$ has at most one $5(5)$-neighbour.
  • $(g)$
    If $m_{3}(v)=5$, $n_{4}(v)=4$, $v$ has a support neighbour, and $v$ has both a $5(4)$- and a $5(5)$-neighbour, then $v$ must be adjacent to a $7$-vertex.
  • $(h)$
    If $m_{3}(v)=5$, $n_{4}(v)\geq 3$, $v$ has a support neighbour, and $v$ has a $3$-neighbour, then $v$ has at most four poor neighbours.
  • $(i)$
    If $m_{3}(v)=5$, $n_{4}(v)=5$, then $v$ has at most one $5(4)$-neighbour. In particular, if $v$ has a $5(4)$-neighbour or two $6(5)$-neighbours, then $v$ is incident to two $5^{+}$-faces.
  • $(j)$
    If $m_{3}(v)=5$, $n_{4}(v)=6$, then $v$ is incident to two $5^{+}$-faces; in particular, $v$ has no bad $5$-neighbour.
Proof.
$(a)$ Let $v_{1}$ be a support neighbour of $v$. Then $v_{1}$ is a $4$-vertex adjacent to both $v_{2}$ and $v_{7}$. Assume for a contradiction that $d_{2}(v)\leq 20$. Take $G^{\prime}=G-v+\{v_{1}v_{3},\,v_{1}v_{4},\,v_{1}v_{5},\,v_{1}v_{6}\}.$ Then $G^{\prime}$ is proper with respect to $G$, contradicting Remark 2.1. Thus $d_{2}(v)\geq 21$.
$(b)$ Let $4\leq m_{3}(v)\leq 5$. If all neighbours of $v$ were $4$-vertices, then $d_{2}(v)\geq 20$; and since $m_{3}(v)\geq 4$, at least one neighbour must be a support neighbour. By part (a), this yields a contradiction. Hence $v$ has at most six $4$-neighbours.
$(c)$ Assume $m_{3}(v)=5$ and $n_{4}(v)\leq 2$. Suppose for a contradiction that $v$ has five poor neighbours. By Lemma 2.26(b), a $5(5)$-vertex is adjacent to at most one $5(5)$-vertex. Since $m_{3}(v)=5$, $v$ can have at most three $5(5)$-neighbours. If $v$ has five poor neighbours, then it must have exactly three $5(5)$-neighbours and two $4$-neighbours, contradicting Lemma 2.26(a), which states that $5(5)$-vertices have no $4$-neighbours.
$(d)$ Let $v$ be a $7(6)$-vertex. By Lemma 2.26(a),(b), a $5(5)$-vertex has no $4$-neighbour and is adjacent to at most one bad $5$-vertex. Also, by Lemma 2.21(b), a $4(1^{+})$-vertex is adjacent to at most one $4$-vertex. These restrictions prevent $v$ from having six poor neighbours.
$(e)$ Let $n_{4}(v)=3$. Since a $5(5)$-vertex has no $4$-neighbour and is adjacent to at most one bad $5$-vertex (Lemma 2.26(a),(b)), $v$ can have at most two $5(5)$-neighbours.
$(f)$ Let $n_{4}(v)=3$. Then similar reasoning as in (d) implies that $v$ can have at most one $5(5)$-neighbour.
$(g)$ Assume $m_{3}(v)=5$, $n_{4}(v)=4$, $v$ has a support neighbour, and $v$ has both a $5(4)$- and a $5(5)$-neighbour. If the seventh neighbour of $v$ were a $6^{-}$-vertex, then $d_{2}(v)\leq 20$ by Lemma 2.2, contradicting part (a). Hence the remaining neighbour must be a $7$-vertex.
$(h)$ Assume the stated conditions. If $v$ had five poor neighbours, then $d_{2}(v)\leq 20$ by Lemma 2.2, contradicting (a).
$(i)$ Assume $m_{3}(v)=5$ and $n_{4}(v)=5$. Then $v$ has a support neighbour. If $v$ had two $5(4)$-neighbours, then $d_{2}(v)\leq 20$ by Lemma 2.2, contradicting (a). Thus $v$ has at most one such neighbour. Similarly, if $v$ has one $5(4)$-neighbour or two $6(5)$-neighbours, then $v$ must be incident to two $5^{+}$-faces.
$(j)$ Assume $m_{3}(v)=5$ and $n_{4}(v)=6$. Then $v$ has a support neighbour. If $v$ were not incident to two $5^{+}$-faces, then $d_{2}(v)\leq 20$ by Lemma 2.2, contradicting (a). Thus $v$ is incident to two $5^{+}$-faces. Similarly we deduce that $v$ has no $5(4)$- and $5(5)$-neighbours.. ∎
We now apply discharging to show that $G$ does not exist. We use the same initial charges as in the case $\Delta=6$, and redistribute charge according to the following rules.
Discharging Rules
We apply the following discharging rules.
  • R1:
    Every $3$-face receives $\frac{1}{3}$ from each of its incident vertices.
  • R2:
    Every $5^{+}$-face gives $\frac{1}{5}$ to each of its incident vertices.
  • R3:
    Every $6(3^{-})$-vertex gives $\frac{1}{6}$ to each of its neighbours.
  • R4:
    Every $6(4)$-vertex gives $\frac{1}{9}$ to each of its neighbours.
  • R5:
    Every $6(5)$-vertex gives $\frac{1}{9}$ to each of its bad $5$-neighbours.
  • R6:
    Every $7(3^{-})$-vertex gives $\frac{2}{7}$ to each of its neighbours.
  • R7:
    Let $v$ be a $7(4)$- or $7(5)$-vertex. Then $v$ gives
    • $(a)$
      $\frac{1}{5}$ to each of its $3$-neighbours,
    • $(b)$
      $\frac{1}{4}$ to each of its bad $4$-neighbours,
    • $(c)$
      $\frac{2}{9}$ to each of its $5(5)$-neighbours,
    • $(d)$
      $\frac{1}{9}$ to each of its $5(4)$-neighbours,
    • $(e)$
      $\frac{1}{18}$ to each of its $6(5)$-neighbours.
  • R8:
    Let $v$ be a $7(6)$-vertex. Then $v$ gives
    • $(a)$
      $\frac{1}{6}$ to each of its $5(5)$-neighbours,
    • $(b)$
      $\frac{1}{9}$ to each of its $5(4)$-neighbours,
    • $(c)$
      $\frac{1}{6}$ to each of its bad $4$-neighbours.
  • R9:
    Let $v$ be a $7(7)$-vertex. Then $v$ gives
    • $(a)$
      $\frac{1}{6}$ to each of its $5(5)$-neighbours,
    • $(b)$
      $\frac{1}{12}$ to each of its $5(4)$-neighbours.
Checking $\mu^{*}(v),\mu^{*}(f)\geq 0$ for $v\in V(G),f\in F(G)$.
First we show that $\mu^{*}(f)\geq 0$ for each $f\in F(G)$. Recall that every face $f$ has initial charge $\mu(f)=\ell(f)-4$. If $f$ is a $3$-face, then by R1 it receives $\frac{1}{3}$ from each of its incident vertices, and hence $\mu^{*}(f)\geq-1+3\times\frac{1}{3}=0$. If $f$ is a $4$-face, then $\mu(f)=0$ and it neither sends nor receives charge, so $\mu^{*}(f)=\mu(f)=0$. Let $f$ be a $5^{+}$-face. By R2, the face $f$ sends $\frac{1}{5}$ to each of its incident vertices. Therefore, $\mu^{*}(f)=\ell(f)-4-\ell(f)\times\frac{1}{5}=\frac{4\ell(f)}{5}-4\geq 0$. Consequently, every face $f\in F(G)$ satisfies $\mu^{*}(f)\geq 0$.
Next, let $v\in V(G)$ be a vertex of degree $d(v)=k$. By Lemma 2.18, we have $k\geq 3$.
(1). Let $k=3$. The initial charge of $v$ is $\mu(v)=d(v)-4=-1$. By Lemma 2.19, $m_{3}(v)=0$, $m_{4}(v)\leq 1$ and each neighbour of $v$ is a $7(5^{-})$-vertex. This implies that $v$ is incident to at least two $5^{+}$-faces. Then $v$ receives $\frac{1}{5}$ from each of its incident $5^{+}$-faces by R2, and at least $\frac{1}{5}$ from each of its $7(5^{-})$-neighbours by R6, R7(a). Hence, $\mu^{*}(v)\geq-1+2\times\frac{1}{5}+3\times\frac{1}{5}=0$.
(2). Let $k=4$. The initial charge of $v$ is $\mu(v)=d(v)-4=0$. We have $m_{3}(v)\leq 3$ by Lemma 2.20. If $m_{3}(v)=0$, then $\mu^{*}(v)\geq 0$ since $v$ does not give a charge to its any incident faces. So we may assume that $1\leq m_{3}(v)\leq 3$.
(2.1). Let $m_{3}(v)=1$. If $m_{4}(v)\leq 1$, then $v$ is incident to two $5^{+}$-faces, and by R2, $v$ receives $\frac{1}{5}$ from each of those $5^{+}$-faces. Thus, $\mu^{*}(v)\geq 2\times\frac{1}{5}-\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to its incident $3$-face by R1. If $m_{4}(v)=2$, then $v$ has either a $7(5^{-})$-neighbour or two $7(6)$-neighbours by Lemma 2.22(a). In both cases, $v$ receives at least $\frac{1}{4}=\min\{\frac{1}{4},2\times\frac{1}{6}\}$ from its $7$-neighbours by R6-R8. On the other hand, $v$ receives $\frac{1}{5}$ from its incident $5^{+}$-face by R2. Thus, $\mu^{*}(v)\geq\frac{1}{4}+\frac{1}{5}-\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to its incident $3$-face by R1. If $m_{4}(v)=3$, then $v$ has two $7(6^{-})$-neighbours by Lemma 2.22(b). By applying R6-R8, $v$ receives totally at least $\frac{1}{3}$ from its $7$-neighbours. Thus, $\mu^{*}(v)\geq\frac{1}{3}-\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to its incident $3$-face by R1.
(2.2). Let $m_{3}(v)=2$. Since $0\leq m_{4}(v)\leq 2$, we consider the following cases:
If $m_{4}(v)=0$, then $v$ has either a $7(5^{-})$-neighbour and a $6(4^{-})$-neighbour or two $7(6^{-})$-neighbours by Lemma 2.23(a). It follows that $v$ receives totally at least $\frac{1}{3}=\min\{\frac{1}{4}+\frac{1}{9},2\times\frac{1}{6}\}$ from its $6^{+}$-neighbours by R3-R4 and R6-R8. In addition, $v$ is incident to two $5^{+}$-faces, and by R2, $v$ receives $\frac{1}{5}$ from each of those $5^{+}$-faces. Thus, $\mu^{*}(v)\geq\frac{1}{3}+2\times\frac{1}{5}-2\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
If $m_{4}(v)=1$, then $v$ has either two $7(6^{-})$-neighbours and two $6(4^{-})$-neighbours or three $7(6^{-})$-neighbours by Lemma 2.23(b). In both cases, $v$ receives totally at least $\frac{1}{2}=\min\{2\times\frac{1}{6}+2\times\frac{1}{9},3\times\frac{1}{6}\}$ from its $6^{+}$-neighbours by R3-R4 and R6-R8. Also, $v$ receives $\frac{1}{5}$ from its incident $5^{+}$-face by R2. Thus, $\mu^{*}(v)\geq\frac{1}{2}+\frac{1}{5}-2\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
If $m_{4}(v)=2$, then $v$ has either two $7(5^{-})$-neighbours and two $7(6)$-neighbours or three $7(5^{-})$-neighbours by Lemma 2.23(c). In both cases, $v$ receives totally at least $\frac{3}{4}=\min\{2\times\frac{1}{4}+2\times\frac{1}{6},3\times\frac{1}{4}\}$ from its $7^{+}$-neighbours by R6-R8. Thus, $\mu^{*}(v)\geq\frac{3}{4}-2\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
(2.3). Let $m_{3}(v)=3$. Suppose first that $m_{4}(v)=0$. It then follows from Lemma 2.24(a) that then $v$ has either four $7(5^{-})$-neighbours or three $7(5^{-})$-neighbours and a $7(6)$-neighbour or three $7(5^{-})$-neighbours and a $6(4^{-})$-neighbour or two $7(5^{-})$-neighbours and two $7(6)$-neighbours. In each case, $v$ receives at least $\frac{5}{6}=\min\{4\times\frac{1}{4},3\times\frac{1}{4}+\frac{1}{6},3\times\frac{1}{4}+\frac{1}{9},2\times\frac{1}{4}+2\times\frac{1}{6}\}$ from its $6^{+}$-neighbours by R3-R8. In addition, $v$ receives $\frac{1}{5}$ from its incident $5^{+}$-face by R2. Thus, $\mu^{*}(v)\geq\frac{5}{6}+\frac{1}{5}-3\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1. Now, we suppose that $m_{4}(v)=1$. By Lemma 2.24(b), all neighbours of $v$ are $7(5^{-})$-vertices, and so $v$ receives totally $4\times\frac{1}{4}$ from its $7(5^{-})$-neighbours by R6-R7. Thus, $\mu^{*}(v)\geq 4\times\frac{1}{4}-3\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
(3). Let $k=5$. The initial charge of $v$ is $\mu(v)=d(v)-4=1$. We distinguish three cases according to the number of $3$-faces incident to $v$ as follows.
(3.1). Let $m_{3}(v)\leq 3$. Obviously, $\mu^{*}(v)\geq 1-3\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
(3.2). Let $m_{3}(v)=4$. Suppose first that $m_{4}(v)=0$. By Lemma 2.25(d), $v$ has two neighbours different from $5^{-}$- and $6(6)$-vertices, i.e., $v$ has two neighbours consisting of $6(5^{-})$- or $7$-vertices. It follows that $v$ receives totally at least $\frac{1}{6}=\min\{2\times\frac{1}{9},\frac{1}{9}+\frac{1}{12},2\times\frac{1}{12}\}$ from its $6^{+}$-neighbours by R3-R9, and $\frac{1}{5}$ from its incident $5^{+}$-face by R2. Thus, $\mu^{*}(v)\geq 1+\frac{1}{6}+\frac{1}{5}-4\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
Next we suppose that $m_{4}(v)=1$. By Lemma 2.25(e), $v$ has at least three $6^{+}$-neighbours different from $6(6)$-vertex. If $v$ has no $7(7)$-neighbour, then $v$ receives totally at least $3\times\frac{1}{9}$ from its $6^{+}$-neighbours by R3-R8. If $v$ has a $7(7)$-neighbours, then $v$ has four $6^{+}$-neighbours different from $6(6)$-vertex by Lemma 2.25(f), and so $v$ receives totally at least $4\times\frac{1}{12}$ from its $6^{+}$-neighbours by R3-R9. Thus, $\mu^{*}(v)\geq 1+\frac{1}{3}-4\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
(3.3). Let $m_{3}(v)=5$. Note that $v$ has no $4^{-}$-neighbour, and $v$ has at most two $5$-neighbours by Lemma 2.26(a).
(3.3.1). Let $n_{5}(v)=0$. By Lemma 2.26(b), $v$ has at most one $6(6)$-neighbour, i.e., $v$ has at least four $6^{+}$-neighbours different from $6(6)$-vertex. First, suppose that $v$ has a $6(6)$-neighbour. If $v$ has no $6(5^{-})$-neighbour, then all the neighbours of $v$ different from $6(6)$-vertex are $7$-vertices, and by applying R6-R9, $v$ receives totally at least $4\times\frac{1}{6}$ from its $7$-neighbours. So, $\mu^{*}(v)\geq 1+4\times\frac{1}{6}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1. On the other hand, if $v$ has a $6(5^{-})$-neighbour, then, by Lemma 2.26(d), $v$ has three $7$-neighbours, at least one of which is a $7(5^{-})$-vertex. Then $v$ receives at least $\frac{1}{9}$ from its $6(5^{-})$-neighbour by R3-R5, at least $\frac{2}{9}$ from each of its $7(5^{-})$-neighbours by R6-R7, at least $\frac{1}{6}$ from each of its other $7$-neighbours by R8-R9. So, $v$ receives totally at least $\frac{1}{9}+\frac{2}{9}+2\times\frac{1}{6}=\frac{2}{3}$ from its $6^{+}$-neighbours. Thus, $\mu^{*}(v)\geq 1+\frac{2}{3}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
Suppose now that $v$ has no $6(6)$-neighbour. That is, all neighbours of $v$ are $6^{+}$-vertices different from $6(6)$-vertex. By Lemma 2.26(b), $v$ has at least one $7$-neighbour, say $x$. Note that if $v$ has exactly four $6$-neighbours, then $x$ must be a $7(5^{-})$-vertex by Lemma 2.26(c). This means that $v$ has either four $6$-neighbours and one $7(5^{-})$-neighbour or at most three $6$-neighbours and two $7$-neighbours. In each case, $v$ receives totally at least $\frac{2}{3}=\min\{4\times\frac{1}{9}+\frac{2}{9},3\times\frac{1}{9}+2\times\frac{1}{6}\}$ from its $6^{+}$-neighbours by R3-R9. Thus, $\mu^{*}(v)\geq 1+\frac{2}{3}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
(3.3.2). Let $n_{5}(v)=1$. Notice that $v$ has no $6(6)$-neighbours by Lemma 2.26(e), i.e., all neighbours of $v$ but one are $7$- or $6(5^{-})$-vertices. Moreover, $v$ has at most two $6(5^{-})$-neighbours by Lemma 2.26(e). We distinguish three cases according to the number of $6(5^{-})$-vertices adjacent to $v$ as follows.
First suppose that $v$ has no $6(5^{-})$-neighbour. So, $v$ has four $7$-neighbours, and by applying R6-R9, $v$ receives totally at least $4\times\frac{1}{6}$ from its $7$-neighbours. So, $\mu^{*}(v)\geq 1+4\times\frac{1}{6}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
Suppose next that $v$ has exactly one $6(5^{-})$-neighbour. Then $v$ has three $7$-neighbours, and by Lemma 2.26(f), one of which is a $7(5^{-})$-vertex. Thus, $v$ receives totally at least $\frac{1}{9}+2\times\frac{1}{6}+\frac{2}{9}$ from its $6^{+}$-neighbours by R3-R9. Hence, $\mu^{*}(v)\geq 1+\frac{2}{3}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
Finally, suppose that $v$ has exactly two $6(5^{-})$-neighbours. Then, by Lemma 2.26(g), $v$ has two $7(5^{-})$-neighbours. So, $v$ receives totally at least $2\times\frac{1}{9}+2\times\frac{2}{9}$ from its $6^{+}$-neighbours by R3-R7. Hence, $\mu^{*}(v)\geq 1+\frac{2}{3}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
(3.3.3). Let $n_{5}(v)=2$. Then $v$ has three $7(5^{-})$-neighbours by Lemma 2.26(h), and so $v$ receives $\frac{2}{9}$ from each of its $7(5^{-})$-neighbours by R6-R7. Thus, $\mu^{*}(v)\geq 1+3\times\frac{2}{9}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
(4). Let $k=6$. The initial charge of $v$ is $\mu(v)=d(v)-4=2$. If $m_{3}(v)\leq 3$, then $\mu^{*}(v)\geq 2-3\times\frac{1}{3}-6\times\frac{1}{6}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and $\frac{1}{6}$ to each of its neighbours by R3. Similarly, if $m_{3}(v)=4$, then $\mu^{*}(v)\geq 2-4\times\frac{1}{3}-6\times\frac{1}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and $\frac{1}{9}$ to each of its neighbours by R4. On the other hand, if $m_{3}(v)=6$, then $\mu^{*}(v)\geq 2-6\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1. Therefore, we further assume that $m_{3}(v)=5$. By Lemma 2.27, $v$ has at most four bad $5$-neighbours. Notice that if $v$ has at most three bad $5$-neighbours, then $\mu^{*}(v)\geq 2-5\times\frac{1}{3}-3\times\frac{1}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and $\frac{1}{9}$ to each of its bad $5$-neighbours by R5. Next we assume that $v$ has exactly four bad $5$-neighbours. If $m_{4}(v)=0$, then $v$ receives $\frac{1}{5}$ from its incident $5^{+}$-face by R2, and so $\mu^{*}(v)\geq 2+\frac{1}{5}-5\times\frac{1}{3}-4\times\frac{1}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-face by R1, and $\frac{1}{9}$ to each of its bad $5$-neighbours by R5. If $m_{4}(v)=1$, then $v$ has two $7(5^{-})$-neighbours by Lemma 2.27, and so $v$ receives at least $\frac{1}{18}$ from each of its $7(5^{-})$-neighbours by R6 and R7(e). Thus, $\mu^{*}(v)\geq 2+2\times\frac{1}{18}-5\times\frac{1}{3}-4\times\frac{1}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-face by R1, and $\frac{1}{9}$ to each of its bad $5$-neighbours by R5.
(5). Let $k=7$. The initial charge of $v$ is $\mu(v)=d(v)-4=3$. Notice first that if $m_{3}(v)\leq 3$, then we have $\mu^{*}(v)\geq 3-3\times\frac{1}{3}-7\times\frac{2}{7}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and $\frac{2}{7}$ to each of its neighbours by R6. Therefore, we may assume that $4\leq m_{3}(v)\leq 7$.
(5.1). Let $m_{3}(v)=4$. By Lemma 2.28(b), $v$ has at most six $4$-neighbours. If $n_{4}(v)\leq 4$, then $v$ sends $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), and at most $\frac{2}{9}$ to each of its other neighbours by R7. Consequently, $\mu^{*}(v)\geq 3-4\times\frac{1}{4}-3\times\frac{2}{9}-4\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1. Suppose now that $5\leq n_{4}(v)\leq 6$. Clearly, $v$ has no $5(5)$-neighbour, since a $4$ vertex is not adjacent to any $5(5)$-vertex by Lemma 2.21(a). If $v$ has a $3$-neighbour $x$, then the edge $xv$ is contained in at least one $5^{+}$-face by Lemma 2.19. So, $v$ receives at least $\frac{1}{5}$ from its incident $5^{+}$-face by R2. Thus $\mu^{*}(v)\geq 3+\frac{1}{5}-6\times\frac{1}{4}-\frac{1}{5}-4\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), at most $\frac{1}{5}$ to each of its neighbours other than $4$-vertex by R7, and $\frac{1}{3}$ to each of its incident $3$-faces by R1. On the other hand, if $v$ has no $3$-neighbour, then $\mu^{*}(v)\geq 3-6\times\frac{1}{4}-\frac{1}{9}-4\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), at most $\frac{1}{9}$ to each of its neighbours other than $4$-vertex by R7, and $\frac{1}{3}$ to each of its incident $3$-faces by R1.
(5.2). Let $m_{3}(v)=5$. Notice that $v$ has at most one $3$-neighbour, since a $3$-vertex does not incident to any $3$-faces by Lemma 2.19. First we assume that $v$ has a $3$-neighbour $x$. In such a case, all $3$-faces incident to $v$ must have a consecutive ordering. Moreover, the edge $xv$ is contained in at least one $5^{+}$-face by Lemma 2.19. So, $v$ receives at least $\frac{1}{5}$ from its incident $5^{+}$-face by R2. Observe first that if $v$ has three $4$-neighbours, then at least one of them must be a support neighbour of $v$. In such a case, $v$ has at most four poor neighbours by Lemma 2.28(h), and recall that a poor vertex receives at most $\frac{1}{4}$ from $v$ by R7(b),(c). It then follows that $\mu^{*}(v)\geq 3+\frac{1}{5}-5\times\frac{1}{3}-\frac{1}{5}-4\times\frac{1}{4}-2\times\frac{1}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{5}$ to its $3$-neighbour $x$ by R7(a), at most $\frac{1}{4}$ to each of its poor neighbours by R7(b),(c), and at most $\frac{1}{9}$ to each of its other neighbours by R7(d),(e). We now suppose that $v$ has at most two $4$-neighbours. It then follows from Lemma 2.28(c) that $v$ has at most four poor neighbours. Thus, similarly as above, we have $\mu^{*}(v)\geq 3+\frac{1}{5}-5\times\frac{1}{3}-\frac{1}{5}-2\times\frac{1}{4}-2\times\frac{2}{9}-2\times\frac{1}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{5}$ to its $3$-neighbour $x$ by R7(a), at most $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), $\frac{2}{9}$ to each of its $5(5)$-neighbours by R7(c), and at most $\frac{1}{9}$ to each of its other neighbours by R7(d),(e).
We may further assume that $v$ has no $3$-neighbours. By Lemma 2.28(b), $v$ has at most six $4$-neighbours, i.e., $n_{4}(v)\leq 6$. On the other hand, recall that $m_{3}(v)=5$. Therefore, if $v$ has more than four $4$-neighbours, then there exists $v_{i}\in N(v)$ such that $vv_{i}$ is contained in two $3$-faces, i.e., $v_{i}$ is a support neighbour.
(5.2.1). Let $n_{4}(v)\leq 2$. Note that $v$ has at most four poor neighbours by Lemma 2.28(c). Then $\mu^{*}(v)\geq 3-5\times\frac{1}{3}-2\times\frac{1}{4}-2\times\frac{2}{9}-3\times\frac{1}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), $\frac{2}{9}$ to each of its $5(5)$-neighbours by R7(c), at most $\frac{1}{9}$ to each of its other neighbours by R7(d),(e).
(5.2.2). Let $n_{4}(v)=3$. By Lemma 2.28(e), $v$ has at most two $5(5)$-neighbours. Suppose first that $v$ has exactly two $5(5)$-neighbours $x,y$. Since a $4$-vertex has no $5(5)$-neighbours by Lemma 2.21(a), we deduce that $x$ and $y$ are adjacent. On the other hand, a $5(5)$-vertex has at most one bad $5$-neighbour by Lemma 2.26(b). This implies that the neighbours of $v$ other than $4$- and $5(5)$-vertices are $5(3^{-})$- or $6^{+}$-vertices. Thus we have $\mu^{*}(v)\geq 3-5\times\frac{1}{3}-3\times\frac{1}{4}-2\times\frac{2}{9}-2\times\frac{1}{18}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), $\frac{2}{9}$ to each of its $5(5)$-neighbours by R7(c), at most $\frac{1}{18}$ to each of its other neighbours by R7(e). Suppose now that $v$ has at most one $5(5)$-neighbour. In such a case, we have again $\mu^{*}(v)\geq 3-5\times\frac{1}{3}-3\times\frac{1}{4}-\frac{2}{9}-3\times\frac{1}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), $\frac{2}{9}$ to each of its $5(5)$-neighbours by R7(c), at most $\frac{1}{9}$ to each of its other neighbours by R7(d),(e).
(5.2.3). Let $n_{4}(v)=4$. By Lemma 2.28(f), $v$ has at most one $5(5)$-neighbour. First, suppose that $v$ has no $5(5)$-neighbours, then $\mu^{*}(v)\geq 3-5\times\frac{1}{3}-4\times\frac{1}{4}-3\times\frac{1}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), at most $\frac{1}{9}$ to each of its other neighbours by R7(d)-(e). We may further assume that $v$ has exactly one $5(5)$-neighbour, say $v_{1}$. Recall that each of $v_{2},v_{7}$ is a $5^{+}$-vertex by Lemmas 2.19 and 2.21(a). Moreover, for $i\in\{2,7\}$, if $v_{i}$ is adjacent to a $4$-vertex, then $v_{i}$ is different from $5(4)$-vertex by Lemma 2.25(b). Recall also that a $4(1^{+})$-vertex is adjacent to at most one $4$-vertex by Lemma 2.21(b).
Suppose first that $v$ has a support neighbour $x$. If $v$ has no $5(4)$-neighbour, then $\mu^{*}(v)\geq 3-5\times\frac{1}{3}-4\times\frac{1}{4}-\frac{2}{9}-2\times\frac{1}{18}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), $\frac{2}{9}$ to its $5(5)$-neighbour by R7(c), at most $\frac{1}{18}$ to each of its other neighbours by R7(e). If $v$ has a $5(4)$-neighbour, then $v$ has also a $7$-neighbour by Lemma 2.28(g). Thus, $\mu^{*}(v)\geq 3-5\times\frac{1}{3}-4\times\frac{1}{4}-\frac{2}{9}-\frac{1}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), $\frac{2}{9}$ to its $5(5)$-neighbour by R7(c), $\frac{1}{9}$ to its $5(4)$-neighbour by R7(d).
Next, suppose that $v$ has no support neighbour. This means that each $4$-neighbour of $v$ is incident to a $4^{+}$-face containing $v$. Obviously, $v$ has one of the two configurations depicted in Figure 4, where $v_{i},v_{j},v_{k},v_{\ell}$ are $4$-vertices. By Lemma 2.21(a), a $4$-vertex has no $5(5)$-neighbours, so the configuration in Figure 4(a) is not possible for $v$. So the neighbours of $v$ can only form as depicted in Figure 4(b). We then infer that $v$ cannot have any $5(4)$-neighbour by Lemma 2.25(b). Thus $\mu^{*}(v)\geq 3-5\times\frac{1}{3}-4\times\frac{1}{4}-\frac{2}{9}-2\times\frac{1}{18}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), $\frac{2}{9}$ to its $5(5)$-neighbour by R7(c), at most $\frac{1}{18}$ to each of its other neighbours by R7(e).
[…]
Figure 4. Two possible configurations of a $7(5)$-vertex $v$.
(5.2.4). Let $n_{4}(v)=5$. Then $v$ has no $5(5)$-neighbours by Lemma 2.21(a). Also, $v$ has at most one $5(4)$-neighbour by Lemma 2.28(i). Suppose first that $v$ has a $5(4)$-neighbour or two $6(5)$-neighbours. Then $v$ is incident to two $5^{+}$-face by Lemma 2.28(i). By applying R2, $v$ receives $\frac{1}{5}$ from each of its incident $5^{+}$-faces. In such a case, we have $\mu^{*}(v)\geq 3+2\times\frac{1}{5}-5\times\frac{1}{3}-5\times\frac{1}{4}-\frac{1}{9}-\frac{1}{18}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), $\frac{1}{9}$ to its $5(4)$-neighbour by R7(d), $\frac{1}{18}$ to each of its $6(5)$-neighbours by R7(e). Now we suppose that $v$ has no $5(4)$-neighbour, and $v$ has at most one $6(5)$-neighbour. Then $\mu^{*}(v)\geq 3-5\times\frac{1}{3}-5\times\frac{1}{4}-\frac{1}{18}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), $\frac{1}{18}$ to its $6(5)$-neighbour by R7(e).
(5.2.5). Let $n_{4}(v)=6$. By Lemma 2.28(j), $v$ is incident to two $5^{+}$-faces, in particular, $v$ has no $5(4)$- or $5(5)$-neighbours. By R2, $v$ receives $\frac{1}{5}$ from each of its incident $5^{+}$-faces. Thus we have $\mu^{*}(v)\geq 3+2\times\frac{1}{5}-5\times\frac{1}{3}-6\times\frac{1}{4}-\frac{1}{18}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), $\frac{1}{18}$ to its $6(5)$-neighbour by R7(e).
(5.3). Let $m_{3}(v)=6$. By Lemma 2.28(d), $v$ has at most five poor neighbours. If $v$ has at most four poor neighbours, then $\mu^{*}(v)\geq 3-6\times\frac{1}{3}-4\times\frac{1}{6}-3\times\frac{1}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{6}$ to each of its poor neighbours by R8(a),(c), $\frac{1}{9}$ to each of its $5(4)$-neighbours by R8(b). Suppose further that $v$ has exactly five poor neighbours. First, observe that $v$ has no $4(0)$-neighbours as $m_{3}(v)=6$, i.e., all $4$-neighbours of $v$ are bad $4$-vertices. Note that, by Lemma 2.25(a),(b), a $5(4)$-neighbour of $v$ is not adjacent to two $5(5)$-vertices, and is not adjacent to both a $5(5)$- and $4$-vertices. Moreover, a $4$-vertex has no $5(5)$-neighbours by Lemma 2.21(a). Furthermore, a bad $4$-vertex has at most one $4$-neighbour by Lemma 2.21(b), and has no both $4$- and $5(4)$-neighbour by Lemma 2.21(c). In addition, by Lemma 2.26(b), a $5(5)$-vertex is adjacent to at most one $5(5)$-vertex. All those facts imply that $v$ has at most one $5(4)$-neighbour. Hence, $\mu^{*}(v)\geq 3-6\times\frac{1}{3}-5\times\frac{1}{6}-\frac{1}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{6}$ to each of its poor neighbours by R8(a),(c), $\frac{1}{9}$ to its $5(4)$-neighbour by R8(b).
(5.4). Let $m_{3}(v)=7$. By Lemma 2.26(b), a $5(5)$-vertex is adjacent to at most one $5(5)$-vertex. So, we deduce that $v$ has at most four $5(5)$-neighbours. In particular, if $v$ has four such neighbours, then $v$ cannot have any $5(4)$-neighbours by Lemmas 2.25(a) and 2.26(b). Thus $\mu^{*}(v)\geq 3-7\times\frac{1}{3}-4\times\frac{1}{6}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1 and $\frac{1}{6}$ to each of its $5(5)$-neighbours by R9(a). If $v$ has three $5(5)$-neighbours, then $v$ has at most two $5(4)$-neighbours by Lemmas 2.25(a) and 2.26(b), and so $\mu^{*}(v)\geq 3-7\times\frac{1}{3}-3\times\frac{1}{6}-2\times\frac{1}{12}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-face by R1, $\frac{1}{6}$ to each of its $5(5)$-neighbour by R9(a), and $\frac{1}{12}$ to each of its $5(4)$-neighbours by R9(b). Suppose now that $v$ has at most two $5(5)$-neighbours. By Lemma 2.26(b), a $5(5)$-vertex has at most one bad $5$-neighbour. We then deduce that $v$ has either seven $5(4)$-neighbours or at most six bad $5$-neighbours. In the worst case, $\mu^{*}(v)\geq 3-7\times\frac{1}{3}-2\times\frac{1}{6}-4\times\frac{1}{12}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-face by R1, $\frac{1}{6}$ to each of its $5(5)$-neighbour by R9(a), and $\frac{1}{12}$ to each of its $5(4)$-neighbours by R9(b).
2.3. The case $\Delta=8$
Recall that $G$ does not admit any $2$-distance $23$-coloring, whereas any planar graph $G^{\prime}$ obtained from $G$ with smaller value of $|V(G^{\prime})|+|E(G^{\prime})|$ admits a $2$-distance $23$-coloring.
We begin by establishing several structural properties of $G$, analogous to those obtained for the cases $\Delta=6$ and $\Delta=7$. The proof of the following lemma is omitted, as it follows from the same arguments used in Lemma 2.3.
Lemma 2.29.
$\delta(G)\geq 3$.
Lemma 2.30.
Let $v$ be a $3$-vertex. Then $m_{3}(v)=0$ and $m_{4}(v)\leq 1$. In particular,
  • $(a)$
    $v$ has no $6^{-}$-neighbour,
  • $(b)$
    if $m_{4}(v)=0$, then $v$ has two $8(6^{-})$-neighbours,
  • $(c)$
    if $m_{4}(v)=1$, then $v$ has three $8(6^{-})$-neighbours.
Proof.
If $v$ is incident to a $3$-face, say $f_{1}=v_{1}vv_{2}$, then we set $G^{\prime}=G-v+\{v_{1}v_{3}\}$. If $v$ is incident to two $4$-faces, say $f_{1}=v_{1}vv_{2}x$ and $f_{2}=v_{2}vv_{3}y$, then we again set $G^{\prime}=G-v+\{v_{1}v_{3}\}$. If $v$ is adjacent to a $7^{-}$-vertex, say $v_{1}$, then we set $G^{\prime}=G-v+\{v_{1}v_{2},v_{1}v_{3}\}$. In each case, $G^{\prime}$ remains proper with respect to $G$, and moreover $d_{2}(v)\leq 22$ by Lemma 2.2. This yields a contradiction by Remark 2.1. Thus, none of the above configurations can occur. ∎
Since a $3$-vertex cannot have any $6^{-}$-neighbour by Lemma 2.30(a), it follows that every vertex of degree $4$, 55, or $6$ has no $3$-neighbour. This fact will be assumed in the remainder of this section.
Lemma 2.31.
Let $v$ be a $4$-vertex with $m_{3}(v)\geq 1$.
  • $(a)$
    If $m_{3}(v)=1$, then $v$ has no $5(5)$-neighbour. In particular, $v$ has at most one $4$-neighbour.
  • $(b)$
    If $m_{3}(v)=2$, then $v$ has at most one $5^{-}$-neighbour.
  • $(c)$
    If $m_{3}(v)\geq 3$, then $v$ has neither a $4$-neighbour nor a $5(4^{+})$-neighbour.
Proof.
$(a)$ Let $m_{3}(v)=1$. Clearly, $v$ cannot have any $5(5)$-neighbour since $m_{3}(v)=1$. Now assume that $v$ has two $4$-neighbours. Then $d_{2}(v)\leq 22$ by Lemma 2.2. If we set $G^{\prime}=G-v+\{v_{1}v_{2},v_{1}v_{3},v_{1}v_{4}\}$ for a $4$-neighbour $v_{1}$ of $v$, then $G^{\prime}$ is proper with respect to $G$. By Remark 2.1, this yields a contradiction.
$(b)$ Let $m_{3}(v)=2$. Suppose that $v$ has two neighbours that are $4$- or $5$-vertices. Then $d_{2}(v)\leq 22$ by Lemma 2.2, and the same reduction as in $(a)$ leads to a contradiction.
$(c)$ Let $m_{3}(v)\geq 3$. If $v$ has a $4$-neighbour or a $5(4^{+})$-neighbour, then again $d_{2}(v)\leq 22$ by Lemma 2.2. As in the previous cases, this contradicts Remark 2.1. ∎
Lemma 2.32.
Let $v$ be a $4(1)$-vertex. If $m_{4}(v)\geq 2$, then $v$ has two $7^{+}$-neighbours.
Proof.
Since $m_{3}(v)=1$, the vertex $v$ is incident to a $3$-face $f_{1}=v_{1}vv_{2}$. Assume that $m_{4}(v)\geq 2$. Suppose, for a contradiction, that $v$ has at most one $7^{+}$-neighbour. Then at least one of $v_{1},v_{2}$ must be a $6^{-}$-vertex, say $v_{1}$. In particular, $d_{2}(v)\leq 22$ by Lemma 2.2. If we set $G^{\prime}=G-v+\{v_{1}v_{3},v_{1}v_{4}\}$, then $G^{\prime}$ is proper with respect to $G$, a contradiction to Remark 2.1. ∎
Lemma 2.33.
Let $v$ be a $4(2)$-vertex.
  • $(a)$
    If $m_{4}(v)=0$, then $v$ has two $7^{+}$-neighbours that are neither $7(7)$- nor $8(8)$-vertices.
  • $(b)$
    If $m_{4}(v)=1$, then $v$ has either two $8(7^{-})$-neighbours or three $7^{+}$-neighbours different from $7(7)$- and $8(8)$-vertices.
  • $(c)$
    If $m_{4}(v)=2$, then $v$ has one of the following: one $8(7^{-})$- and three $7(6^{-})$-neighbours, or two $8(7^{-})$- and one $7(6^{-})$-neighbour, or three $8(7^{-})$-neighbours.
Proof.
Let $f_{i}=v_{i}vv_{i+1}$ and $f_{j}=v_{j}vv_{j+1}$ with $i<j$ be the two $3$-faces incident to $v$. Notice first that if $d_{2}(v)\leq 22$, then we set $G^{\prime}=G-v+\{v_{2}v_{4}\}$ when $(i,j)=(1,2)$; we set $G^{\prime}=G-v+\{v_{2}v_{3},\,v_{1}v_{4}\}$ when $(i,j)=(1,3)$. In each case, $G^{\prime}$ is proper with respect to $G$. By Remark 2.1, this yields a contradiction. Thus we further assume that $d_{2}(v)\geq 23$. We also note that if $v$ has a $7(7)$- or $8(8)$-neighbour, then at least two edges in $E(v)$ lie in two $3$-faces. Moreover, $v$ cannot have two neighbours consisting of $7(7)$-, or $8(8)$-vertices since $m_{3}(v)=2$.
$(a)$ Let $m_{4}(v)=0$. Assume for a contradiction that $v$ has at most one $7^{+}$-neighbour that is not $7(7)$- or $8(8)$-vertices; equivalently, $v$ has three neighbours consisting of $6^{-}$-, $7(7)$-, or $8(8)$-vertices. Then $d_{2}(v)\leq 22$ by Lemma 2.2, a contradiction.
$(b)$ Let $m_{4}(v)=1$. First suppose that $v$ has no $7(7)$- or $8(8)$-neighbours. If $v$ has either two $6^{-}$-neighbours and one $7$-neighbour, or three $6^{-}$-neighbours, then $d_{2}(v)\leq 22$ by Lemma 2.2, giving a contradiction. Hence $v$ has either two $8(7^{-})$-neighbours or three $7^{+}$-neighbours. Now suppose $v$ has a $7(7)$- or $8(8)$-neighbour. If $v$ has two $6^{-}$-neighbours, then $d_{2}(v)\leq 21$, and if $v$ has one $6^{-}$- and one $7$-neighbour (other than $7(7)$), then $d_{2}(v)\leq 22$; in both cases we reach a contradiction as before. Thus the statement follows.
$(c)$ Let $m_{4}(v)=2$. We first show that $v$ has at least one $8(7^{-})$-neighbour. Indeed, if all neighbours of $v$ were $7^{-}$- or $8(8)$-vertices, then $d_{2}(v)\leq 22$ by Lemma 2.2, a contradiction. Thus $v$ has at least one $8(7^{-})$-neighbour. If $v$ has exactly one $8(7^{-})$-neighbour and exactly two $7(6^{-})$-neighbours, then $d_{2}(v)\leq 22$. Likewise, if $v$ has exactly two $8(7^{-})$-neighbours and no $7(6^{-})$-neighbours, then again $d_{2}(v)\leq 22$. In each case we obtain a contradiction. Therefore, the remaining possibilities are those listed in the statement: $v$ has either one $8(7^{-})$- and three $7(6^{-})$-neighbours, or two $8(7^{-})$- and one $7(6^{-})$-neighbour, or three $8(7^{-})$-neighbours. ∎
Lemma 2.34.
Let $v$ be a $4(3)$-vertex.
  • $(a)$
    If $m_{4}(v)=0$, then $v$ has one of the following: one $8(7^{-})$- and three $7(6^{-})$-neighbours, or two $8(7^{-})$- and one $7(6^{-})$-neighbour, or two $8(7^{-})$- and two $8(8)$-neighbours, or three $8(7^{-})$-neighbours.
  • $(b)$
    If $m_{4}(v)=1$, then $v$ has either two $8(7^{-})$- and two $7(6^{-})$-neighbours, or three $8(7^{-})$-neighbours.
Proof.
Let $f_{i}=v_{i}vv_{i+1}$ for $i\in[3]$ be $3$-faces incident to $v$. Notice first that if $d_{2}(v)\leq 22$, then $G^{\prime}=G-v+\{v_{1}v_{4}\}$ a graph that is proper with respect to $G$, contradicting Remark 2.1. Thus we assume that $d_{2}(v)\geq 23$.
$(a)$ Let $m_{4}(v)=0$. We first show that $v$ has at least one $8(7^{-})$-neighbour. Indeed, if all neighbours of $v$ were $7^{-}$- or $8(8)$-vertices, then $d_{2}(v)\leq 22$ by Lemma 2.2, a contradiction. Thus $v$ has at least one $8(7^{-})$-neighbour.
Suppose first that $v$ has exactly one $8(7^{-})$-neighbour. If $v$ has no $7(6^{-})$-neighbours, then $d_{2}(v)\leq 22$. If it has exactly one $7(6^{-})$-neighbour and two neighbours consisting of $6^{-}$-, $7(7)$-, or $8(8)$-vertices, then again $d_{2}(v)\leq 22$. Similarly, if $v$ has two $7(6^{-})$-neighbours and one neighbour consisting of $6^{-}$-, $7(7)$-, or $8(8)$-vertices, we again obtain $d_{2}(v)\leq 22$. Each case contradicts Remark 2.1. Hence when $v$ has exactly one $8(7^{-})$-neighbour, it must have three $7(6^{-})$-neighbours.
Suppose next that $v$ has exactly two $8(7^{-})$-neighbours. If $v$ has no $7(6^{-})$-neighbours, then $v$ must have two $8(8)$-neighbours; otherwise $d_{2}(v)\leq 22$. Thus, when $v$ has two $8(7^{-})$-neighbours, it has either one $7(6^{-})$-neighbour or two $8(8)$-neighbours.
$(b)$ Let $m_{4}(v)=1$. If $v$ had at most one $8(7^{-})$-neighbour, then $d_{2}(v)\leq 22$ by Lemma 2.2, yielding a contradiction. Thus $v$ has at least two $8(7^{-})$-neighbours. If $v$ had two neighbours consisting of $6^{-}$-, $7(7)$-, or $8(8)$-vertices, then again $d_{2}(v)\leq 22$. Similarly, if $v$ had two $8(7^{-})$-neighbours, one $7(6^{-})$-neighbour, and one neighbour consisting of $6^{-}$-, $7(7)$-, or $8(8)$-vertices, then $d_{2}(v)\leq 22$. Both situations contradict with our assumption. ∎
Lemma 2.35.
Let $v$ be a $4(4)$-vertex. Then $v$ has either three $8(7^{-})$- and one $7(5^{-})$-neighbour, or four $8(7^{-})$-neighbours.
Proof.
Note that if $d_{2}(v)\leq 22$, then $G^{\prime}=G-v$ is proper with respect to $G$, contradicting Remark 2.1. Thus we assume that $d_{2}(v)\geq 23$. If $v$ has a neighbour consisting of $6^{-}$-, $7(6^{+})$-, or $8(8)$-vertex, then $d_{2}(v)\leq 22$ by Lemma 2.2, a contradiction. Furthermore, if $v$ has two $7^{-}$-neighbours, then $d_{2}(v)\leq 22$ once again. Thus all neighbours of $v$ must be $8(7^{-})$- or $7(5^{-})$-vertices, and at most one neighbour can be of type $7(5^{-})$. Hence $v$ has either three $8(7^{-})$- and one $7(5^{-})$-neighbour, or four $8(7^{-})$-neighbours. ∎
Lemma 2.36.
Let $v$ be a $5(4)$-vertex.
  • $(a)$
    If $m_{4}(v)=0$, then $v$ has two $6^{+}$-neighbours that are not $6(6)$-vertices.
  • $(b)$
    If $m_{4}(v)=1$, then $v$ has three $6^{+}$-neighbours that are not $6(6)$-vertices.
Proof.
Let $f_{i}=v_{i}vv_{i+1}$ for $i\in[4]$.
$(a)$ Assume $m_{4}(v)=0$, and suppose that $v$ has at most one $6^{+}$-neighbour other than a $6(6)$-vertex. Thus all but one neighbour of $v$ are $6(6)$- or $5^{-}$-vertices. Observe that $v$ cannot have four $6(6)$-neighbours. We now check all possibilities: three $6(6)$-neighbours; two $6(6)$- and one $5^{-}$-neighbour; one $6(6)$- and two $5^{-}$-neighbours; or four $5^{-}$-neighbours. In every case $d_{2}(v)\leq 22$ by Lemma 2.2. Setting $G^{\prime}=G-v+\{v_{1}v_{5}\}$ yields a graph that is proper with respect to $G$, contradicting Remark 2.1. Therefore $v$ must have at least two $6^{+}$-neighbours distinct from $6(6)$-vertices.
$(b)$ Assume $m_{4}(v)=1$, and suppose that $v$ has three neighbours $x,y,z$ that are either $5^{-}$- or $6(6)$-vertices. Checking the possibilities (all of $x,y,z$ are $5^{-}$-vertices; two $5^{-}$- and one $6(6)$-vertices; one $5^{-}$- and two $6(6)$-vertices; or all $6(6)$-vertices) gives $d_{2}(v)\leq 22$ in every case by Lemma 2.2. This again contradicts Remark 2.1. Thus $v$ must have at least three $6^{+}$-neighbours that are not $6(6)$-vertices. ∎
Lemma 2.37.
Let $v$ be a $5(5)$-vertex.
  • $(a)$
    If $v$ is adjacent to another $5(5)$-vertex, then $v$ has no $4$- or $5(4)$-neighbours.
  • $(b)$
    $v$ has at most one neighbour that is a $4$- or $5(5)$-vertex.
Proof.
$(a)$ Suppose that $v$ is a $5(5)$-vertex adjacent to a $5(5)$-vertex. If $v$ also has a $4$- or $5(4)$-neighbour, then $d_{2}(v)\leq 21$ by Lemma 2.2. Setting $G^{\prime}=G-v$ yields a graph that is proper with respect to $G$, contradicting Remark 2.1. Thus $v$ has no such neighbours.
$(b)$ Suppose that $v$ has two neighbours consisting of $4$- or $5(5)$-vertices. If one of them is a $5(5)$-vertex, part (a) already gives a contradiction. Otherwise, $v$ has two $4$-neighbours, implying $d_{2}(v)\leq 22$ by Lemma 2.2. As above, this yields a contradiction. Thus $v$ has at most one such neighbour. ∎
Lemma 2.38.
Let $v$ be a $5(5)$-vertex.
  • $(a)$
    If $n_{4}(v)=1$, then $v$ has three $7^{+}$-neighbours that are not $7(7)$-vertices.
  • $(b)$
    If $n_{4}(v)=n_{5}(v)=0$, then $v$ has either three $8(7^{-})$-neighbours or four $6^{+}$-neighbours different from $6(6)$-vertices such that two of which are $7^{+}$-vertices different from $7(7)$-vertex.
  • $(c)$
    If $n_{4}(v)=0$ and $n_{5}(v)=1$, then $v$ has either three $7^{+}$-neighbours different from $7(7)$-vertex or two $6(5^{-})$- and two $7^{+}$-neighbours different from $7(7)$-vertex.
  • $(d)$
    If $n_{4}(v)=0$ and $n_{5}(v)=2$, then $v$ has three $7^{+}$-neighbours different from $7(7)$-vertex.
Proof.
Let $f_{i}=v_{i}vv_{i+1}$ for $i\in[5]$ in cyclic order.
$(a)$ Assume $n_{4}(v)=1$ and that $v$ has at most two $7^{+}$-neighbours different from $7(7)$-vertices. Then $v$ has two neighbours consisting of $5$-, 66-, or $7(7)$-vertices, which forces $d_{2}(v)\leq 22$ by Lemma 2.2. Removing $v$ yields a graph that is proper with respect to $G$, contradicting Remark 2.1. Thus $v$ must have three such $7^{+}$-neighbours.
$(b)$ Let $n_{4}(v)=n_{5}(v)=0$. Assume that $v$ has at most two $8(7^{-})$-neighbours. If $v$ has at most three $6^{+}$-neighbours other than $6(6)$-vertices, then it must have two $6(6)$-neighbours. This forces $d_{2}(v)\leq 22$ by Lemma 2.2, a contradiction. Hence $v$ has four $6^{+}$-neighbours different from $6(6)$-vertices. Among these four, if only one is a $7^{+}$-vertex different from $7(7)$-vertex, then again $d_{2}(v)\leq 22$ by Lemma 2.2, yielding the same contradiction. Therefore at least two must be such $7^{+}$-vertices.
$(c)$ Let $n_{4}(v)=0$ and $n_{5}(v)=1$. If $v$ has at most one $7^{+}$-neighbour different from $7(7)$-vertex, then $d_{2}(v)\leq 22$, a contradiction. Thus $v$ has at least two such $7^{+}$-neighbours.
Assume now that $v$ has exactly two $7^{+}$-neighbours different from $7(7)$-vertex. If $v$ has at most one $6(5^{-})$-neighbour, then again $d_{2}(v)\leq 22$ by Lemma 2.2. Thus $v$ must have two $6(5^{-})$-neighbours in this configuration.
$(d)$ Let $n_{4}(v)=0$ and $n_{5}(v)=2$. If $v$ has at most two $7^{+}$-neighbours different from $7(7)$-vertices, then $d_{2}(v)\leq 22$, again contradicting Remark 2.1. Hence $v$ has three such neighbours. ∎
Lemma 2.39.
Let $v$ be a $6(5)$-vertex. Then $v$ has at most four $5(4^{+})$-neighbours. In particular, if $m_{4}(v)=1$ and $v$ has four $5(4^{+})$-neighbours, then $v$ has two $8(6^{-})$-neighbours.
Proof.
Let $f_{i}=v_{i}vv_{i+1}$ for $i\in[5]$. Assume first that $v$ has five $5(4^{+})$-neighbours. Then at least three edges of $E(v)$ are contained in two $3$-faces, and Lemma 2.2 gives $d_{2}(v)\leq 20$. Since $v$ has five $5$-neighbours, one of $v_{1},v_{6}$ must be a $5$-vertex, say $v_{1}$, and one of $v_{3},v_{5}$ is also a $5$-vertex, say $v_{3}$. Let $G^{\prime}=G-v+\{\,v_{1}v_{6},\ v_{3}v_{1},\ v_{3}v_{5}\,\}$. Clearly, $G^{\prime}$ is proper with respect to $G$, contradicting Remark 2.1. Thus, $v$ has at most four $5(4^{+})$-neighbours.
Now suppose that $m_{4}(v)=1$ and that $v$ has four $5(4^{+})$-neighbours. Then there are at least two edges in $E(v)$ each contained in two $3$-faces. If $v$ has at most one $8(6^{-})$-neighbour, then $d_{2}(v)\leq 22$ by Lemma 2.2, and the same reduction as above yields a contradiction. Hence $v$ must have at least two $8(6^{-})$-neighbours. ∎
Let $v$ be a $7$-vertex (resp. an $8$-vertex), and let $x$ be a $5^{-}$-neighbour (resp. a $4^{-}$-neighbour) of $v$. If $v$ and $x$ have two common neighbours, (equivalently, if the edge $vx$ lies in two $3$-faces), then $x$ is called a support neighbour of $v$.
Lemma 2.40.
Let $v$ be a $7^{+}$-vertex.
  • $(a)$
    If $v$ has a support neighbour, then $d_{2}(v)\geq 23$.
  • $(b)$
    If $v$ is a $7(6)$-vertex, and has five $4(1^{+})$-neighbours, then $v$ cannot have any $5$-neighbour.
  • $(c)$
    If $v$ is a $7(6)$-vertex, and has a $5(5)$-neighbour, then $v$ has at most four neighbours consisting of $4$- or $5(4)$-vertices.
  • $(d)$
    If $v$ is a $7(6)$-vertex, and has two $5(5)$-neighbours, then $v$ has at most three neighbours consisting of $4$- or $5(4)$-vertices.
  • $(e)$
    If $v$ is a $7(6)$-vertex, and has three $5(5)$-neighbours, then $v$ has at most one $4$-neighbour. In particular, $v$ has at most two neighbours consisting of $4$- or $5(4)$-vertices.
  • $(f)$
    If $v$ is a $7(7)$-vertex, then $v$ has at most five $5(4^{+})$-neighbours.
  • $(g)$
    If $v$ is a $8(7)$-vertex, and has a support neighbour, then $v$ has at most six neighbours consisting of $4(2^{-})$- or $5(5)$-vertices.
Proof.
$(a)$ Let $d(v)=k$ and suppose that $v_{1}$ is a support neighbour of $v$. By definition, $v_{1}$ is adjacent to both $v_{2}$ and $v_{k}$. Assume for a contradiction that $d_{2}(v)\leq 22$. Consider $G^{\prime}=G-v+\{v_{1}v_{3},v_{1}v_{4},\ldots,v_{1}v_{k-1}\}.$ The graph $G^{\prime}$ is proper with respect to $G$, contradicting Remark 2.1. Thus, $d_{2}(v)\geq 23$.
$(b)$ Suppose that $v$ is a $7(6)$-vertex, and $v$ has five $4(1^{+})$-neighbours. Clearly, $v$ has a support neighbour. If $v$ had a $5$-neighbour, then it would be $d_{2}(v)\leq 21$, contradicting (a).
$(c)$ Suppose that $v$ is a $7(6)$-vertex, and $v$ has a $5(5)$-neighbour. Clearly, $v$ has a support neighbour. If $v$ had five neighbours consisting of $4$- or $5(4)$-vertices, then it would be $d_{2}(v)\leq 22$, contradicting (a).
$(d)$ Suppose that $v$ is a $7(6)$-vertex with two $5(5)$-neighbours. Clearly, $v$ has a support neighbour. If $v$ had four neighbours consisting of $4$- or $5(4)$-vertices, then $d_{2}(v)\leq 22$, contradicting part (a).
$(e)$ Suppose that $v$ is a $7(6)$-vertex with three $5(5)$-neighbours. Clearly, $v$ has a support neighbour. Since a $5(5)$-vertex is adjacent to at most one $5(5)$-vertex by Lemma 2.37(b), $v$ can have at most one $4$-neighbour; otherwise $d_{2}(v)\leq 22$, contradicting (a). Similarly, $v$ cannot have three neighbours from $\{4\text{-vertices},\,5(4)\text{-vertices}\}$, as this would again imply $d_{2}(v)\leq 22$, contradicting (a).
$(f)$ Assume, for a contradiction, that $v$ has six $5(4^{+})$-neighbours. Clearly, at least one of such $5(4^{+})$-neighbour is a support neighbour of $v$, and Lemma 2.2 then gives $d_{2}(v)\leq 21$, contradicting (a).
$(g)$ Suppose that $v$ has seven such neighbours. Then Lemma 2.2 yields $d_{2}(v)\leq 22$, again contradicting (a). ∎
We now apply discharging to show that $G$ does not exist. We use the same initial charges as before, together with the following discharging rules.
Discharging Rules
We apply the following discharging rules.
  • R1:
    Every $3$-face receives $\frac{1}{3}$ from each of its incident vertices.
  • R2:
    Every $5^{+}$-face gives $\frac{1}{5}$ to each of its incident vertices.
  • R3:
    Every $3$-vertex receives $\frac{1}{5}$ from each of its $8(6^{-})$-neighbour.
  • R4:
    Every $4(4)$-vertex receives $\frac{1}{3}$ from each of its $7(5^{-})$-neighbours.
  • R5:
    Every $6(5)$-vertex receives $\frac{1}{18}$ from each of its $8(6^{-})$-neighbours.
  • R6:
    Every $6(5^{-})$-vertex gives $\frac{1}{9}$ to each of its bad $5$-neighbours.
  • R7:
    Let $v$ be a $7(6^{-})$-vertex. Then, $v$ gives
    • $(a)$
      $\frac{1}{6}$ to each of its $4(1)$-, $4(2)$- and $4(3)$-neighbours,
    • $(b)$
      $\frac{1}{9}$ to each of its $5(4)$-neighbours,
    • $(c)$
      $\frac{2}{9}$ to each of its $5(5)$-neighbours,
  • R8:
    Every $7(7)$-vertex gives $\frac{1}{9}$ to each of its bad $5$-neighbours,
  • R9:
    Let $v$ be a $8(7^{-})$-vertex. Then, $v$ gives
    • $(a)$
      $\frac{1}{6}$ to each of its $4(1)$-neighbours,
    • $(b)$
      $\frac{1}{4}$ to each of its $4(2)$-neighbours,
    • $(c)$
      $\frac{1}{3}$ to each of its $4(3^{+})$-neighbours,
    • $(d)$
      $\frac{1}{9}$ to each of its $5(4)$-neighbours,
    • $(e)$
      $\frac{2}{9}$ to each of its $5(5)$-neighbours,
  • R10:
    Let $v$ be an $8(8)$-vertex. Then, $v$ gives
    • $(a)$
      $\frac{1}{9}$ to each of its $4(3)$-, and 5(4)5(4)–neighbours,
    • $(b)$
      $\frac{2}{9}$ to each of its $5(5)$-neighbours,
Checking $\mu^{*}(v),\mu^{*}(f)\geq 0$ for $v\in V(G),f\in F(G)$.
First we show that $\mu^{*}(f)\geq 0$ for each $f\in F(G)$. Given a face $f\in F(G)$, if $f$ is a $3$-face, then it receives $\frac{1}{3}$ from each of its incident vertices by R1, and so $\mu^{*}(f)=\ell(f)-4+3\times\frac{1}{3}=0$. If $f$ is a $4$-face, then $\mu(f)=\mu^{*}(f)=0$. Let $f$ be a $5^{+}$-face. By applying R2, $f$ sends $\frac{1}{5}$ to each of its incident vertices. It then follows that $\mu^{*}(f)\geq\frac{4\ell(f)}{5}-4\geq 0$. Consequently, $\mu^{*}(f)\geq 0$ for each $f\in F(G)$.
We now pick a vertex $v\in V(G)$ with $d(v)=k$. By Lemma 2.29, we have $k\geq 3$.
(1). Let $k=3$. The initial charge of $v$ is $\mu(v)=d(v)-4=-1$. By Lemma 2.30, $m_{3}(v)=0$ and $m_{4}(v)\leq 1$. If $m_{4}(v)=0$, then $v$ is adjacent to two $8(6^{-})$-vertices by Lemma 2.30(b), and if $m_{4}(v)=1$, then $v$ is adjacent to three $8(6^{-})$-neighbours by Lemma 2.30(c). Thus $\mu^{*}(v)\geq-1+3\times\frac{1}{5}+2\times\frac{1}{5}=0$ after $v$ receives $\frac{1}{5}$ from each of its incident $5^{+}$-faces by R2, and $\frac{1}{5}$ from each of its $8(6^{-})$-neighbours by R3.
(2). Let $k=4$. The initial charge of $v$ is $\mu(v)=d(v)-4=0$. If $m_{3}(v)=0$, then $\mu^{*}(v)\geq 0$, since $v$ does not give any charge to its incident faces. Therefore, we may assume that $1\leq m_{3}(v)\leq 4$.
(2.1). Let $m_{3}(v)=1$. If $m_{4}(v)\leq 1$, then $v$ is incident to two $5^{+}$-faces, and by R2, $v$ receives $\frac{1}{5}$ from each of those $5^{+}$-faces. Thus, $\mu^{*}(v)\geq 2\times\frac{1}{5}-\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to its incident $3$-face by R1. If $m_{4}(v)\geq 2$, then $v$ has two $7^{+}$-neighbours by Lemma 2.32, which are clearly different from $7(7)$- and $8(8)$-vertices as $m_{3}(v)=1$. It follows that $v$ receives $\frac{1}{6}$ from each of its $7(6^{-})$- and $8(7^{-})$-neighbours by R7(a), R9(a). Thus, $\mu^{*}(v)\geq 2\times\frac{1}{6}-\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to its incident $3$-face by R1.
(2.2). Let $m_{3}(v)=2$. Since $0\leq m_{4}(v)\leq 2$, we consider the following cases:
If $m_{4}(v)=0$, then $v$ has two $7^{+}$-neighbours different from $7(7)$- and $8(8)$-vertices by Lemma 2.33(a). It follows that $v$ receives totally at least $2\times\frac{1}{6}$ from its $7(6^{-})$- and $8(7^{-})$-neighbours by R7(a), R9(b). In addition, $v$ is incident to two $5^{+}$-faces, and by R2, $v$ receives $\frac{1}{5}$ from each of those $5^{+}$-faces. Thus, $\mu^{*}(v)\geq 2\times\frac{1}{6}+2\times\frac{1}{5}-2\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
If $m_{4}(v)=1$, then $v$ has either two $8(7^{-})$-neighbours or three $7^{+}$-neighbours different from $7(7)$- and $8(8)$-vertices by Lemma 2.33(b). In each case, $v$ receives totally at least $\frac{1}{2}$ from its $7(6^{-})$- and $8(7^{-})$-neighbours by R7(a), R9(b). Also, $v$ receives $\frac{1}{5}$ from its incident $5^{+}$-face by R2. Thus, $\mu^{*}(v)\geq\frac{1}{2}+\frac{1}{5}-2\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
If $m_{4}(v)=2$, then $v$ has either one $8(7^{-})$- and three $7(6^{-})$-neighbours or two $8(7^{-})$- and one $7(6^{-})$-neighbours or three $8(7^{-})$-neighbours by Lemma 2.33(c). In each case, $v$ receives totally at least $\frac{2}{3}=\min\{\frac{1}{4}+3\times\frac{1}{6},2\times\frac{1}{4}+\frac{1}{6},3\times\frac{1}{4}\}$ from its $7(6^{-})$- and $8(7^{-})$-neighbours by R7(a), R9(b). Thus, $\mu^{*}(v)\geq\frac{2}{3}-2\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
(2.3). Let $m_{3}(v)=3$. Suppose first that $m_{4}(v)=0$. It then follows from Lemma 2.34(a) that $v$ has either one $8(7^{-})$- and three $7(6^{-})$-neighbours or two $8(7^{-})$- and one $7(6^{-})$-neighbours or two $8(7^{-})$- and two $8(8)$-neighbours or three $8(7^{-})$-neighbours. In each case, $v$ receives totally at least $\frac{5}{6}=\min\{\frac{1}{3}+3\times\frac{1}{6},2\times\frac{1}{3}+\frac{1}{6},2\times\frac{1}{3}+2\times\frac{1}{9},3\times\frac{1}{3}\}$ from its $7(6^{-})$- and $8$-neighbours by R7(a), R9(c), R10(a). In addition, $v$ receives $\frac{1}{5}$ from its incident $5^{+}$-face by R2. Thus, $\mu^{*}(v)\geq\frac{5}{6}+\frac{1}{5}-3\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1. Now, we suppose that $m_{4}(v)=1$. By Lemma 2.34(b), $v$ has either two $8(7^{-})$- and two $7(6^{-})$-neighbours or three $8(7^{-})$-neighbours. In both cases, $v$ receives totally at least $1=\min\{2\times\frac{1}{3}+2\times\frac{1}{6},3\times\frac{1}{3}\}$ from its $7(6^{-})$- and $8(7^{-})$-neighbours by R7(a), R9(c). Thus, $\mu^{*}(v)\geq 1-3\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
(2.4). Let $m_{3}(v)=4$. By Lemma 2.35, $v$ has either three $8(7^{-})$- and one $7(5^{-})$-neighbour or four $8(7^{-})$-neighbours. By applying R4 and R9(c), $v$ receives $\frac{1}{3}$ from each of its $7(5^{-})$- and $8(7^{-})$-neighbours. Thus, $\mu^{*}(v)\geq 4\times\frac{1}{3}-4\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
(3). Let $k=5$. The initial charge of $v$ is $\mu(v)=d(v)-4=1$. If $m_{3}(v)\leq 3$, then $\mu^{*}(v)\geq 1-3\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1. Therefore, we may assume that $4\leq m_{3}(v)\leq 5$.
(3.1). Let $m_{3}(v)=4$. Suppose first that $m_{4}(v)=0$. By Lemma 2.36(a), $v$ has two $6^{+}$-neighbours different from $6(6)$-vertex. It follows that $v$ receives $\frac{1}{9}$ from each of its $6^{+}$-neighbours different from $6(6)$-vertex by R6-R10, and $\frac{1}{5}$ from its incident $5^{+}$-face by R2. Thus, $\mu^{*}(v)\geq 1+2\times\frac{1}{9}+\frac{1}{5}-4\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1. Next we suppose that $m_{4}(v)=1$. By Lemma 2.36(b), $v$ has three $6^{+}$-neighbours different from $6(6)$-vertex. So, $v$ receives $\frac{1}{9}$ from each of its $6^{+}$-neighbours different from $6(6)$-vertex by R6-R10. Thus, $\mu^{*}(v)\geq 1+3\times\frac{1}{9}-4\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
(3.2). Let $m_{3}(v)=5$. We distinguish two subcases according to the number of $4$-vertices adjacent to $v$ as follows. Recall that $0\leq n_{4}(v)\leq 1$ by Lemma 2.37(b).
(3.2.1). Let $n_{4}(v)=0$. If $n_{5}(v)=0$, then, by Lemma 2.38(b), $v$ has either three $8(7^{-})$-neighbours or four $6^{+}$-neighbours different from $6(6)$-vertex, two of which are $7^{+}$-vertices different from $7(7)$-vertex. It follows from applying R6-R10 that $v$ receives totally at least $\frac{2}{3}=\min\{3\times\frac{2}{9},2\times\frac{1}{9}+2\times\frac{2}{9}\}$ from its $6^{+}$-neighbours. Thus, $\mu^{*}(v)\geq 1+\frac{2}{3}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
If $n_{5}(v)=1$, then, by Lemma 2.38(c), $v$ has either three $7^{+}$-neighbours different from $7(7)$-vertex or two $6(5^{-})$- and two $7^{+}$-neighbours different from $7(7)$-vertex. So, $v$ receives totally at least $\frac{2}{3}=\min\{3\times\frac{2}{9},2\times\frac{1}{9}+2\times\frac{2}{9}\}$ from its $6^{+}$-neighbours by R6-R10. Thus, $\mu^{*}(v)\geq 1+\frac{2}{3}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
If $n_{5}(v)=2$, then $v$ has three $7^{+}$-neighbours different from $7(7)$-vertex by Lemma 2.38(d). It follows that $v$ receives $\frac{2}{9}$ from each of its $7^{+}$-neighbours different from $7(7)$-vertex by R7(c), R9(e), R10(b). Thus, $\mu^{*}(v)\geq 1+3\times\frac{2}{9}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
(3.2.2). Let $n_{4}(v)=1$. By Lemma 2.38(a), $v$ has three $7^{+}$-neighbours different from $7(7)$-vertex. It follows that $v$ receives $\frac{2}{9}$ from each of its $7^{+}$-neighbours different from $7(7)$-vertex by R7(c), R9(e), R10(b). So, $\mu^{*}(v)\geq 1+3\times\frac{2}{9}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
(4). Let $k=6$. The initial charge of $v$ is $\mu(v)=d(v)-4=2$. If $m_{3}(v)\leq 4$, then $\mu^{*}(v)\geq 2-4\times\frac{1}{3}-6\times\frac{1}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and $\frac{1}{9}$ to each of its $5(4^{+})$-neighbours by R6. On the other hand, if $m_{3}(v)=6$, then $\mu^{*}(v)\geq 2-6\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1. Therefore, we may assume that $m_{3}(v)=5$. By Lemma 2.39, $v$ has at most four $5(4^{+})$-neighbours. If $v$ has at most three $5(4^{+})$-neighbours, then $\mu^{*}(v)\geq 2-5\times\frac{1}{3}-3\times\frac{1}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and $\frac{1}{9}$ to each of its $5(4^{+})$-neighbours by R6. Next we assume that $v$ has exactly four $5(4^{+})$-neighbours. If $m_{4}(v)=0$, then $v$ receives $\frac{1}{5}$ from its incident $5^{+}$-face by R2, and so $\mu^{*}(v)\geq 2+\frac{1}{5}-5\times\frac{1}{3}-4\times\frac{1}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and $\frac{1}{9}$ to each of its $5(4^{+})$-neighbours by R6. If $m_{4}(v)=1$, then $v$ has two $8(6^{-})$-neighbours by Lemma 2.39, and so $v$ receives $\frac{1}{18}$ from each of its $8(6^{-})$-neighbours by R5. Thus, $\mu^{*}(v)\geq 2+2\times\frac{1}{18}-5\times\frac{1}{3}-4\times\frac{1}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and $\frac{1}{9}$ to each of its $5(4^{+})$-neighbours by R6.
(5). Let $k=7$. The initial charge of $v$ is $\mu(v)=d(v)-4=3$. Notice first that if $m_{3}(v)\leq 3$, then we have $\mu^{*}(v)\geq 3-3\times\frac{1}{3}-2\times\frac{1}{3}-5\times\frac{2}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its $4(4)$-neighbours by R4, and at most $\frac{2}{9}$ to each of its other neighbours by R7, where we note that $v$ can have at most two $4(4)$-neighbours due to $m_{3}(v)\leq 3$. Therefore, we may assume that $4\leq m_{3}(v)\leq 7$.
(5.1). Let $m_{3}(v)=4$. Observe that $v$ can have at most two neighbours consisting of $4(4)$- or $5(5)$-vertices by Lemmas 2.31(c) and 2.37(b). If $v$ has at most one $4(4)$-neighbour, then $\mu^{*}(v)\geq 3-4\times\frac{1}{3}-\frac{1}{3}-6\times\frac{2}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to its $4(4)$-neighbour by R4, and at most $\frac{2}{9}$ to each of its other neighbours by R7. If $v$ has exactly two $4(4)$-neighbours, then $v$ is not adjacent to any $5(5)$-vertex as stated earlier, and so $\mu^{*}(v)\geq 3-4\times\frac{1}{3}-2\times\frac{1}{3}-5\times\frac{1}{6}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its $4(4)$-neighbours by R4, and at most $\frac{1}{6}$ to each of its other neighbours by R7(a)-(b).
(5.2). Let $m_{3}(v)=5$. By Lemma 2.31(c), $v$ has at most two $4(4)$-neighbours. If $v$ has no $4(4)$-neighbour, then $v$ is adjacent to at most three $5(5)$-neighbour by Lemma 2.37(b), and so $\mu^{*}(v)\geq 3-5\times\frac{1}{3}-3\times\frac{2}{9}-4\times\frac{1}{6}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{2}{9}$ to each of its $5(5)$-neighbours by R7(c), and at most $\frac{1}{6}$ to each of its other neighbours by R7(a)-(b). If $v$ has exactly one $4(4)$-neighbour, say $v_{i}$, then $v$ is adjacent to at most two $5(5)$-neighbours by Lemmas 2.31(c) and 2.37(b). Moreover, the neighbours of $v_{i}$ belonging to $G[N(v)]$, say $v_{j}$ and $v_{k}$, are neither $4$-vertices nor $5(4^{+})$-vertices by Lemma 2.31(c). Clearly, neither $v_{j}$ nor $v_{k}$ receives any charge from $v$. Therefore, $\mu^{*}(v)\geq 3-5\times\frac{1}{3}-\frac{1}{3}-2\times\frac{2}{9}-2\times\frac{1}{6}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to its $4(4)$-neighbour by R4, $\frac{2}{9}$ to each of its $5(5)$-neighbours by R7(c), and at most $\frac{1}{6}$ to each of its $4(1^{+})$- and $5(4)$-neighbours different from $4(4)$-vertex by R7(a)-(b). If $v$ has exactly two $4(4)$-neighbours, then $v$ is not adjacent to any $5(5)$-vertices by Lemma 2.31(c). Moreover, similarly as above, $v$ has at least three neighbours $v_{i}$, $v_{j}$, and $v_{k}$ that are neither $4$-vertices nor $5(4^{+})$-vertices by Lemma 2.31(c), since $v$ has two $4(4)$-neighbors. Clearly, none of $v_{i}$, $v_{j}$, or $v_{k}$ receives any charge from $v$. Thus, $\mu^{*}(v)\geq 3-5\times\frac{1}{3}-2\times\frac{1}{3}-2\times\frac{1}{6}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its $4(4)$-neighbours by R4, and at most $\frac{1}{6}$ to each of its $4(1^{+})$- and $5(4)$-neighbours different from $4(4)$-vertex by R7(a)-(b).
(5.3). Let $m_{3}(v)=6$. Notice first that if $v$ has at most four neighbours forming $4$- or $5$-vertices, then we have $\mu^{*}(v)\geq 3-6\times\frac{1}{3}-4\times\frac{2}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and at most $\frac{2}{9}$ to each of its $4(1)$-, $4(2)$-, $4(3)$-, and $5(4^{+})$-neighbours by R7. Therefore we may further assume that $v$ has at least five neighbours forming $4$- or $5$-vertices. By Lemma 2.31, we conclude that a $4(1^{+})$-vertex has at most one $4$-neighbour. In addition, a $5(5)$-vertex has at most one neighbour consisting of $4$- or $5(5)$-vertices by Lemma 2.37(b). This implies that $v$ has at most four $5(5)$-neighbours; in particular, $v$ has at most five $4(1^{+})$-neighbours.
Suppose first that $v$ has no $5(5)$-neighbours. If $v$ has at most four $4(1^{+})$-neighbours, then $\mu^{*}(v)\geq 3-6\times\frac{1}{3}-4\times\frac{1}{6}-3\times\frac{1}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{6}$ to each of its $4(1^{+})$-neighbours by R7(a), and $\frac{1}{9}$ to each of its $5(4)$-neighbours by R7(b). If $v$ has five $4(1^{+})$-neighbours, then $v$ cannot have any $5(4)$-neighbours by Lemma 2.40(b). So, $\mu^{*}(v)\geq 3-6\times\frac{1}{3}-5\times\frac{1}{6}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and $\frac{1}{6}$ to each of its $4(1^{+})$-neighbours by R7(a).
Next we suppose that $v$ has exactly one $5(5)$-neighbour. By Lemma 2.40(c), $v$ has at most four $4$-neighbours. If $v$ has at most two $4$-neighbours, then $\mu^{*}(v)\geq 3-6\times\frac{1}{3}-\frac{2}{9}-2\times\frac{1}{6}-4\times\frac{1}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{2}{9}$ to its $5(5)$-neighbour by R7(c), $\frac{1}{6}$ to each of its $4(1^{+})$-neighbours by R7(a), and $\frac{1}{9}$ to each of its $5(4)$-neighbours by R7(b). If $v$ has at least three $4$-neighbours, then $v$ is adjacent to at most one $5(4)$-vertices by Lemma 2.40(c), and so $\mu^{*}(v)\geq 3-6\times\frac{1}{3}-\frac{2}{9}-4\times\frac{1}{6}-\frac{1}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{2}{9}$ to its $5(5)$-neighbour by R7(c), $\frac{1}{6}$ to each of its $4(1^{+})$-neighbours by R7(a), and $\frac{1}{9}$ to each of its $5(4)$-neighbours by R7(b).
Suppose now that $v$ has exactly two $5(5)$-neighbours. Then, by Lemma 2.40(d), $v$ has at most three neighbours consisting of $4$- or $5(4)$-vertices, and so $\mu^{*}(v)\geq 3-6\times\frac{1}{3}-2\times\frac{2}{9}-3\times\frac{1}{6}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{2}{9}$ to each of its $5(5)$-neighbours by R7(c), and at most $\frac{1}{6}$ to each of its $4(1^{+})$ and $5(4)$-neighbours by R7(a),(b).
Suppose that $v$ has exactly three $5(5)$-neighbours. It follows that $v$ has at most one $4$-neighbour by Lemma 2.40(e). By the same reason, $v$ has at most two neighbours consisting of $4$- or $5(4)$-vertices. Thus, we have $\mu^{*}(v)\geq 3-6\times\frac{1}{3}-3\times\frac{2}{9}-\frac{1}{6}-\frac{1}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{2}{9}$ to its $5(5)$-neighbour by R7(c), $\frac{1}{6}$ to each of its $4(1^{+})$-neighbours by R7(a), and $\frac{1}{9}$ to each of its $5(4)$-neighbours by R7(b).
Finally, suppose that $v$ has four $5(5)$-neighbours. Then $v$ has neither $4$- nor $5(4)$-neighbours by Lemma 2.37(a),(b), and so $\mu^{*}(v)\geq 3-6\times\frac{1}{3}-4\times\frac{2}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{2}{9}$ to each of its $5(5)$-neighbours by R7(c).
(5.4). Let $m_{3}(v)=7$. By Lemma 2.40(f), $v$ has at most five $5(4^{+})$-neighbours, and so $\mu^{*}(v)\geq 3-7\times\frac{1}{3}-5\times\frac{1}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-face by R1, $\frac{1}{9}$ to each of its $5(4^{+})$-neighbour by R8.
(6). Let $k=8$. The initial charge of $v$ is $\mu(v)=d(v)-4=4$. Notice first that if $m_{3}(v)\leq 4$, then we have $\mu^{*}(v)\geq 4-4\times\frac{1}{3}-8\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and at most $\frac{1}{3}$ to each of its neighbours by R3, R5, R9. Therefore, we may assume that $5\leq m_{3}(v)\leq 8$.
(6.1). Let $m_{3}(v)=5$. By Lemma 2.31(c), a $4(3^{+})$-vertex cannot be adjacent to any $4$-vertex. It follows that $v$ has at most four $4(3^{+})$-neighbours. Then $\mu^{*}(v)\geq 4-5\times\frac{1}{3}-4\times\frac{1}{3}-4\times\frac{1}{4}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its $4(3^{+})$-neighbours by R9(c), and at most $\frac{1}{4}$ to each of its other neighbours by R3, R5, R9.
(6.2). Let $m_{3}(v)=6$. By Lemma 2.31(c), $v$ has at most four $4(3^{+})$-neighbours. Suppose first that $v$ has no $4(3^{+})$-neighbour. Then $\mu^{*}(v)\geq 4-6\times\frac{1}{3}-8\times\frac{1}{4}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and at most $\frac{1}{4}$ to each of its neighbours by R3, R5, R9.
Next suppose that $v$ has exactly one $4(3^{+})$-neighbour. It follows from Lemmas 2.30(a) and 2.31(c) that $v$ has at most six neighbours consisting of $3$-, $4(2^{-})$- or $5(4^{+})$-vertices. Thus $\mu^{*}(v)\geq 4-6\times\frac{1}{3}-\frac{1}{3}-6\times\frac{1}{4}-\frac{1}{18}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to its $4(3^{+})$-neighbour by R9(c), at most $\frac{1}{4}$ to each of its $3$-, $4(1)$-, $4(2)$- and $5(4^{+})$-neighbours by R3, R9, and $\frac{1}{18}$ to each of its $6(5)$-neighbours by R5.
Suppose now that $v$ has exactly two $4(3^{+})$-neighbours. In such a case, $v$ has at most five neighbours consisting of $3$-, $4(2^{-})$- or $5(4^{+})$-vertices by Lemmas 2.30(a) and 2.31(c). Thus $\mu^{*}(v)\geq 4-6\times\frac{1}{3}-2\times\frac{1}{3}-5\times\frac{1}{4}-\frac{1}{18}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its $4(3^{+})$-neighbours by R9(c), at most $\frac{1}{4}$ to each of its $3$-, $4(1)$-, $4(2)$- and $5(4^{+})$-neighbours by R3, R9, and $\frac{1}{18}$ to each of its $6(5)$-neighbours by R5.
Suppose that $v$ has exactly three $4(3^{+})$-neighbours. In such a case, $v$ can have at most three neighbours consisting of $3$-, $4(2^{-})$- or $5(4^{+})$-vertices by Lemmas 2.30(a) and 2.31(c). Thus $\mu^{*}(v)\geq 4-6\times\frac{1}{3}-3\times\frac{1}{3}-3\times\frac{1}{4}-2\times\frac{1}{18}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its $4(3^{+})$-neighbours by R9(c), at most $\frac{1}{4}$ to each of its $3$-, $4(1)$-, $4(2)$- and $5(4^{+})$-neighbours by R3, R9, and $\frac{1}{18}$ to each of its $6(5)$-neighbours by R5.
Finally suppose that $v$ has exactly four $4(3^{+})$-neighbours. In such a case, $v$ can have at most one neighbour consisting of $3$-, $4(2^{-})$- or $5(4^{+})$-vertices by Lemmas 2.30(a) and 2.31(c). Thus $\mu^{*}(v)\geq 4-6\times\frac{1}{3}-4\times\frac{1}{3}-\frac{1}{4}-3\times\frac{1}{18}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its $4(3^{+})$-neighbours by R9(c), and at most $\frac{1}{4}$ to its $3$-, $4(1)$-, $4(2)$- and $5(4^{+})$-neighbours by R3, R9, and $\frac{1}{18}$ to each of its $6(5)$-neighbours by R5.
(6.3). Let $m_{3}(v)=7$. Let $f_{i}=v_{i}vv_{i+1}$ for $i\in[7]$. Clearly, $v$ has no $3$-neighbour by Lemma 2.30.
Let us first claim that $v$ has a support neighbour. By contradiction, assume that $v$ does not have. Then, $v_{2},v_{3},\ldots,v_{7}$ are $5^{+}$-vertices. Recall that a $5(5)$-vertex has at most one $5(5)$-neighbour by Lemma 2.37(b). Thus we infer that $v$ has at most four $5(5)$-neighbour. If one of $v_{1},v_{8}$ is a $5^{+}$-vertex, then $\mu^{*}(v)\geq 4-7\times\frac{1}{3}-\frac{1}{3}-4\times\frac{2}{9}-3\times\frac{1}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, at most $\frac{1}{3}$ to each of its $4$-neighbours by R9(a)-(c), $\frac{2}{9}$ to each of its $5(5)$-neighbours by R9(e), $\frac{1}{9}$ to each of its $5(4)$-neighbours by R9(d). Thus, we assume that $v_{1}$ and $v_{8}$ are $4$-vertices. By Lemma 2.31, this infer that $v$ has at most three $5(5)$-neighbours. It then follows that $\mu^{*}(v)\geq 4-7\times\frac{1}{3}-2\times\frac{1}{3}-3\times\frac{1}{9}-3\times\frac{2}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, at most $\frac{1}{3}$ to each of its $4$-neighbours by R9(a)-(c), $\frac{1}{9}$ to each of its $5(4)$-neighbours by R9(d), $\frac{2}{9}$ to each of its $5(5)$-neighbours by R9(e). Consequently, we deduce that $v$ has a support neighbour, i.e., one of $v_{2},v_{3},\ldots,v_{7}$ is a $4^{-}$-vertex.
By Lemma 2.31(c), a $4(3^{+})$-vertex has no $4$- and $5(4^{+})$-neighbours. So, we deduce that $v$ has at most four $4(3^{+})$-neighbours. Indeed, if $v$ has four such neighbours, then $v$ is not adjacent to any $4(2^{-})$- or $5(4^{+})$-vertices by Lemma 2.31(c). Thus $\mu^{*}(v)\geq 4-7\times\frac{1}{3}-4\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its $4(3^{+})$-neighbours by R9(c). We may further assume that $v$ has at most three $4(3^{+})$-neighbours.
Suppose first that $v$ has no $4(3^{+})$-neighbour. Note that $v$ has at most six neighbours consisting of $4(2^{-})$- or $5(5)$-vertices by Lemma 2.40(g). Moreover, if $v$ has six such neighbours, then $v$ cannot be adjacent to any $5(4)$-vertex by Lemmas 2.31 and 2.37, and so $\mu^{*}(v)\geq 4-7\times\frac{1}{3}-6\times\frac{1}{4}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, at most $\frac{1}{4}$ to each of its $4(1)$-, $4(2)$- and $5(5)$-neighbours by R9(a), (b), (e). If $v$ has at most five neighbours consisting of $4(2^{-})$- or $5(5)$-vertices, then $\mu^{*}(v)\geq 4-7\times\frac{1}{3}-5\times\frac{1}{4}-3\times\frac{1}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, at most $\frac{1}{4}$ to each of its $4(1)$-, $4(2)$- and $5(5)$-neighbours by R9(a),(b),(e), and $\frac{1}{9}$ to each of its $5(4)$-neighbours by R9(d).
Next we suppose that $v$ has exactly one $4(3^{+})$-neighbour. By Lemma 2.31(c), $v$ has at least one neighbour different from $4$- and $5(4^{+})$-vertices, which does not receive any charge from $v$. On the other hand, a $4(t)$-vertex with $1\leq t\leq 2$ has at most one neighbour consisting of $4$- or $5(5)$-vertices by Lemma 2.31(a),(b). It then follows from Lemma 2.31(c) and Lemma 2.37(b) that $v$ has at most four neighbours consisting of $4(1)$-, $4(2)$- or $5(5)$-vertices. Thus $\mu^{*}(v)\geq 4-7\times\frac{1}{3}-\frac{1}{3}-4\times\frac{1}{4}-2\times\frac{1}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its $4(3^{+})$-neighbours by R9(c), at most $\frac{1}{4}$ to each of its $4(1)$-, $4(2)$- and $5(5)$-neighbours by R9(a),(b),(e), and $\frac{1}{9}$ to each of its $5(4)$-neighbours by R9(d).
Suppose that $v$ has exactly two $4(3^{+})$-neighbours. By Lemma 2.31(c), a $4(3^{+})$-vertex has neither $4$-neighbour nor $5(4^{+})$-neighbour. This means that $v$ has at least two neighbours different from $4$- and $5(4^{+})$-vertices, which do not receive any charge from $v$. Thus $\mu^{*}(v)\geq 4-7\times\frac{1}{3}-2\times\frac{1}{3}-4\times\frac{1}{4}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its $4(3^{+})$-neighbours by R9(c), and at most $\frac{1}{4}$ to each of its $4(1)$-, $4(2)$- and $5(4^{+})$-neighbours by R9(a),(b),(d),(e).
Finally suppose that $v$ has exactly three $4(3^{+})$-neighbours. It follows from Lemma 2.31(c) that $v$ has at least three neighbours different from $4$- and $5(4^{+})$-vertices, which do not receive any charge from $v$. Thus $\mu^{*}(v)\geq 4-7\times\frac{1}{3}-3\times\frac{1}{3}-2\times\frac{1}{4}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its $4(3^{+})$-neighbours by R9(c), at most $\frac{1}{4}$ to each of its $4(1)$-, $4(2)$- and $5(4^{+})$-neighbours by R9(a),(b),(d),(e).
(6.4). Let $m_{3}(v)=8$. By Lemma 2.37(b), $v$ has at most five $5(5)$-neighbours. If $v$ has at most four such neighbours, then $\mu^{*}(v)\geq 4-8\times\frac{1}{3}-4\times\frac{2}{9}-4\times\frac{1}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{2}{9}$ to each of its $5(5)$-neighbours by R10(b), and at most $\frac{1}{9}$ to each of its $4(3)$- and $5(4)$-neighbours by R10(a). If $v$ has five $5(5)$-neighbours, then $v$ has neither $4$-neighbour nor $5(4)$-neighbour by Lemma 2.37(a). Thus, $\mu^{*}(v)\geq 4-8\times\frac{1}{3}-5\times\frac{2}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and $\frac{2}{9}$ to each of its $5(5)$-neighbours by R10(b).
Acknowledgement
The author extends their gratitude to Kengo Aoki for pointing out a significant error in a previous proof of Lemma 2.3.
Declarations
Conflict of interest The author has no conflicts of interest to declare that are relevant to the content of this article.
Availability of data and material Data sharing not applicable to this article as no datasets were generated or analysed during the current study.